### v2

# FairPO: Robust Preference Optimization for Fair Multi-Label Learning

Abstract
Multi-label classification (MLC) often exhibits performance disparities, especially for infrequent or sensitive label categories. We propose FairPO, a novel framework that integrates preference-based loss formulations with group-robust optimization to improve fairness in multi-label classification (MLC), particularly targeting underperforming and sensitive label groups. FairPO partitions labels into a privileged set, targeted for enhanced performance, and a non-privileged set, where baseline performance is maintained. For privileged labels, a preference-based loss, inspired by Direct Preference Optimization (DPO), encourages model scores for true positives to be significantly higher than for their confusing negative counterparts, and scores for true negatives to be significantly lower than for their confusing positive counterparts—addressing hard examples that challenge standard classifiers. For non-privileged labels, a constrained objective ensures performance does not degrade substantially below a reference model. These group-specific objectives are balanced using a Group Robust Preference Optimization (GRPO) formulation, adaptively mitigating bias. Our experiments also include FairPO variants built on recent reference-free preference optimization techniques, namely Contrastive Preference Optimization (CPO) and Simple Preference Optimization (SimPO), which further highlight FairPO’s versatility [^1] Our code is available at GitHub: https://anonymous.4open.science/r/FairPO.
1 Introduction
Multi-label classification (MLC), the task of assigning a relevant subset of labels from a universe $\mathcal{T}$ to an instance $x$, is pervasive in modern machine learning, from image annotation (Wei et al., 2016) to document categorization (Kiritchenko and Matwin, 2006) and bioinformatics (Schietgat et al., 2010). Standard MLC models, parameterized by $\mathbf{w}$, are often trained by minimizing an aggregate loss, such as the sum of per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2013; Sorower, 2010; Tsoumakas and Katakis, 2007). However, such global optimization can obscure or even exacerbate significant fairness issues (Mehrabi et al., 2021; Chouldechova, 2017). Real-world label sets frequently include categories with imbalanced frequencies, differing societal importance, or varying inherent difficulty levels, which can result in models with disparate performance across labels (Zhang and Wu, 2021). For instance, a system might excel on common labels while consistently failing on rare but critical ones. This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
A key challenge lies in the model’s ability to discriminate effectively, especially for "hard" examples. Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions. For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a ‘confusing negative’ that the model mistakenly ranks above the true label. Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive). Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2012) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization. Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2023), FairPO recasts parts of the MLC task as learning explicit preferences. We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
Figure 1: FairPO’s methodology: Labels are split into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$). Preference loss ($\mathcal{L}_{\mathcal{P}}$) targets $\mathcal{P}$ labels; constrained loss ($\mathcal{L}_{\bar{\mathcal{P}}}$) maintains $\bar{\mathcal{P}}$ performance. GRPO adaptively balances these for fairer, robust MLC.
For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective. If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied. This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$. If no such confusing examples are found for $(x_{i},l)$, a standard BCE loss for label $l$ is applied to ensure consistent learning. For non-privileged labels $j\in\bar{\mathcal{P}}$, FairPO enforces a constraint that their classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ does not substantially exceed that of a reference model $\hat{\mathbf{w}}_{j}$, i.e., $\ell(\mathbf{w}_{j})\leq\ell(\hat{\mathbf{w}}_{j})+\epsilon$.
To manage the trade off between these two distinct objectives, FairPO leverages a group specific robust optimization technique—Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024). The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$. This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting robust fairness across the label partition (see Figure 1).
Our contributions are: (1) FairPO, a novel framework for fair MLC using preference signals and group robustness; (2) A conditional, preference-based objective for privileged labels targeting both confusing negatives and confusing positives, with a fallback to standard classification signals; (3) A robust optimization strategy to balance performance across privileged and non-privileged groups. Our comprehensive experiments, evaluating DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), and CPO (Xu et al., 2024) inspired preference objectives within FairPO, demonstrate its efficacy. Notably, FairPO, particularly the CPO-variant, achieves significant improvements in mAP and Exact Match Ratio for privileged labels—for instance, gaining up to 3.44% mAP on COCO (Lin et al., 2015) for the least frequent 20% of labels—while robustly preserving the performance metrics on non-privileged labels over the standard baseline. This underscores FairPO’s ability to effectively address challenging fairness-performance trade-offs in MLC. In Appendix A, we also briefly touch upon the framework’s applicability to fair multi-label attribute generation.
2 Related Work
Fairness in Multi-Label Classification: Ensuring fairness in MLC (Zhang and Wu, 2021; Tantithamthavorn et al., 2018) is complex due to multi-faceted label relationships. Recent efforts address MLC-specific fairness challenges, such as tackling label imbalance impacting tail labels (Wu et al., 2023), learning instance and class-level subjective fairness (Gao et al., 2023), or incorporating fairness in dynamic learning settings like class-incremental MLC (Masud et al., 2023). FairPO contributes by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets, applying distinct fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and managing them via a robustness framework. This targeted approach to enhancing performance for pre-defined critical labels, while safeguarding others, differentiates our work.
Preference Optimization: Preference optimization, especially for aligning LLMs (Ouyang et al., 2022; Christiano et al., 2017), has rapidly advanced. Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its reference-free variants like CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies directly from preference pairs. The field continues to evolve with methods like Identity Preference Optimization (IPO) for stability (Azhar et al., 2023), Kahneman-Tversky-based optimization (KTO) (Ethayarajh et al., 2024), simple yet effective Rejection Sampling Fine-Tuning (RFT) (Dong et al., 2023), and ongoing theoretical analyses of these methods (Zhao et al., 2024). We use preferences not to rank entire outputs, but to specifically differentiate true label scores from those of their confusing positive/negative counterparts within the privileged label set, thereby sharpening the model’s decision boundaries for critical distinctions.
Group Robust Optimization: Distributionally Robust Optimization (DRO) (Ben-Tal et al., 2009), particularly Group DRO (Sagawa et al., 2019), aims to improve worst-case performance across predefined data groups, enhancing fairness and robustness. This concept sees continued development, for instance, in improving its practicality (Zhang et al., 2022). Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024) extended this to LLM preference learning, balancing performance across preference groups. FairPO directly employs GRPO’s adaptive optimization strategy. However, our groups are defined by the label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$), and GRPO balances their distinct, custom-formulated loss objectives. This provides a principled mechanism for robustly managing the specific fairness-performance trade-offs in our MLC context.
3 Preliminaries: Preference Optimization Methods
Our framework builds upon recent preference optimization techniques. We briefly review the key formulations.
Direct Preference Optimization (DPO): DPO (Rafailov et al., 2023) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$. Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
$h_{\pi_{\theta}}(x,y_{w},y_{l})$ $=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)},$ (1)
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(h_{\pi_{\theta}}(x,y_{w},y_{l})\right)\right].$ (2)
where $\sigma$ is the sigmoid function and $\beta$ controls the deviation from $\pi_{\text{ref}}$.
Group Robust Preference Optimization (GRPO): GRPO (Ramesh et al., 2024) extends preference optimization to handle diverse preference groups $\{D_{g}\}_{g=1}^{K}$. Instead of minimizing the average loss, GRPO minimizes the worst-case loss across groups using a robust objective:
$\min_{\pi_{\theta}}\max_{\alpha\in\Delta_{K-1}}\sum_{g=1}^{K}\alpha_{g}L_{\text{Pref}}(\pi_{\theta};\pi_{\text{ref}},D_{g}),$ (3)
where $L_{\text{Pref}}$ is a base preference loss (like $L_{\text{DPO}}$), and $\alpha=(\alpha_{1},...,\alpha_{K})$ are adaptive weights in the probability simplex $\Delta_{K-1}$. The optimization dynamically increases weights $\alpha_{g}$ for groups with higher current loss, focusing learning on the worst-performing groups.
Simple Preference Optimization (SimPO): SimPO (Meng et al., 2024) aims to align the implicit reward with generation metrics and eliminates the need for $\pi_{\text{ref}}$. It uses the length-normalized average log-likelihood as the reward: $r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y|x)$. It also introduces a target margin $\gamma>0$ into the preference model. The resulting SimPO loss is:
$L_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$ (4)
Contrastive Preference Optimization (CPO): CPO (Xu et al., 2024) also removes the dependency on $\pi_{\text{ref}}$ for efficiency, approximating the DPO objective. It combines a reference-free preference loss with a negative log-likelihood (NLL) regularizer on preferred responses $y_{w}$ to maintain generation quality:
$L_{\text{prefer}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)\right]$ (5)
$L_{\text{NLL}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w})\sim D}[\log\pi_{\theta}(y_{w}|x)]$ (6)
$L_{\text{CPO}}(\pi_{\theta})$ $=L_{\text{prefer}}+L_{\text{NLL}}.$ (7)
This formulation avoids the computational cost of the reference model pass.
4 Methodology: Fair Preference Optimization (FairPO)
4.1 Problem Setup and Fairness Goals
Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is an input instance and $y_{i}=[y_{i1},\dots,y_{iT}]$ is the ground truth binary vector for $T$ labels, we aim to learn a model that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$, using parameters $\mathbf{w}_{t}$. These parameters are fine-tuned from a pretrained policy $\pi_{\theta}$, and we leverage a reference model with parameters $\hat{\mathbf{w}}_{t}$ producing scores $m(x_{i};\hat{\mathbf{w}}_{t})$. In practice, a pretrained model $\pi_{\theta}$ (e.g., an LLM) yields an input representation $\mathbf{z}_{i}=\pi_{\theta}(x_{i})\in\mathbb{R}^{d}$. A classifier head (e.g., FFN) with parameters $\mathbf{w}_{t}$ then computes the score $m(x_{i};\mathbf{w}_{t})$ for label $t$. The reference parameters $\hat{\mathbf{w}}$ are typically obtained from standard supervised fine-tuning (SFT) of $\pi_{\theta}$ on $\mathcal{D}$.
Our fairness goals are as follows: For any label $l\in\mathcal{P}$ (Privileged Group), the model should accurately discriminate its true state ($y_{il}=+1$ or $y_{il}=0$) from confusing alternatives. For any label $j\in\bar{\mathcal{P}}$ (Non-Privileged Group), the classification performance under $\mathbf{w}_{j}$ should remain close to that of $\hat{\mathbf{w}}_{j}$. To formalize the challenges for privileged labels, we define two types of confusing set for an instance $x_{i}$ and a privileged label $l\in\mathcal{P}$: If $y_{il}=+1$ (true positive), $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ is the set of confusing negatives. If $y_{il}=0$ (true negative), $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ is the set of confusing positives. The overall confusing set for $(x_{i},l)$ is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
4.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2023) preference loss. If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$). The DPO term is:
$h_{\mathbf{w}}(x_{i},l,k)=\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right).$ (8)
The loss contribution is $-\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
If $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$ is sampled, we prefer $l$ over $k$ (i.e., the true negative $l$ is preferred over the confusing positive $k$, meaning $m(x_{i};\mathbf{w}_{l})$ should be lower than $m(x_{i};\mathbf{w}_{k})$ relative to the reference). The DPO term is (note the swapped order for $k$ and $l$ to reflect $k\succ l$ effectively aiming for $m(x_{i};\mathbf{w}_{l})<m(x_{i};\mathbf{w}_{k})$):
$h_{\mathbf{w}}(x_{i},k,l)=\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right).$ (9)
The loss contribution is $-\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
No Confusing Examples ($S_{il}=\emptyset$): We revert to a standard base classification loss (e.g., Binary Cross-Entropy, BCE) for label $l$. This BCE fallback ensures that the model continues to receive learning signals even for labels without any identified confusing counterparts, thereby avoiding training stagnation.
$\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})=-[y_{il}\log m(x_{i};\mathbf{w}_{l})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}))].$ (10)
The overall Privileged Loss $\mathcal{L}_{\mathcal{P}}$ is the expectation over these conditional losses:
$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right.$
$\quad\left.+\mathbf{1}_{S_{il}=\emptyset}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\right]$ (11)
where $h_{\text{DPO}}$ refers to the appropriate term from Eq. 8 or Eq. 9. The hyperparameter $\beta$ controls the preference strength.
SimPO-Inspired Loss: We also adapt insights from Simple Preference Optimization (SimPO) (Meng et al., 2024). Original SimPO for generative models uses length-normalized log-likelihoods and a target margin $\gamma$ in its preference scoring. Since our multi-label classification (MLC) setup deals with individual label scores $m(x_{i};\mathbf{w}_{t})$ where sequence length is inapplicable, we adapt SimPO’s core concept of a target margin $\gamma$ to our context. This results in a reference-free preference component that aims to ensure a minimum separation for preferred label scores, omitting length normalization but retaining the margin:
$L_{\text{pref}}^{\text{SimPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)-\gamma\right),$ (12)
where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$. The term $\beta\log(\cdot)$ captures the relative preference, while $-\gamma$ enforces a desired separation margin. This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq. 11 when $S_{il}\neq\emptyset$, with the BCE fallback (Eq. 10) for $S_{il}=\emptyset$ remaining. More broadly, this adaptation leverages SimPO’s margin mechanism for more distinct score separation in challenging MLC cases.
CPO-Inspired Loss: Adapting Contrastive Preference Optimization (Xu et al., 2024). If $S_{il}\neq\emptyset$, the preference component is:
$L_{\text{pref}}^{\text{CPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)\right).$ (13)
(using the same preferred/dispreferred logic as SimPO). CPO also includes an NLL regularizer. Thus, the overall CPO-inspired privileged loss, $\mathcal{L}_{\mathcal{P}}^{\text{CPO}}$, could be formulated by combining this preference term (when $S_{il}\neq\emptyset$) with a consistent NLL/BCE component for all true ground truth labels within $\mathcal{P}$:
$\mathcal{L}_{\mathcal{P}}^{\text{CPO}}(\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},l)\text{ s.t. }l\in\mathcal{P}}\Big{[}$ $\mathbf{1}_{S_{il}\neq\emptyset}\cdot L_{\text{pref}}^{\text{CPO}}(x_{i},l,k_{s})+\lambda_{\text{CPO}}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\Big{]},$ (14)
where $k_{s}$ is a confusing example sampled from $S_{il}$, and $\lambda_{\text{CPO}}$ is a weighting hyperparameter. Note that the $\ell_{\text{BCE}}$ term is always applied, serving as both the NLL regularizer and the fallback for $S_{il}=\emptyset$ if the first term becomes zero. These alternatives offer different ways to model preferences. The non-privileged loss $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq. 17) remains unchanged regardless of the privileged loss choice.
4.3 Constrained Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$)
For non-privileged labels $j\in\bar{\mathcal{P}}$, our goal is to maintain performance relative to the reference model $\hat{\mathbf{w}}_{j}$, preventing degradation due to the focus on privileged labels. We use a standard base classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ (e.g., Binary Cross-Entropy) defined as:
$\ell(\mathbf{w}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\mathbf{w}_{j})+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}))],$ (15)
$\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\hat{\mathbf{w}}_{j})+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))].$ (16)
The Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$ employs a hinge mechanism, penalizing the model only if its loss $\ell(\mathbf{w}_{j})$ exceeds the reference model’s loss $\ell(\hat{\mathbf{w}}_{j})$ by more than a predefined slack margin $\epsilon\geq 0$:
$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|\exists t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t. }j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
This ensures $\mathbf{w}_{j}$ is primarily updated only if performance on label $j$ drops significantly below the reference baseline plus $\epsilon$.
4.4 Group Robust Optimization Formulation
To effectively balance the objective for privileged labels ($\mathcal{L}_{\mathcal{P}}$) and the constraint for non-privileged labels ($\mathcal{L}_{\bar{\mathcal{P}}}$), we employ the Group Robust Preference Optimization (GRPO) framework (Ramesh et al., 2024). We define two distinct groups corresponding to our label partitions:
  • $G_{\mathcal{P}}$: Associated with the Privileged Loss $\mathcal{L}_{\mathcal{P}}$, focusing on triplets $(x_{i},l,k)$.
  • $G_{\bar{\mathcal{P}}}$: Associated with the Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$, focusing on pairs $(x_{i},j)$.
The FairPO objective becomes the GRPO minimax problem:
[…]
Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss. The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}$ that perform well even under this worst-case weighting. This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.
4.5 Optimization Algorithm
The FairPO framework solves the minimax objective (Eq. 18) iteratively. A high-level overview is in Algorithm 1, with full details in Algorithm 2 (see Appendix B). In each training iteration, an instance and a label $r$ are sampled. The corresponding loss ($\mathcal{L}_{\mathcal{P}}^{(s)}$ or $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ via Eq. 11 or Eq. 17) and its gradient are computed based on whether $r$ is in the privileged set $\mathcal{P}$ or non-privileged set $\bar{\mathcal{P}}$.
The group weights $\alpha^{(s+1)}$ are updated using a mirror ascent step. A crucial aspect for the stability and effectiveness of this step is loss scaling. Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version. Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted ¯$\bar{\mathcal{L}}_{g}^{\text{avg}}$. The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability. This normalization prevents instability due to scale disparities in group loss values and makes the optimization more sensitive to relative improvements or regressions, rather than absolute magnitudes.
This scaled loss, say $\Delta\tilde{\mathcal{L}}_{g}^{(s)}$, is then used in the exponential update: $\alpha_{g}^{\text{new}}\leftarrow\alpha_{g}^{(s)}\exp(\eta_{\alpha}\Delta\tilde{\mathcal{L}}_{g}^{(s)})$, followed by normalization so $\sum_{g}\alpha_{g}^{\text{new}}=1$. This scaling normalizes loss magnitudes across groups, makes weight updates sensitive to significant performance changes relative to recent history, and improves overall stability of the $\alpha$ dynamics. Subsequently, model parameters $\mathbf{w}^{(s+1)}$ are updated via mirror descent, using a gradient that combines group gradients weighted by the adaptively balanced $\alpha^{(s+1)}$.
Algorithm 1 FairPO Training Overview (DPO-inspired)
1:Initialize: Model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(0)}$ (e.g., from supervised fine-tuning), set group weights $\alpha_{\mathcal{P}}^{(0)},\alpha_{\bar{\mathcal{P}}}^{(0)}$
2:Input: Dataset $\mathcal{D}$, reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$, hyperparameters $\beta,\epsilon,\eta_{\mathbf{w}},\eta_{\alpha}$.
3:For each training iteration $s=0,\dots,S-1$:
4:      Sample instance $(x_{i},y_{i})\sim\mathcal{D}$ and a label $r\in\mathcal{T}$.
5:      If $r\in\mathcal{P}$:
6:          Compute privileged loss $\mathcal{L}_{\mathcal{P}}^{(s)}$ for $(x_{i},r)$ (Eq. 11, conditionally using preference loss like Eq. 8/9 or BCE Eq. 10).
7:      Else if $r\in\bar{\mathcal{P}}$:
8:          Compute non-privileged loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ for $(x_{i},r)$ (Eq. 17).
9:      Update group weights $\alpha^{(s+1)}$ via mirror ascent using $\mathcal{L}_{\mathcal{P}}^{(s)},\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ (GRPO step).
10:      Update model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(s+1)}$ via mirror descent using weighted gradients.
11:End For
12:Return Optimized parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(S)}$.
13:Full details are in Algorithm 2 (see Appendix B).
5 Experimental Setup
Datasets and Preprocessing: We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014222MS-COCO: https://cocodataset.org/ (80 object categories) (Lin et al., 2015) and NUS-WIDE [^3] NUS-WIDE: https://github.com/NExTplusplus/NUS-WIDE/ (81 concept labels) (Chua et al., 2009). For partitioning, we designate the 20% least frequent labels in the training set of each dataset as the privileged group ($\mathcal{P}$), with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$). We use standard train/validation/test splits for both datasets (see Appendix C for more details).
Model and Implementation Details: Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021) pretrained on ImageNet [^4] ViT Model: https://huggingface.co/docs/transformers/en/model_doc/vit. The FairPO framework, including the classifier head and label-specific parameters $\mathbf{w}_{t}$, is fine-tuned on top of the frozen ViT features. The reference parameters $\hat{\mathbf{w}}$ for FairPO are obtained by standard supervised fine-tuning (SFT) of the same ViT architecture with a BCE loss on the target MLC dataset. Hyperparameters like learning rates ($\eta_{\mathbf{w}},\eta_{\alpha}$), DPO’s $\beta$, SimPO’s $\gamma$, CPO’s $\lambda_{\text{CPO}}$, and the non-privileged slack $\epsilon$ are tuned via a grid search on the validation set (see Appendix F for more details).
Evaluation Metrics: Model performance is assessed using established multi-label metrics, calculated separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ label sets. These include Mean Average Precision (mAP), which measures the average precision across all labels; Sample F1 Score (F1-Sample), the harmonic mean of precision and recall averaged per sample; Exact Match Ratio (EMR), the proportion of samples with perfectly predicted label sets; and Multi-label Accuracy (Accuracy), representing element-wise prediction correctness averaged over all labels and instances.
Baselines: FairPO is compared against: (1) BCE-SFT, standard per-label Binary Cross-Entropy (BCE) fine-tuning, which also provides reference parameters $\hat{\mathbf{w}}$. (2) BCE-SFT + Privileged Re-Weighting (RW), which statically increases loss weights for the privileged group $\mathcal{P}$. (3) Group DRO + BCE (Sagawa et al., 2019), applying Group DRO to minimize the worst-group BCE loss between $\mathcal{P}$ and $\bar{\mathcal{P}}$ (see Appendix D for more details).
FairPO Variants: We evaluate three main variants of our FairPO framework, each differing in the specific preference loss formulation used for the privileged group $\mathcal{P}$ (as detailed in Section 4.2): FairPO-DPO utilizes the DPO-inspired conditional preference loss (Eq. 11). FairPO-SimPO employs a SimPO-inspired (Meng et al., 2024) reference-free loss, adapting the preference term from Eq. 12 within the conditional structure of Eq. 11. Lastly, FairPO-CPO uses a CPO-inspired (Xu et al., 2024) reference-free loss combined with an NLL-like regularization component, as formulated in Eq. 14 (see Appendix E for more details).
6 Results and Analysis
Overview and Performance Summary: We evaluate FairPO against baselines on MS-COCO (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009). Tables 1 and 2 detail performance for privileged ($\mathcal{P}$, 20% least frequent) and non-privileged ($\bar{\mathcal{P}}$, 80% most frequent) label sets. The $\Delta\text{mAP}(\mathcal{P})$ column highlights mAP gains for $\mathcal{P}$ labels over the BCE SFT baseline. A consistent trend emerges: FairPO variants, particularly FairPO-CPO, significantly improve $\mathcal{P}$ set performance while robustly maintaining $\bar{\mathcal{P}}$ set metrics. This addresses the common trade-off where gains in one group detrimentally affect the other, a challenge often faced by simpler baselines.
Table 1: Performance comparison on MS-COCO. $\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%). Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 86.32 90.65 61.43 64.89 94.89 98.12 35.78 36.98 Ref BCE SFT + RW 87.25 89.85 62.68 64.11 95.95 97.93 47.43 33.81 +0.93 GDRO + BCE 87.92 90.41 62.31 64.75 95.72 98.05 46.12 35.16 +1.60 FairPO-DPO 88.02 89.97 63.45 63.65 97.89 98.04 62.19 35.12 +1.70 FairPO-SimPO 87.67 88.76 62.34 63.12 95.69 97.78 45.32 32.34 +1.35 FairPO-CPO 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44
Table 2: Performance comparison on NUS-WIDE. $\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%). Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 63.53 70.24 48.12 55.83 91.51 95.22 19.32 11.56 Ref BCE SFT + RW 65.12 69.14 49.51 54.73 92.33 94.81 21.23 10.32 +1.59 GDRO + BCE 64.84 69.91 49.13 55.62 92.11 95.13 21.02 11.34 +1.31 FairPO-DPO 66.34 69.05 51.71 54.52 93.92 95.04 27.91 11.21 +2.81 FairPO-SimPO 64.11 68.03 48.82 53.81 91.94 94.52 20.18 10.19 +0.58 FairPO-CPO 67.12 69.83 52.21 55.24 94.31 95.12 31.87 11.25 +3.59
Comparative Analysis of FairPO Variants: The FairPO variants differ in their preference mechanisms: FairPO-DPO utilizes a reference model $\hat{\mathbf{w}}$ in its log-ratio preference term. This relative formulation can stabilize training by anchoring scores to a pre-trained baseline, making the preference signal robust to the absolute scale of model outputs. This likely contributes to its strong EMR-$\mathcal{P}$ gains, indicating better resolution of complex multi-label confusion. FairPO-CPO, often the top performer, employs a reference-free contrastive term $\log(m(x_{i};\text{preferred})/m(x_{i};\text{dispreferred}))$. Its strength is amplified by an NLL-like regularizer (realized as $\ell_{\text{BCE}}$ for all privileged labels in Eq. 14). This dual focus on relative preference and absolute correctness for $\mathcal{P}$ underpins its superior $\mathcal{P}$ metrics while maintaining robust $\bar{\mathcal{P}}$ performance. FairPO-SimPO, another reference-free method using a fixed margin $\gamma$, generally shows improvements but lags behind DPO and CPO variants. This suggests that a fixed margin without DPO’s reference anchoring or CPO’s explicit NLL-like regularization might be less robust or harder to tune optimally for diverse preference pairs in MLC.
Impact on Exact Match Ratio: The substantial improvements in EMR-$\mathcal{P}$ by FairPO-DPO and FairPO-CPO are particularly significant. EMR, requiring perfect prediction for the $\mathcal{P}$ label subset, is highly sensitive to confusion. FairPO’s preference optimization, by directly addressing ambiguities between true labels and their specific confusing counterparts, enhances discriminative power for these challenging cases. This leads to more accurate complete label set predictions for the privileged group. The GRPO mechanism ensures these targeted gains do not disproportionately penalize the $\bar{\mathcal{P}}$ group, demonstrating FairPO’s capacity to learn a fairer and more robust multi-label classifier.
7 Ablation Studies
We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions. Results are in Table 3 and Table 4.
Table 3: Ablation on core components of FairPO-CPO (MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT. Parentheses show change vs Full FairPO-CPO.
Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08)
Table 3 shows that removing the preference loss for $\mathcal{P}$ (using BCE instead) significantly reduces $\mathcal{P}$ gains, underscoring its criticality. Removing the $\bar{\mathcal{P}}$ constraint causes a notable drop in $\bar{\mathcal{P}}$ performance, affirming its protective role. Using fixed weights instead of GRPO degrades performance for both groups, highlighting GRPO’s superior adaptive balancing.
Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT (86.32). Parentheses show change vs Full FairPO-CPO.
Preference Detail Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 (Conf. Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 22.54 35.10 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) (-42.89) (-0.13) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 60.75 34.90 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07) (-4.68) (-0.33)
Table 4 details the impact of specific choices within the $\mathcal{P}$ loss formulation. Restricting preference learning to only Confusing Negatives (when $y_{il}=+1$) severely curtails its effectiveness. Given that positive instances of privileged labels can be extremely rare, this makes the preference signal exceptionally sparse. This approach neglects to train the model on ranking true negatives ($y_{il}=0$) against confusing positives—a scenario that forms the vast majority of instances for these labels. This omission of crucial learning signals for most situations involving $\mathcal{P}$ labels results in a profound degradation in performance for this set. Omitting the BCE Fallback for non-confusing $\mathcal{P}$ instances (i.e., no loss if $S_{il}=\emptyset$) also clearly degrades performance. This standard classification signal for easier cases is vital for model stability and reinforcing correct predictions within $\mathcal{P}$ when direct preference learning is not active. Both aspects of our full preference formulation are thus shown to be essential. This demonstrates that preference loss alone is insufficient for stable training, and standard losses play a critical complementary role even under targeted preference optimization.
8 Discussion
Conclusion: We introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification. Our experiments, particularly with the FairPO-CPO variant, underscore the efficacy of this approach in navigating complex fairness-performance trade-offs, highlighting the value of nuanced preference signals for challenging discrimination tasks.
Limitations: Despite its promising results, FairPO presents technical limitations. The dynamic construction of the ‘confusing set’ $S_{il}$ based on evolving model scores can lead to optimization instability or sparse preference signals, especially for rare/easily-separated privileged labels. Furthermore, the efficacy of FairPO-DPO and the non-privileged constraint hinges on a well-calibrated reference model $\hat{\mathbf{w}}$; its deficiencies can distort DPO log-ratios or misguide the constraint. Finally, GRPO’s adaptive balancing of heterogeneous loss terms ($\mathcal{L}_{\mathcal{P}}$ combining preference and BCE, $\mathcal{L}_{\bar{\mathcal{P}}}$ with a hinge loss) may require careful tuning of $\eta_{\alpha}$ and loss scaling for stable convergence, the theoretical properties of which remain an open area. While these limitations highlight the complexity of fair MLC optimization, they also suggest promising directions for refinement and extension, which we talk about next.
Future Work: Future work will focus on several key directions. A primary goal is to extend FairPO’s principles to multi-label attribute generation (see Appendix A for a conceptual outline), aiming to foster fair and diverse outputs. We also plan comprehensive empirical validation across a wider array of datasets and modalities, including text, to assess generalizability. Further research will involve in-depth analysis of different label partitioning strategies beyond frequency, and pursuing theoretical insights into convergence properties for FairPO.
Acknowledgments and Disclosure of Funding
The authors would like to express their gratitude to the members of the CSE Department and CMInDS at IIT Bombay for their support and for providing a conducive research environment. We also thank the anonymous reviewers for their valuable feedback which helped improve the quality of this manuscript. We acknowledge the use of high-performance computing resources that were instrumental in conducting the experiments presented in this work.
Funding Disclosure This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors for the submitted work, beyond the general support and infrastructure provided by IIT Bombay to its faculty and students.
Competing Interests The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Adapting FairPO for Multi-Attribute Generation
This section outlines a key direction for future research: adapting the FairPO framework to multi-attribute generation tasks. The methodologies described here are conceptual and represent a planned extension of our work, building directly upon the principles established in the main paper for classification. Experimental validation of this approach is part of our future agenda.
In such a generative setting, the goal would be to produce an output sequence $y$ from a prompt $x$ using a single generative policy $\pi_{\mathbf{w}}(y|x)$, such that the generation aligns with fairness goals defined over a set of attributes $\mathcal{A}$. As in the classification setting, a potential approach involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets. The core GRPO minimax structure (Eq. 18) could be retained, with group losses defined based on a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$ and the attributes $j_{i}$ driving the preferences.
A primary consideration in this adaptation would be how preference losses are applied, particularly for the privileged group, as DPO would then operate on the log-probabilities of entire generated sequences rather than individual label scores.
Proposed Privileged Loss ($\mathcal{L}_{\mathcal{P}}$): For future generative applications, the aim for privileged attributes $j\in\mathcal{P}$ would be to ensure that the learned policy $\pi_{\mathbf{w}}$ strongly reflects observed preferences $y_{w}\succ y_{l}$ when the preference was established based on such an attribute. This could translate to assigning significantly higher relative log-probability to the preferred sequence $y_{w}$ compared to the dispreferred sequence $y_{l}$, relative to a reference policy $\pi_{\text{ref}}$.
This could be achieved by utilizing a standard DPO loss formulation, averaged over the subset of the preference data $\mathcal{D}_{pref}$ where the driving attribute $j$ belongs to the privileged set $\mathcal{P}$. The privileged loss could then be defined as:
$\mathcal{L}_{\mathcal{P}}(\pi_{\mathbf{w}},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l},j)\sim\mathcal{D}_{pref}\mid j\in\mathcal{P}}\left[-\log\sigma\left(\beta\cdot h_{\pi_{\mathbf{w}}}(x,y_{w},y_{l})\right)\right]$ (19)
Minimizing such a loss would directly encourage the model $\pi_{\mathbf{w}}$ to favor preferred sequences over dispreferred ones for preferences driven by privileged attributes, relative to the reference model $\pi_{\text{ref}}$.
Proposed Non-Privileged Loss ($\mathcal{L}_{\bar{\mathcal{P}}}$): The objective for non-privileged attributes $k\in\bar{\mathcal{P}}$ would remain analogous to the classification setting: to prevent significant degradation compared to a baseline. The non-privileged loss could then utilize a hinge formulation based on the DPO loss for preferences driven by attributes $k\in\bar{\mathcal{P}}$:
$\mathcal{L}_{\bar{\mathcal{P}}}(\pi_{\mathbf{w}},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l},k)\sim\mathcal{D}_{pref}\mid k\in\bar{\mathcal{P}}}\left[\max\left(0,\mathcal{L}_{DPO}(\pi_{\mathbf{w}},\pi_{\text{ref}};x,y_{w},y_{l})-(\log 2)-\epsilon^{\prime}\right)\right].$ (20)
This would penalize the model only if its preference modeling for non-privileged attributes degrades substantially beyond the baseline performance (represented by $\log 2$ for random preference) plus a slack $\epsilon^{\prime}$.
The overall FairPO objective for generation would then employ the GRPO minimax formulation to balance $\mathcal{L}_{\mathcal{P}}$ (Eq. 19) and $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq. 20), aiming for robust alignment across both privileged and non-privileged attributes identified in the preference data.
Planned Task and Potential Datasets: Future work will focus on evaluating FairPO for generating text $y$ based on a prompt $x$, aiming for alignment with attributes $\mathcal{A}=\mathcal{P}\cup\bar{\mathcal{P}}$. This would require preference datasets $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}$ where the driving attribute $j_{i}$ for the preference is known. Potential datasets for such an investigation include:
  • Publicly available RLHF datasets (e.g., subsets of Anthropic HH-RLHF [Bai et al., 2022], Summarization preferences [Stiennon et al., 2020]), where attributes like ‘Helpfulness’, ‘Harmlessness’, ‘Conciseness’, ‘Faithfulness’ would need to be partitioned into $\mathcal{P}$ and $\bar{\mathcal{P}}$.
  • Synthetic or newly annotated preference data where attribute labels are explicitly assigned during collection, tailored for this specific fairness investigation.
Prospective Model and Implementation Strategy: We plan to use pretrained generative language models (e.g., Gemma [Team et al., 2024], Llama [Touvron et al., 2023], Mistral [Jiang et al., 2023]) as the base policy $\pi_{\theta}$. These models would then be fine-tuned using the FairPO objective proposed for generation (Equations 19 and 20). The reference policy $\pi_{\text{ref}}$ would typically be the SFT version of the base model.
Proposed Evaluation and Baselines: Evaluation in this future work would primarily rely on preference-based metrics. Comparisons would include:
  • Win rates (human or model-based, e.g., using GPT-4) for preferences driven by privileged attributes $j\in\mathcal{P}$.
  • Win rates for preferences driven by non-privileged attributes $k\in\bar{\mathcal{P}}$.
  • The worst-group win rate across the two attribute groups.
Baselines for comparison would likely include standard non-robust DPO fine-tuning and the SFT reference model.
Future experiments within this generative setting would also investigate the impact of hyperparameters like $\beta$ and $\epsilon^{\prime}$, and the potential effectiveness of adapting the alternative SimPO and CPO loss formulations to the FairPO framework for generation.
Appendix B FairPO Algorithm
The FairPO framework is trained iteratively to solve the minimax objective presented in Eq. 18. The detailed procedure, inspired by the DPO-based variant of FairPO, is provided in Algorithm 2.
The core training loop proceeds as follows:
  1. 1.
    Initialization: Model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}$ are initialized (e.g., from a reference model $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$), and the adaptive group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are typically set to uniform values (e.g., 0.5 each).
  2. 2.
    Iterative Updates: For each training step:
    • An instance $(x_{i},y_{i})$ is sampled from the dataset $\mathcal{D}$. A label $r\in\mathcal{T}$ is then randomly selected for processing from this instance.
    • Privileged Label Processing (if $r\in\mathcal{P}$): The algorithm first identifies if any confusing examples $S_{il}$ exist for the chosen privileged label $l=r$ with respect to other labels (Lines 11-17 in Alg. 2). If $S_{il}\neq\emptyset$, a DPO-inspired preference loss (Eq. 8 or 9) is computed between label $l$ and a sampled confusing example $k\in S_{il}$ (Lines 19-27). This loss encourages the model to correctly rank $l$ against $k$. If $S_{il}=\emptyset$ (no confusing examples found), a standard base classification loss (e.g., BCE, Eq. 10) is used for label $l$ to ensure continued learning (Lines 29-31). The resulting loss contributes to $\mathcal{L}_{\mathcal{P}}^{(s)}$.
    • Non-Privileged Label Processing (if $r\in\bar{\mathcal{P}}$): For a chosen non-privileged label $j=r$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed as per Eq. 17, penalizing deviations from the reference model’s performance beyond a slack $\epsilon$ (Lines 34-37).
    • Adaptive Weight Update (GRPO): The group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated using a mirror ascent step (exponential weighting based on the current step’s group losses $\mathcal{L}_{\mathcal{P}}^{(s)}$ and $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ should be scaled for stable training as explained in Section 4) and then normalized to sum to one (Lines 39-41). This dynamically adjusts the focus on the group currently performing worse.
    • Model Parameter Update: The model parameters $\mathbf{w}_{t}$ for all labels $t\in\mathcal{T}$ are updated via a mirror descent step (gradient descent), using a combined gradient weighted by the updated $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ (Line 42).
  3. 3.
    The process repeats for a predefined number of iterations or until convergence.
This iterative procedure allows FairPO to dynamically balance the specialized objectives for privileged labels against the performance constraints for non-privileged labels, guided by the GRPO mechanism to achieve robust fairness. For variants like FairPO-SimPO or FairPO-CPO, the DPO-inspired preference loss component (Lines 21-26) for privileged labels is replaced with their respective preference formulations (e.g., Eq. 12 or 13, possibly combined with an NLL-like term as in Eq. 14), while the overall GRPO structure and non-privileged handling remain consistent.
Algorithm 2 FairPO Algorithm for Multi-Label Classification (DPO-inspired)
1:Initialize: model parameters $\{\mathbf{w}_{t}^{(0)}\in\mathbb{R}^{d}|\forall t\in\mathcal{T}\}$ (e.g., copy $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$), group weights $\alpha_{\mathcal{P}}^{(0)}\leftarrow 0.5,\alpha_{\bar{\mathcal{P}}}^{(0)}\leftarrow 0.5$.
2:Choose: learning rates $\eta_{\mathbf{w}},\eta_{\alpha}$, DPO hyperparameter $\beta$, reference parameters $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$, constraint slack $\epsilon$.
3:for $s=0$ to $S$ (MaxIterations) do
4:      Sample an example: $(x_{i},[y_{i1},\dots,y_{iT}])\in\mathcal{D}\sim p_{\mathcal{D}}(.)$.
5:      Initialize group losses for this step: $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow 0$, $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow 0$.
6:      Initialize gradients: $g^{t}_{\mathcal{P}}\leftarrow\vec{0}$, $g^{t}_{\bar{\mathcal{P}}}\leftarrow\vec{0}\quad\forall t\in\mathcal{T}$.
7:      Forward pass: $m(x_{i};\mathbf{w}_{t}^{(s)})\leftarrow\sigma(\mathbf{w}_{t}^{(s)^{T}}\mathbf{z}_{i})$ where $\mathbf{z}_{i}\leftarrow\pi_{\theta}(x_{i})\quad\forall t\in\mathcal{T}$.
8:      Sample a label: $r\in\mathcal{T}\sim Uniform(\frac{1}{|\mathcal{T}|})$.
9:      if $r\in\mathcal{P}$ then $\triangleright$ Handle privileged label
10:          $l\leftarrow r$
11:          $S_{il}^{\text{neg}}\leftarrow\emptyset$, $S_{il}^{\text{pos}}\leftarrow\emptyset$
12:          if $y_{il}=+1$ then $\triangleright$ True Positive case for privileged label $l$
13:              $S_{il}^{\text{neg}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\geq m(x_{i};\mathbf{w}_{l}^{(s)})\}$
14:              $S_{il}\leftarrow S_{il}^{\text{neg}}$
15:          else if $y_{il}=0$ then $\triangleright$ True Negative case for privileged label $l$
16:              $S_{il}^{\text{pos}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\leq m(x_{i};\mathbf{w}_{l}^{(s)})\}$
17:              $S_{il}\leftarrow S_{il}^{\text{pos}}$
18:          end if
19:          if $S_{il}\neq\emptyset$ then $\triangleright$ Confusing examples exist, use DPO-inspired loss
20:              Sample $k\in S_{il}\sim Uniform(\frac{1}{|S_{il}|})$
21:              if $y_{il}=+1$ then $\triangleright$ DPO for True Positive $l$ vs Confusing Negative $k$
22:                  $h_{\mathbf{w}^{(s)}}(x_{i},l,k)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)$.
23:                  $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},l,k)\right)$
24:              else if $y_{il}=0$ then $\triangleright$ DPO for True Negative $l$ vs Confusing Positive $k$
25:                  $h_{\mathbf{w}^{(s)}}(x_{i},k,l)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)$.
26:                  $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},k,l)\right)$
27:              end if
28:              $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{pref}}$
29:              $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{pref}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
30:          else $\triangleright$ No confusing examples, use BCE loss for privileged label $l$
31:              $\mathcal{L}_{\text{BCE}}\leftarrow-[y_{il}\log m(x_{i};\mathbf{w}_{l}^{(s)})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}^{(s)}))]$
32:              $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{BCE}}$
33:              $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{BCE}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
34:          end if
35:      else if $r\in\bar{\mathcal{P}}$ then $\triangleright$ Handle non-privileged label
36:          $j\leftarrow r$
37:          $\ell(\mathbf{w}_{j}^{(s)})\leftarrow-[y_{ij}\log(m(x_{i};\mathbf{w}_{j}^{(s)}))+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}^{(s)}))]$
38:          $\ell(\hat{\mathbf{w}}_{j})\leftarrow-[y_{ij}\log(m(x_{i};\hat{\mathbf{w}}_{j}))+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))]$
39:          $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow\max\left(0,\,\ell(\mathbf{w}_{j}^{(s)})-\ell(\hat{\mathbf{w}}_{j})-\epsilon\right)$
40:          $g_{\bar{\mathcal{P}}}^{t}\leftarrow g_{\bar{\mathcal{P}}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
41:      end if
42:      $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\alpha_{\mathcal{P}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\mathcal{P},scaled}^{(s)})$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\alpha_{\bar{\mathcal{P}}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\bar{\mathcal{P}},scaled}^{(s)})$triangleright Mirror ascent
43:      $Z\leftarrow\alpha_{\mathcal{P}}^{(s+1)}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}$
44:      $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\frac{\alpha_{\mathcal{P}}^{(s+1)}}{Z}$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\frac{\alpha_{\bar{\mathcal{P}}}^{(s+1)}}{Z}$triangleright Weight normalization
45:      $\mathbf{w}_{t}^{(s+1)}\leftarrow\mathbf{w}_{t}^{(s)}-\eta_{\mathbf{w}}(\alpha_{\mathcal{P}}^{(s+1)}g_{\mathcal{P}}^{t}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}g_{\bar{\mathcal{P}}}^{t})$$\forall t\in\mathcal{T}$$\triangleright$ Mirror descent
46:end for
47:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$
Appendix C Dataset and Preprocessing Details
MS-COCO 2014 [Lin et al., 2015]: We used the official 2014 train/val splits. The training set contains 82,783 images and the validation set (used as our test set) contains 40,504 images. There are 80 object categories. The privileged set $\mathcal{P}$ consisted of the 16 labels (20% of 80) with the lowest frequency in the training set. The remaining 64 labels formed $\bar{\mathcal{P}}$.
NUS-WIDE [Chua et al., 2009]: This dataset contains 269,648 images with 81 concept labels. We used the common split of 161,789 images for training and 107,859 for testing. The privileged set $\mathcal{P}$ consisted of the 16 labels (approx. 20% of 81) with the lowest frequency in the training set. The remaining 65 labels formed $\bar{\mathcal{P}}$.
Image Preprocessing: For both datasets, images were resized to $224\times 224$ pixels and normalized using the standard ImageNet mean and standard deviation, consistent with the ViT pretraining. Standard data augmentations like random horizontal flips and random resized crops were applied during training.
Appendix D Baseline Experimental Details
This section provides further details on the implementation and hyperparameter tuning for the baseline methods used in our experiments. Unless otherwise specified, all baselines were trained using the same base Vision Transformer (ViT) architecture, optimizer (AdamW), number of epochs, and batch size as the main FairPO experiments for fair comparison. Hyperparameters specific to each baseline were tuned on the validation set of MS-COCO and NUS-WIDE. The privileged group $\mathcal{P}$ comprised the 20% least frequent labels, and the non-privileged group $\bar{\mathcal{P}}$ comprised the remaining 80%.
D.1 BCE-SFT (Reference Model)
The BCE-SFT baseline represents standard supervised fine-tuning.
  • Objective: The model is trained by minimizing the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels. For an instance $(x_{i},y_{i})$, the loss is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$, where $m(x_{i};\mathbf{w}_{t})$ is the model’s predicted probability for label $t$.
  • Implementation: A separate non-linear classifier head (same MLP as used in FairPO’s main experiments) is trained on top of the frozen ViT features for each label $t$, with parameters $\mathbf{w}_{t}$.
  • Hyperparameters: The primary hyperparameter tuned was the learning rate for the classifier heads. This was selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ based on the best overall mean Average Precision (mAP) on the validation set. Weight decay was set consistent with FairPO experiments (e.g., 0.01).
  • Role: This model serves two purposes:
    1. 1.
      It acts as a direct performance baseline.
    2. 2.
      The final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used by the FairPO framework (for DPO-based losses and the non-privileged constraint).
  • Training: Trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
D.2 BCE-SFT + Privileged Re-Weighting (RW)
This baseline modifies the standard BCE-SFT by applying static loss re-weighting.
  • Objective: The goal is to improve performance on the privileged labels $\mathcal{P}$ by assigning a higher weight to their contribution in the total loss. The loss function is:
    $\mathcal{L}_{\text{RW}}=\lambda_{\mathcal{P}}\sum_{l\in\mathcal{P}}\text{BCE}(m(x_{i};\mathbf{w}_{l}),y_{il})+\lambda_{\bar{\mathcal{P}}}\sum_{j\in\bar{\mathcal{P}}}\text{BCE}(m(x_{i};\mathbf{w}_{j}),y_{ij})$
    Typically, $\lambda_{\bar{\mathcal{P}}}$ is set to $1$, and $\lambda_{\mathcal{P}}>1$.
  • Implementation: Same architecture and per-label classifiers as BCE-SFT. The loss computation is modified to incorporate the weights.
  • Hyperparameters:
    • Learning rate: Tuned as for BCE-SFT.
    • Privileged group weight $\lambda_{\mathcal{P}}$: This is the key hyperparameter. It was tuned via grid search on the validation set, selecting from values such as $\{2,3,5,8,10\}$. The value was chosen to maximize mAP on the privileged group $\mathcal{P}$ without causing an excessive drop in mAP on the non-privileged group $\bar{\mathcal{P}}$, or based on the best overall mAP. For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective. $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$.
  • Training: Trained similarly to BCE-SFT.
D.3 Group DRO + BCE
This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss.
  • Objective: To minimize the worst-case BCE loss across the predefined groups (privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$). The objective is:
    $\min_{\{\mathbf{w}_{t}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\geq 0,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\text{BCE}}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})\right]$
    where $\mathcal{L}_{\text{BCE}}(\mathcal{G})=\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}\text{BCE}(m(x_{i};\mathbf{w}_{g}),y_{ig})$ is the average BCE loss for labels within group $\mathcal{G}$.
  • Implementation:
    • Architecture: Same as BCE-SFT.
    • Loss Calculation: Per-label BCE losses are computed. These are then aggregated (averaged) separately for all labels in $\mathcal{P}$ to get $\mathcal{L}_{\text{BCE}}(\mathcal{P})$ for the current batch/sample, and similarly for $\bar{\mathcal{P}}$ to get $\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})$.
    • Optimization: Model parameters $\{\mathbf{w}_{t}\}$ are updated using AdamW. The group weights $\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step (exponential weighting based on the current group losses, followed by normalization):
      $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$
      The final loss for backpropagation through model parameters is $\alpha_{\mathcal{P}}^{(s+1)}\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}^{(s+1)}\mathcal{L}_{\text{BCE}}^{(s)}(\bar{\mathcal{P}})$.
  • Hyperparameters:
    • Learning rate for model parameters $\{\mathbf{w}_{t}\}$: Tuned as for BCE-SFT.
    • Learning rate for group weights $\eta_{\alpha}$: Tuned from $\{0.01,0.05,0.1,0.2\}$ on the validation set, similar to FairPO’s $\eta_{\alpha}$. The selection aimed for stable convergence and good worst-group performance. For Group DRO with BCE, the raw group losses might be directly used in the exponential update, or a similar loss scaling/normalization approach as discussed for FairPO’s GRPO step could be employed for improved stability if raw loss magnitudes vary significantly between groups or over time. For the reported results, $\eta_{\alpha}$ was set to $0.05$ after tuning.
  • Training: Trained similarly to BCE-SFT and FairPO regarding epochs and batching.
This section provides further details on the experimental setup for the main FairPO variants (FairPO-DPO, FairPO-SimPO, and FairPO-CPO), complementing the information in Section 5 of the main paper.
Appendix E FairPO Experimental Details
E.1 Common Setup for All FairPO Variants
Unless specified otherwise, the following setup details apply to all FairPO variants:
  • Base Model: A Vision Transformer (ViT) model, specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k, was used as the feature extractor. The ViT backbone (i.e., the ‘model.vit‘ component) was kept frozen during fine-tuning with FairPO, except for its final encoder layer, which was made trainable to allow for some adaptation of higher-level features.
  • Datasets and Preprocessing: Experiments were conducted on MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009]. Images were resized to $224\times 224$ pixels and normalized using ImageNet statistics. Standard data augmentations (random horizontal flips, random resized crops) were applied during training. Details on dataset splits and privileged/non-privileged partitioning (20% least frequent labels as privileged) are provided in Section 5 and Appendix C.
  • Optimizer: The AdamW optimizer [Loshchilov and Hutter, 2019] was used for updating the trainable parameters (classifier heads and the last ViT encoder layer).
  • Training Schedule: Models were trained for a maximum of 25 epochs with a batch size of 32. Early stopping was employed based on the overall mAP (average of mAP on privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets) on the validation set, with a patience of 5 epochs.
  • Reference Model ($\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$): The reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$ required by FairPO-DPO and for the non-privileged loss constraint (Eq. 17) in all variants were obtained from a BCE-SFT model (see Appendix D). This SFT model used the same ViT architecture and per-label MLP classifier heads (detailed below) and was trained with a standard BCE loss.
E.2 Per-Label Non-Linear MLP Classifier Head
For each of the $T$ labels in the dataset, a dedicated, independent non-linear Multi-Layer Perceptron (MLP) head was used to predict the probability of that label being positive. This allows each label to have its own specialized classifier operating on the shared image features extracted by the ViT.
  • Architecture: Each MLP head takes the $d$-dimensional output feature vector from the ViT’s [CLS] token (where $d$ is the hidden size of ViT, e.g., 768 for ViT-Base) as input and outputs a single logit for the corresponding label. The architecture of each MLP head is as follows:
    1. 1.
      Linear Layer: $d\rightarrow 256$ neurons
    2. 2.
      ReLU Activation
    3. 3.
      Linear Layer: $256\rightarrow 64$ neurons
    4. 4.
      ReLU Activation
    5. 5.
      Linear Layer: $64\rightarrow 16$ neurons
    6. 6.
      ReLU Activation
    7. 7.
      Linear Layer: $16\rightarrow 4$ neurons
    8. 8.
      ReLU Activation
    9. 9.
      Linear Layer (Output): $4\rightarrow 1$ neuron (producing the logit)
    The final probability score $m(x_{i};\mathbf{w}_{t})$ for label $t$ is obtained by applying a sigmoid function to this output logit.
  • Parameters: The parameters $\mathbf{w}_{t}$ for each label $t$’s MLP head are specific to that label and are learned during the fine-tuning process. The set of all such parameters is denoted by $\{\mathbf{w}_{t}|\forall t\in\mathcal{T}\}$.
  • Trainability: All parameters of these MLP heads were trainable during both SFT (for the reference model) and FairPO fine-tuning.
  • Motivation: Using separate MLP heads rather than a single shared linear layer allows for more complex, non-linear decision boundaries tailored to each label’s specific characteristics and inter-dependencies with the input features. This can be particularly beneficial for labels with varying difficulty or those requiring nuanced feature interpretations.
E.3 FairPO Variant-Specific Hyperparameters
The core FairPO learning rates ($\eta_{\mathbf{w}}$ for model parameters, $\eta_{\alpha}$ for GRPO weights) were tuned via grid search. For $\eta_{\mathbf{w}}$ (AdamW LR for classifier heads and last ViT layer), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$. For $\eta_{\alpha}$, values from $\{0.01,0.05,0.1,0.2\}$ were explored. The final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$ after tuning.
Specific hyperparameters for each FairPO variant’s privileged loss component were tuned as per Table 5 in Appendix F. All hyperparameter tuning was performed on the designated validation set for each dataset. The GRPO mechanism with scaled loss updates (as described in Section 4) was used for adaptively balancing $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$.
Appendix F Hyperparameter Details
F.1 Hyperparameters Tuning
For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019]. The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$ and the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$. A batch size of 32 was used. The ViT model used was vit-base-patch16-224555https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k. All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP (average of $\mathcal{P}$ and $\bar{\mathcal{P}}$ mAP), with a patience of 5 epochs  [^6] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
Specific hyperparameters for FairPO variants were tuned via grid search on the validation set:
  • FairPO-DPO: $\beta$ (DPO strength parameter, see Eq. 8, 9) was chosen from $\{0.1,0.5,1.0\}$. The non-privileged slack $\epsilon$ (Eq. 17) was chosen from $\{0.01,0.05,0.1\}$.
  • FairPO-SimPO: $\beta$ (SimPO preference term scaling, see Eq. 12) was chosen from $\{0.1,0.5,1.0\}$. The margin $\gamma$ (SimPO, Eq. 12) was chosen from $\{0.05,0.1,0.2\}$. The non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
  • FairPO-CPO: $\beta$ (CPO preference term scaling, see Eq. 13) was chosen from $\{0.1,0.5,1.0\}$. The NLL regularizer weight $\lambda_{\text{CPO}}$ (Eq. 14) was chosen from $\{0.1,0.5,1.0\}$. The non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE. Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$. Base learning rate $\eta_{\mathbf{w}}$ was $41e-$ 4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
Dataset Method $\beta$ / $\lambda_{\text{CPO}}$ $\gamma$ (SimPO only) $\epsilon$
MS-COCO FairPO-DPO $\beta=0.5$ N/A $0.05$
FairPO-SimPO $\beta=0.5$ $0.1$ $0.05$
FairPO-CPO $\beta=0.5,\lambda_{\text{CPO}}=0.5$ N/A $0.01$
NUS-WIDE FairPO-DPO $\beta=0.1$ N/A $0.05$
FairPO-SimPO $\beta=0.1$ $0.05$ $0.1$
FairPO-CPO $\beta=0.1,\lambda_{\text{CPO}}=0.5$ N/A $0.05$
For the baseline BCE-SFT + Privileged Re-Weighting, the loss weight for privileged labels was set to 5 after tuning from $\{2,3,5,10\}$. For GDRO + BCE, the $\eta_{\alpha}$ was tuned similarly to FairPO.
F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
We investigated the sensitivity of FairPO-CPO’s performance on MS-COCO to variations in the preference strength parameter $\beta$ (from Eq. 13) and the non-privileged slack $\epsilon$ (from Eq. 17).
Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values. Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results. Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed. A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.
Table 6: Sensitivity of FairPO-CPO mAP on MS-COCO to preference strength $\beta$ (with $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$).
$\beta$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
0.05 88.95 90.28
0.1 89.32 90.31
0.5 (Optimal) 89.76 90.34
1.0 89.51 90.25
2.0 88.67 90.11
Table 7: Sensitivity of FairPO-CPO mAP on MS-COCO to non-privileged slack $\epsilon$ (with $\beta=0.5,\lambda_{\text{CPO}}=0.5$).
$\epsilon$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
0.001 89.65 90.40
0.01 (Optimal) 89.76 90.34
0.05 89.81 90.22
0.1 89.70 90.05
0.2 89.55 89.87
Appendix G Computational Resources and Training Time
All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM).
  • Training Time:
    • BCE-SFT (Reference Model Training): Approximately 12-13 hours for MS-COCO and 14-15 hours for NUS-WIDE on a single GPU.
    • FairPO Variants (DPO, SimPO, CPO): Fine-tuning with FairPO typically took slightly longer per epoch than BCE-SFT due to the additional computations for identifying confusing sets and the GRPO updates. FairPO-DPO requires an additional forward pass for the reference model if reference logits are not pre-computed and stored. For FairPO-SimPO and FairPO-CPO, this is not needed. Total training time for FairPO variants was roughly 13-14 hours for MS-COCO and 15-17 hours for NUS-WIDE on a single GPU, including early stopping.
  • Total Compute: The entire research project, including hyperparameter tuning for all baselines and FairPO variants across both datasets, involved multiple runs. A rough estimate for the total GPU-hours would be in the range of 400-500 GPU-hours.
The memory footprint for training was manageable within a single 80GB A100 GPU.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the FairPO framework, its methodology for addressing fairness in MLC, and the key findings regarding improved privileged label performance while maintaining non-privileged performance, which are supported by the experimental results in Section 6.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The Discussion section (Section 8) explicitly addresses technical limitations of the FairPO framework, including aspects like the dynamic confusing set, reference model dependency, and GRPO dynamics.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper primarily presents an empirical framework and experimental results; it does not introduce new theoretical results requiring formal proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 5 details the datasets, model architecture, baselines, and evaluation metrics. Appendix F and C provide further specifics on hyperparameters and data processing. Algorithm 1 and 2 describe the method.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper uses publicly available datasets (MS-COCO, NUS-WIDE) as detailed in Section 5 and Appendix C. An anonymized code repository link is provided in the abstract footnote, with plans for public release upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental settings are described in Section 5. Further details on hyperparameters, their selection via validation, optimizer choice, and data splits are provided in Appendix F and C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The current version of the paper reports mean performance metrics from single runs after hyperparameter tuning. Error bars or statistical significance tests across multiple runs were not included due to computational resource constraints for comprehensive multi-run evaluations of all variants and baselines.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix G provides details on the GPU type, software versions, estimated training times per dataset for baselines and FairPO variants, and an overall estimate of total compute for the project.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research aims to improve fairness in machine learning models, uses public datasets, and does not involve human subjects in a way that would contravene the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The core motivation of FairPO is to address fairness disparities in MLC, which is a positive societal impact. While not explicitly detailing negative impacts, the work contributes to responsible AI by design. The limitations section (Section 8) also implicitly touches on areas where the method might not generalize perfectly, which could have downstream effects if misapplied.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper focuses on a training framework for multi-label classification models using standard public image datasets and a pretrained ViT. It does not introduce new large-scale models or datasets deemed high-risk for misuse requiring special safeguards beyond standard responsible research practices.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The datasets MS-COCO (Section 5, [Lin et al., 2015]) and NUS-WIDE (Section 5, [Chua et al., 2009]) are standard public benchmarks and are cited appropriately. The ViT model [Dosovitskiy et al., 2021] is also cited. These assets are used in accordance with their publicly available terms.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper introduces a new methodological framework (FairPO) and provides its algorithm. It does not introduce new datasets or standalone models as primary assets for release. The code for the framework will be released.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This research does not involve crowdsourcing or direct experiments with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research does not involve direct experiments with human subjects requiring IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used as an important, original, or non-standard component of the core methodology presented in this research. The base model is a Vision Transformer (ViT). However, as mentioned in Appendix A, the future work will be tested on the LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
 1 Introduction
-Multi-label classification (MLC), the task of assigning a relevant subset of labels from a universe $\mathcal{T}$ to an instance $x$, is pervasive in modern machine learning, from image annotation (Wei et al., 2016) to document categorization (Kiritchenko and Matwin, 2006) and bioinformatics (Schietgat et al., 2010).
-Standard MLC models, parameterized by $\mathbf{w}$, are often trained by minimizing an aggregate loss, such as the sum of per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2013; Sorower, 2010; Tsoumakas and Katakis, 2007).
-However, such global optimization can obscure or even exacerbate significant fairness issues (Mehrabi et al., 2021; Chouldechova, 2017).
-Real-world label sets frequently include categories with imbalanced frequencies, differing societal importance, or varying inherent difficulty levels, which can result in models with disparate performance across labels (Zhang and Wu, 2021).
-For instance, a system might excel on common labels while consistently failing on rare but critical ones.
+Multi-label classification (MLC)—assigning a subset of labels from a universe $\mathcal{T}$ to an instance xx—is a pervasive task in fields like image annotation (Wang et al., 2016) and document categorization (Zangari et al., 2024; Schietgat et al., 2010).
+Typically, models are trained by minimizing per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2014; Sorower, 2010; Tarekegn et al., 2024).
+However, this global optimization often creates fairness issues (Mehrabi et al., 2022a; Chouldechova, 2017), as imbalanced label frequencies and varying importance lead to disparate performance (Mehrabi et al., 2022b).
+For instance, a model might excel on common labels while consistently failing on rare but critical ones.
 This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
-A key challenge lies in the model’s ability to discriminate effectively, especially for "hard" examples.
+Figure 1: FairPO’s methodology
+A key challenge lies in the model’s ability to discriminate effectively, especially for hard examples.
 Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions.
-For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a ‘confusing negative’ that the model mistakenly ranks above the true label.
+For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a confusing negative that the model mistakenly ranks above the true label.
 Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive).
-Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2012) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
+Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2011) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
 To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization.
-Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2023), FairPO recasts parts of the MLC task as learning explicit preferences.
+Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2024), FairPO recasts parts of the MLC task as learning explicit preferences.
 We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
-Figure 1: FairPO’s methodology: Labels are split into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$).
-Preference loss ($\mathcal{L}_{\mathcal{P}}$) targets $\mathcal{P}$ labels; constrained loss ($\mathcal{L}_{\bar{\mathcal{P}}}$) maintains $\bar{\mathcal{P}}$ performance.
-GRPO adaptively balances these for fairer, robust MLC.
 For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective.
-If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied.
+If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied.
 This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$.
@@
 The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$.
-This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting robust fairness across the label partition (see Figure 1).
-Our contributions are: (1) FairPO, a novel framework for fair MLC using preference signals and group robustness; (2) A conditional, preference-based objective for privileged labels targeting both confusing negatives and confusing positives, with a fallback to standard classification signals; (3) A robust optimization strategy to balance performance across privileged and non-privileged groups.
-Our comprehensive experiments, evaluating DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024), and CPO (Xu et al., 2024) inspired preference objectives within FairPO, demonstrate its efficacy.
-Notably, FairPO, particularly the CPO-variant, achieves significant improvements in mAP and Exact Match Ratio for privileged labels—for instance, gaining up to 3.44% mAP on COCO (Lin et al., 2015) for the least frequent 20% of labels—while robustly preserving the performance metrics on non-privileged labels over the standard baseline.
-This underscores FairPO’s ability to effectively address challenging fairness-performance trade-offs in MLC.
-In Appendix A, we also briefly touch upon the framework’s applicability to fair multi-label attribute generation.
-2 Related Work
-Fairness in Multi-Label Classification: Ensuring fairness in MLC (Zhang and Wu, 2021; Tantithamthavorn et al., 2018) is complex due to multi-faceted label relationships.
-Recent efforts address MLC-specific fairness challenges, such as tackling label imbalance impacting tail labels (Wu et al., 2023), learning instance and class-level subjective fairness (Gao et al., 2023), or incorporating fairness in dynamic learning settings like class-incremental MLC (Masud et al., 2023).
-FairPO contributes by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets, applying distinct fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and managing them via a robustness framework.
-This targeted approach to enhancing performance for pre-defined critical labels, while safeguarding others, differentiates our work.
-Preference Optimization: Preference optimization, especially for aligning LLMs (Ouyang et al., 2022; Christiano et al., 2017), has rapidly advanced.
-Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its reference-free variants like CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies directly from preference pairs.
-The field continues to evolve with methods like Identity Preference Optimization (IPO) for stability (Azhar et al., 2023), Kahneman-Tversky-based optimization (KTO) (Ethayarajh et al., 2024), simple yet effective Rejection Sampling Fine-Tuning (RFT) (Dong et al., 2023), and ongoing theoretical analyses of these methods (Zhao et al., 2024).
-We use preferences not to rank entire outputs, but to specifically differentiate true label scores from those of their confusing positive/negative counterparts within the privileged label set, thereby sharpening the model’s decision boundaries for critical distinctions.
-Group Robust Optimization: Distributionally Robust Optimization (DRO) (Ben-Tal et al., 2009), particularly Group DRO (Sagawa et al., 2019), aims to improve worst-case performance across predefined data groups, enhancing fairness and robustness.
-This concept sees continued development, for instance, in improving its practicality (Zhang et al., 2022).
-Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024) extended this to LLM preference learning, balancing performance across preference groups.
-FairPO directly employs GRPO’s adaptive optimization strategy.
-However, our groups are defined by the label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$), and GRPO balances their distinct, custom-formulated loss objectives.
-This provides a principled mechanism for robustly managing the specific fairness-performance trade-offs in our MLC context.
+This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting fairness across the labels (see Figure 1).
+Our primary contributions are a novel framework for fair MLC using preference signals, a conditional objective targeting hard examples, and a robust optimization strategy to manage fairness trade-offs.
+Our comprehensive experiments show that FairPO, particularly the CPO-variant, achieves significant gains—for instance, up to 3.44% mAP on the least frequent labels of COCO (Lin et al., 2015)—while robustly maintaining performance on non-privileged labels.
+We also outline the framework’s applicability to attribute generation in Appendix A.
 2 Preliminaries: Preference Optimization Methods
 Our framework builds upon recent preference optimization techniques. We briefly review the key formulations.
-Direct Preference Optimization (DPO): DPO (Rafailov et al., 2023) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$.
+Direct Preference Optimization (DPO): DPO (Rafailov et al., 2024) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$.
 Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
@@
 3.1 Problem Setup and Fairness Goals
-Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ is an input instance and $y_{i}=[y_{i1},\dots,y_{iT}]$ is the ground truth binary vector for $T$ labels, we aim to learn a model that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$, using parameters $\mathbf{w}_{t}$.
-These parameters are fine-tuned from a pretrained policy $\pi_{\theta}$, and we leverage a reference model with parameters $\hat{\mathbf{w}}_{t}$ producing scores $m(x_{i};\hat{\mathbf{w}}_{t})$.
-In practice, a pretrained model $\pi_{\theta}$ (e.g., an LLM) yields an input representation $\mathbf{z}_{i}=\pi_{\theta}(x_{i})\in\mathbb{R}^{d}$.
-A classifier head (e.g., FFN) with parameters $\mathbf{w}_{t}$ then computes the score $m(x_{i};\mathbf{w}_{t})$ for label $t$.
-The reference parameters $\hat{\mathbf{w}}$ are typically obtained from standard supervised fine-tuning (SFT) of $\pi_{\theta}$ on $\mathcal{D}$.
-Our fairness goals are as follows: For any label $l\in\mathcal{P}$ (Privileged Group), the model should accurately discriminate its true state ($y_{il}=+1$ or $y_{il}=0$) from confusing alternatives.
-For any label $j\in\bar{\mathcal{P}}$ (Non-Privileged Group), the classification performance under $\mathbf{w}_{j}$ should remain close to that of $\hat{\mathbf{w}}_{j}$.
-To formalize the challenges for privileged labels, we define two types of confusing set for an instance $x_{i}$ and a privileged label $l\in\mathcal{P}$: If $y_{il}=+1$ (true positive), $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ is the set of confusing negatives.
-If $y_{il}=0$ (true negative), $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ is the set of confusing positives.
-The overall confusing set for $(x_{i},l)$ is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
+Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune parameters $\mathbf{w}_{t}$ for a per-label classifier that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$.
+Our framework leverages a reference model with parameters $\hat{\mathbf{w}}_{t}$, typically obtained from standard supervised fine-tuning (SFT).
+Our framework addresses the problem of group fairness, where the goal is to ensure equitable performance across different predefined groups.
+In our MLC context, we define these groups not by instance attributes (e.g., demographics), but by partitioning the labels themselves into a privileged set $\mathcal{P}$ (e.g., rare, critical, or historically underperforming labels) and a non-privileged set $\bar{\mathcal{P}}$.
+Our fairness objective is a nuanced form of equitable treatment: we aim to significantly improve performance for the underperforming privileged group while ensuring that performance for the non-privileged group is not substantially harmed.
+This goal of targeted improvement without undue harm to other groups is enforced by our robust optimization framework.
+For any privileged label $l\in\mathcal{P}$, the model must accurately discriminate its true state from confusing alternatives.
+For any non-privileged label $j\in\bar{\mathcal{P}}$, performance should not degrade significantly from the reference model.
+To formalize this, we define two types of confusing sets for a privileged label $l\in\mathcal{P}$: the set of confusing negatives, $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=+1$, and the set of confusing positives, $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=0$.
+The overall confusing set is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
 3.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
-Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2023) preference loss.
+Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2024) preference loss.
 If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$).
@@
 $h_{\mathbf{w}}(x_{i},l,k)=\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right).$ (8)
-The loss contribution is $-\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
+The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
 If $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$ is sampled, we prefer $l$ over $k$ (i.e., the true negative $l$ is preferred over the confusing positive $k$, meaning $m(x_{i};\mathbf{w}_{l})$ should be lower than $m(x_{i};\mathbf{w}_{k})$ relative to the reference).
@@
 $h_{\mathbf{w}}(x_{i},k,l)=\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right).$ (9)
-The loss contribution is $-\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
+The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
 No Confusing Examples ($S_{il}=\emptyset$): We revert to a standard base classification loss (e.g., Binary Cross-Entropy, BCE) for label $l$.
@@
 The overall Privileged Loss $\mathcal{L}_{\mathcal{P}}$ is the expectation over these conditional losses:
-$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t.
+$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t.
 }l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right$.
@@
 where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$.
-The term $\beta\log(\cdot)$ captures the relative preference, while $-\gamma$ enforces a desired separation margin.
+The term $\beta\log(\cdot)$ captures the relative preference, while $\gamma$ enforces a desired separation margin.
 This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq.
@@
 Thus, the overall CPO-inspired privileged loss, $\mathcal{L}_{\mathcal{P}}^{\text{CPO}}$, could be formulated by combining this preference term (when $S_{il}\neq\emptyset$) with a consistent NLL/BCE component for all true ground truth labels within $\mathcal{P}$:
-$\mathcal{L}_{\mathcal{P}}^{\text{CPO}}(\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},l)\text{ s.t.
-}l\in\mathcal{P}}\Big{[}$ $\mathbf{1}_{S_{il}\neq\emptyset}\cdot L_{\text{pref}}^{\text{CPO}}(x_{i},l,k_{s})+\lambda_{\text{CPO}}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\Big{]},$ (14)
+[…]
 where $k_{s}$ is a confusing example sampled from $S_{il}$, and $\lambda_{\text{CPO}}$ is a weighting hyperparameter.
@@
 The Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$ employs a hinge mechanism, penalizing the model only if its loss $\ell(\mathbf{w}_{j})$ exceeds the reference model’s loss $\ell(\hat{\mathbf{w}}_{j})$ by more than a predefined slack margin $\epsilon\geq 0$:
-$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|\exists t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t.
+$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t.
 }j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
@@
 The FairPO objective becomes the GRPO minimax problem:
-[…]
+$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\,\geq\,0,\,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\,=\,1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})\right].$ (18)
 Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss.
-The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}$ that perform well even under this worst-case weighting.
+The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$ that perform well even under this worst-case weighting.
 This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.
@@
 Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version.
-Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted ¯$\bar{\mathcal{L}}_{g}^{\text{avg}}$.
+Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted $\bar{\mathcal{L}}_{g}^{\text{avg}}$.
 The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability.
@@
 Algorithm 1 FairPO Training Overview (DPO-inspired)
-1:Initialize: Model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(0)}$ (e.g., from supervised fine-tuning), set group weights $\alpha_{\mathcal{P}}^{(0)},\alpha_{\bar{\mathcal{P}}}^{(0)}$
-2:Input: Dataset $\mathcal{D}$, reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$, hyperparameters $\beta,\epsilon,\eta_{\mathbf{w}},\eta_{\alpha}$.
+1:Initialize: Model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(0)}$ (e.g., from supervised fine-tuning), set group weights $\alpha_{\mathcal{P}}^{(0)},\alpha_{\bar{\mathcal{P}}}^{(0)}$
+2:Input: Dataset $\mathcal{D}$, reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, hyperparameters $\beta,\epsilon,\eta_{\mathbf{w}},\eta_{\alpha}$.
 3:For each training iteration $s=0,\dots,S-1$:
@@
 9:    Update group weights $\alpha^{(s+1)}$ via mirror ascent using $\mathcal{L}_{\mathcal{P}}^{(s)},\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ (GRPO step).
-10:      Update model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(s+1)}$ via mirror descent using weighted gradients.
+10:    Update model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(s+1)}$ via mirror descent using weighted gradients.
 11:End For
-12:Return Optimized parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}^{(S)}$.
+12:Return Optimized parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(S)}$.
 13:Full details are in Algorithm 2 (see Appendix B).
 4 Experimental Setup
-Datasets and Preprocessing: We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014222MS-COCO: https://cocodataset.org/ (80 object categories) (Lin et al., 2015) and NUS-WIDE [^3] NUS-WIDE: https://github.com/NExTplusplus/NUS-WIDE/ (81 concept labels) (Chua et al., 2009).
-For partitioning, we designate the 20% least frequent labels in the training set of each dataset as the privileged group ($\mathcal{P}$), with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$).
-We use standard train/validation/test splits for both datasets (see Appendix C for more details).
-Model and Implementation Details: Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021) pretrained on ImageNet [^4] ViT Model: https://huggingface.co/docs/transformers/en/model_doc/vit.
-The FairPO framework, including the classifier head and label-specific parameters $\mathbf{w}_{t}$, is fine-tuned on top of the frozen ViT features.
-The reference parameters $\hat{\mathbf{w}}$ for FairPO are obtained by standard supervised fine-tuning (SFT) of the same ViT architecture with a BCE loss on the target MLC dataset.
-Hyperparameters like learning rates ($\eta_{\mathbf{w}},\eta_{\alpha}$), DPO’s $\beta$, SimPO’s $\gamma$, CPO’s $\lambda_{\text{CPO}}$, and the non-privileged slack $\epsilon$ are tuned via a grid search on the validation set (see Appendix F for more details).
-Evaluation Metrics: Model performance is assessed using established multi-label metrics, calculated separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ label sets.
-These include Mean Average Precision (mAP), which measures the average precision across all labels; Sample F1 Score (F1-Sample), the harmonic mean of precision and recall averaged per sample; Exact Match Ratio (EMR), the proportion of samples with perfectly predicted label sets; and Multi-label Accuracy (Accuracy), representing element-wise prediction correctness averaged over all labels and instances.
-Baselines: FairPO is compared against: (1) BCE-SFT, standard per-label Binary Cross-Entropy (BCE) fine-tuning, which also provides reference parameters $\hat{\mathbf{w}}$.
-(2) BCE-SFT + Privileged Re-Weighting (RW), which statically increases loss weights for the privileged group $\mathcal{P}$.
-(3) Group DRO + BCE (Sagawa et al., 2019), applying Group DRO to minimize the worst-group BCE loss between $\mathcal{P}$ and $\bar{\mathcal{P}}$ (see Appendix D for more details).
-FairPO Variants: We evaluate three main variants of our FairPO framework, each differing in the specific preference loss formulation used for the privileged group $\mathcal{P}$ (as detailed in Section 4.2): FairPO-DPO utilizes the DPO-inspired conditional preference loss (Eq.
-11).
-FairPO-SimPO employs a SimPO-inspired (Meng et al., 2024) reference-free loss, adapting the preference term from Eq.
-12 within the conditional structure of Eq.
-11.
-Lastly, FairPO-CPO uses a CPO-inspired (Xu et al., 2024) reference-free loss combined with an NLL-like regularization component, as formulated in Eq.
-14 (see Appendix E for more details).
+We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014 (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009).
+For our fairness setup, we define the privileged group ($\mathcal{P}$) as the 20% least frequent labels in each training set, with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$).
+Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021), where we fine-tune label-specific classifier heads on top of its features.
+The reference model parameters $\hat{\mathbf{w}}$ are obtained from a standard supervised fine-tuning (SFT) of this architecture with a Binary Cross-Entropy (BCE) loss.
+Performance is assessed separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ sets using standard metrics: Mean Average Precision (mAP), Sample F1, Exact Match Ratio (EMR), and Accuracy.
+We compare FairPO against three baselines: (1) BCE-SFT, which also serves as our reference model; (2) BCE-SFT + Re-Weighting (RW), a simple loss up-weighting for $\mathcal{P}$ labels; and (3) Group DRO + BCE (Sagawa et al., 2020), which minimizes the worst-group loss.
+We test three variants of our framework, differing in their preference loss for privileged labels: FairPO-DPO (Eq.
+11), FairPO-SimPO (Eq.
+12), and FairPO-CPO (Eq.
+14).
+Comprehensive details on datasets, preprocessing, model architecture, baseline implementations, and hyperparameter tuning are deferred to Appendices C, D, and E.
 5 Results and Analysis
-Overview and Performance Summary: We evaluate FairPO against baselines on MS-COCO (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009).
-Tables 1 and 2 detail performance for privileged ($\mathcal{P}$, 20% least frequent) and non-privileged ($\bar{\mathcal{P}}$, 80% most frequent) label sets.
-The $\Delta\text{mAP}(\mathcal{P})$ column highlights mAP gains for $\mathcal{P}$ labels over the BCE SFT baseline.
-A consistent trend emerges: FairPO variants, particularly FairPO-CPO, significantly improve $\mathcal{P}$ set performance while robustly maintaining $\bar{\mathcal{P}}$ set metrics.
-This addresses the common trade-off where gains in one group detrimentally affect the other, a challenge often faced by simpler baselines.
 Table 1: Performance comparison on MS-COCO.
@@
 Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 63.53 70.24 48.12 55.83 91.51 95.22 19.32 11.56 Ref BCE SFT + RW 65.12 69.14 49.51 54.73 92.33 94.81 21.23 10.32 +1.59 GDRO + BCE 64.84 69.91 49.13 55.62 92.11 95.13 21.02 11.34 +1.31 FairPO-DPO 66.34 69.05 51.71 54.52 93.92 95.04 27.91 11.21 +2.81 FairPO-SimPO 64.11 68.03 48.82 53.81 91.94 94.52 20.18 10.19 +0.58 FairPO-CPO 67.12 69.83 52.21 55.24 94.31 95.12 31.87 11.25 +3.59
-Comparative Analysis of FairPO Variants: The FairPO variants differ in their preference mechanisms: FairPO-DPO utilizes a reference model $\hat{\mathbf{w}}$ in its log-ratio preference term.
-This relative formulation can stabilize training by anchoring scores to a pre-trained baseline, making the preference signal robust to the absolute scale of model outputs.
-This likely contributes to its strong EMR-$\mathcal{P}$ gains, indicating better resolution of complex multi-label confusion.
-FairPO-CPO, often the top performer, employs a reference-free contrastive term $\log(m(x_{i};\text{preferred})/m(x_{i};\text{dispreferred}))$.
-Its strength is amplified by an NLL-like regularizer (realized as $\ell_{\text{BCE}}$ for all privileged labels in Eq.
-14).
-This dual focus on relative preference and absolute correctness for $\mathcal{P}$ underpins its superior $\mathcal{P}$ metrics while maintaining robust $\bar{\mathcal{P}}$ performance.
-FairPO-SimPO, another reference-free method using a fixed margin $\gamma$, generally shows improvements but lags behind DPO and CPO variants.
-This suggests that a fixed margin without DPO’s reference anchoring or CPO’s explicit NLL-like regularization might be less robust or harder to tune optimally for diverse preference pairs in MLC.
-Impact on Exact Match Ratio: The substantial improvements in EMR-$\mathcal{P}$ by FairPO-DPO and FairPO-CPO are particularly significant.
-EMR, requiring perfect prediction for the $\mathcal{P}$ label subset, is highly sensitive to confusion.
-FairPO’s preference optimization, by directly addressing ambiguities between true labels and their specific confusing counterparts, enhances discriminative power for these challenging cases.
-This leads to more accurate complete label set predictions for the privileged group.
-The GRPO mechanism ensures these targeted gains do not disproportionately penalize the $\bar{\mathcal{P}}$ group, demonstrating FairPO’s capacity to learn a fairer and more robust multi-label classifier.
+Tables 1 and 2 detail our findings.
+The standard BCE-SFT baseline confirms the fairness problem central to our motivation: on MS-COCO, a significant performance gap exists between the privileged ($\mathcal{P}$) labels (86.32% mAP) and the non-privileged ($\bar{\mathcal{P}}$) labels (90.65% mAP).
+This disparity highlights how standard aggregate losses can fail to ensure equitable performance for rare or difficult categories.
+Our results demonstrate that FairPO variants, particularly FairPO-CPO, effectively address this challenge by navigating the fairness-performance trade-off in a controlled manner.
+On MS-COCO, FairPO-CPO achieves a substantial +3.44% mAP gain for $\mathcal{P}$ labels, while performance on $\bar{\mathcal{P}}$ labels dips by a negligible 0.31%.
+Similarly, on NUS-WIDE, it yields a +3.59% mAP gain for a minor 0.41% drop.
+This is not an uncontrolled side effect but a managed outcome, enforced by the GRPO mechanism and the constrained loss (Eq.
+17), which together prevent significant harm to the non-privileged group while targeting improvements for the privileged one.
+By directly optimizing the model to rank true labels higher than their specific, dynamically identified confusing counterparts, FairPO sharpens discriminative power for the most challenging cases.
+Among the variants, the reference-free FairPO-CPO proves most effective.
+This directly addresses a potential limitation of DPO-style methods, which can be sensitive to the quality of the reference model.
+FairPO-CPO’s strong performance, likely due to its dual objective of optimizing both relative preference and absolute correctness (via its NLL-like regularizer), demonstrates its robustness and versatility, making it a more practical choice than the reference-dependent FairPO-DPO or the fixed-margin FairPO-SimPO.
 6 Ablation Studies
-We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions. Results are in Table 3 and Table 4.
 Table 3: Ablation on core components of FairPO-CPO (MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT. Parentheses show change vs Full FairPO-CPO.
-Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08)
-Table 3 shows that removing the preference loss for $\mathcal{P}$ (using BCE instead) significantly reduces $\mathcal{P}$ gains, underscoring its criticality.
-Removing the $\bar{\mathcal{P}}$ constraint causes a notable drop in $\bar{\mathcal{P}}$ performance, affirming its protective role.
-Using fixed weights instead of GRPO degrades performance for both groups, highlighting GRPO’s superior adaptive balancing.
+Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 55.20 37.15 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (-0.47) (-1.08) (+0.05) (-10.23) (+1.92)
 Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT (86.32). Parentheses show change vs Full FairPO-CPO.
@@
 Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 22.54 35.10 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) (-42.89) (-0.13) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 60.75 34.90 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07) (-4.68) (-0.33)
-Table 4 details the impact of specific choices within the $\mathcal{P}$ loss formulation.
-Restricting preference learning to only Confusing Negatives (when $y_{il}=+1$) severely curtails its effectiveness.
-Given that positive instances of privileged labels can be extremely rare, this makes the preference signal exceptionally sparse.
-This approach neglects to train the model on ranking true negatives ($y_{il}=0$) against confusing positives—a scenario that forms the vast majority of instances for these labels.
-This omission of crucial learning signals for most situations involving $\mathcal{P}$ labels results in a profound degradation in performance for this set.
-Omitting the BCE Fallback for non-confusing $\mathcal{P}$ instances (i.e., no loss if $S_{il}=\emptyset$) also clearly degrades performance.
-This standard classification signal for easier cases is vital for model stability and reinforcing correct predictions within $\mathcal{P}$ when direct preference learning is not active.
-Both aspects of our full preference formulation are thus shown to be essential.
-This demonstrates that preference loss alone is insufficient for stable training, and standard losses play a critical complementary role even under targeted preference optimization.
+We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions (Tables 3 and 4).
+Our analysis of core components (Table 3) confirms the criticality of each part.
+Replacing the preference loss with BCE reduces privileged gains, removing the $\bar{\mathcal{P}}$ constraint harms non-privileged performance, and using fixed weights instead of GRPO degrades both groups.
+To further isolate the value of our fairness framework, we tested a Global CPO variant that applies the preference objective to all labels without the privileged/non-privileged split or GRPO.
+While this proves the preference objective is a strong general-purpose loss (achieving a +2.23% mAP gain over BCE), the full FairPO framework is decisively superior for the targeted fairness task.
+By applying CPO globally, the performance gains are unfocused, resulting in a 1.21 point lower mAP on the privileged set.
+This demonstrates that the full FairPO architecture is crucial for focusing the model’s capacity to resolve the hardest discrimination challenges within the privileged group.
+Furthermore, the design of the preference objective itself is crucial (Table 4).
+Restricting it to only Confusing Negatives (for $y_{il}=+1$) causes a profound performance drop, as this neglects the far more common scenario of ranking true negatives below confusing positives for rare labels, making the learning signal exceptionally sparse.
+Similarly, removing the BCE Fallback for non-confusing instances degrades performance, proving that a standard classification signal on easier cases is vital for model stability.
+Both aspects of our full formulation are thus essential, demonstrating that while targeted preference optimization is powerful, it must be complemented by standard losses and robust balancing for stable, effective training.
+7 Related Work
+Recent efforts in fair MLC address complex challenges like label imbalance impacting tail labels (Guo and Wang, 2021), subjective fairness (Liu et al., 2023), and class-incremental learning (Dong et al., 2025).
+FairPO contributes a novel approach by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets.
+It applies distinct, fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and manages the trade-off with a robustness framework, differentiating our targeted approach from prior work.
+We adapt recent advances in preference optimization, originally developed for aligning LLMs (Ouyang and Others, 2022; Christiano et al., 2023).
+Techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and its reference-free variants CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies from preference pairs.
+Rather than ranking entire outputs, we repurpose these methods to specifically differentiate true label scores from their dynamically identified confusing counterparts within the privileged set, thereby sharpening critical decision boundaries.
+To balance our objectives, we employ a Group Robust Optimization strategy inspired by Group DRO (Sagawa et al., 2020; Rice et al., 2021) and GRPO (Ramesh et al., 2024).
+While these methods typically balance performance across data or preference groups, FairPO uniquely defines its groups by our label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$).
+It then uses GRPO’s adaptive weighting to balance their distinct, custom-formulated loss objectives, providing a principled mechanism for managing the specific fairness-performance trade-offs in our MLC context.
 8 Discussion
-Conclusion: We introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification.
-Our experiments, particularly with the FairPO-CPO variant, underscore the efficacy of this approach in navigating complex fairness-performance trade-offs, highlighting the value of nuanced preference signals for challenging discrimination tasks.
-Limitations: Despite its promising results, FairPO presents technical limitations.
-The dynamic construction of the ‘confusing set’ $S_{il}$ based on evolving model scores can lead to optimization instability or sparse preference signals, especially for rare/easily-separated privileged labels.
-Furthermore, the efficacy of FairPO-DPO and the non-privileged constraint hinges on a well-calibrated reference model $\hat{\mathbf{w}}$; its deficiencies can distort DPO log-ratios or misguide the constraint.
-Finally, GRPO’s adaptive balancing of heterogeneous loss terms ($\mathcal{L}_{\mathcal{P}}$ combining preference and BCE, $\mathcal{L}_{\bar{\mathcal{P}}}$ with a hinge loss) may require careful tuning of $\eta_{\alpha}$ and loss scaling for stable convergence, the theoretical properties of which remain an open area.
-While these limitations highlight the complexity of fair MLC optimization, they also suggest promising directions for refinement and extension, which we talk about next.
-Future Work: Future work will focus on several key directions.
-A primary goal is to extend FairPO’s principles to multi-label attribute generation (see Appendix A for a conceptual outline), aiming to foster fair and diverse outputs.
-We also plan comprehensive empirical validation across a wider array of datasets and modalities, including text, to assess generalizability.
-Further research will involve in-depth analysis of different label partitioning strategies beyond frequency, and pursuing theoretical insights into convergence properties for FairPO.
-Acknowledgments and Disclosure of Funding
-The authors would like to express their gratitude to the members of the CSE Department and CMInDS at IIT Bombay for their support and for providing a conducive research environment.
-We also thank the anonymous reviewers for their valuable feedback which helped improve the quality of this manuscript.
-We acknowledge the use of high-performance computing resources that were instrumental in conducting the experiments presented in this work.
-Funding Disclosure This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors for the submitted work, beyond the general support and infrastructure provided by IIT Bombay to its faculty and students.
-Competing Interests The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification.
+Our experiments, particularly with the FairPO-CPO variant, highlight the value of nuanced preference signals for navigating complex fairness-performance trade-offs in challenging discrimination tasks.
+While promising, FairPO has technical limitations: its dynamic confusing set can lead to instability or sparse signals for rare labels; the DPO-based variant and non-privileged constraint rely on a well-calibrated reference model; and GRPO’s balancing of heterogeneous losses requires careful tuning and lacks theoretical convergence guarantees.
+These challenges also define our future work, which includes extending FairPO’s principles to multi-label attribute generation (Appendix A), conducting comprehensive empirical validation on more datasets and modalities, analyzing alternative label partitioning strategies, and pursuing theoretical insights into the framework’s convergence properties.
 Appendix A Adapting FairPO for Multi-Attribute Generation
-This section outlines a key direction for future research: adapting the FairPO framework to multi-attribute generation tasks.
-The methodologies described here are conceptual and represent a planned extension of our work, building directly upon the principles established in the main paper for classification.
-Experimental validation of this approach is part of our future agenda.
-In such a generative setting, the goal would be to produce an output sequence $y$ from a prompt $x$ using a single generative policy $\pi_{\mathbf{w}}(y|x)$, such that the generation aligns with fairness goals defined over a set of attributes $\mathcal{A}$.
-As in the classification setting, a potential approach involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets.
-The core GRPO minimax structure (Eq.
-18) could be retained, with group losses defined based on a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$ and the attributes $j_{i}$ driving the preferences.
-A primary consideration in this adaptation would be how preference losses are applied, particularly for the privileged group, as DPO would then operate on the log-probabilities of entire generated sequences rather than individual label scores.
-Proposed Privileged Loss ($\mathcal{L}_{\mathcal{P}}$): For future generative applications, the aim for privileged attributes $j\in\mathcal{P}$ would be to ensure that the learned policy $\pi_{\mathbf{w}}$ strongly reflects observed preferences $y_{w}\succ y_{l}$ when the preference was established based on such an attribute.
-This could translate to assigning significantly higher relative log-probability to the preferred sequence $y_{w}$ compared to the dispreferred sequence $y_{l}$, relative to a reference policy $\pi_{\text{ref}}$.
-This could be achieved by utilizing a standard DPO loss formulation, averaged over the subset of the preference data $\mathcal{D}_{pref}$ where the driving attribute $j$ belongs to the privileged set $\mathcal{P}$.
-The privileged loss could then be defined as:
+This section outlines our planned extension of FairPO to multi-attribute generation, a conceptual direction for future work.
+The goal is to generate a sequence $y$ from a prompt $x$ using a policy $\pi_{\mathbf{w}}(y|x)$ that aligns with fairness goals over a set of attributes $\mathcal{A}$.
+The core idea involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets and retaining the GRPO minimax structure (Eq.
+18).
+The group losses would be defined over a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$, with preference losses like DPO applied to the log-probabilities of entire generated sequences rather than individual label scores.
+Proposed Privileged Loss ($\mathcal{L}_{\mathcal{P}}$): For privileged attributes $j\in\mathcal{P}$, the goal is to ensure the learned policy $\pi_{\mathbf{w}}$ strongly reflects preferences $y_{w}\succ y_{l}$ established by that attribute.
+This is achieved using a standard DPO loss, averaged over the privileged subset of the preference data:
 $\mathcal{L}_{\mathcal{P}}(\pi_{\mathbf{w}},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l},j)\sim\mathcal{D}_{pref}\mid j\in\mathcal{P}}\left[-\log\sigma\left(\beta\cdot h_{\pi_{\mathbf{w}}}(x,y_{w},y_{l})\right)\right]$ (19)
-Minimizing such a loss would directly encourage the model $\pi_{\mathbf{w}}$ to favor preferred sequences over dispreferred ones for preferences driven by privileged attributes, relative to the reference model $\pi_{\text{ref}}$.
-Proposed Non-Privileged Loss ($\mathcal{L}_{\bar{\mathcal{P}}}$): The objective for non-privileged attributes $k\in\bar{\mathcal{P}}$ would remain analogous to the classification setting: to prevent significant degradation compared to a baseline.
-The non-privileged loss could then utilize a hinge formulation based on the DPO loss for preferences driven by attributes $k\in\bar{\mathcal{P}}$:
+Minimizing this loss directly encourages the model to favor preferred sequences for preferences driven by privileged attributes, relative to the reference policy $\pi_{\text{ref}}$.
+Proposed Non-Privileged Loss ($\mathcal{L}_{\bar{\mathcal{P}}}$): For non-privileged attributes $k\in\bar{\mathcal{P}}$, the objective remains analogous to the classification setting: preventing significant performance degradation.
+This is accomplished with a hinge formulation based on the DPO loss:
 $\mathcal{L}_{\bar{\mathcal{P}}}(\pi_{\mathbf{w}},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l},k)\sim\mathcal{D}_{pref}\mid k\in\bar{\mathcal{P}}}\left[\max\left(0,\mathcal{L}_{DPO}(\pi_{\mathbf{w}},\pi_{\text{ref}};x,y_{w},y_{l})-(\log 2)-\epsilon^{\prime}\right)\right].$ (20)
-This would penalize the model only if its preference modeling for non-privileged attributes degrades substantially beyond the baseline performance (represented by $\log 2$ for random preference) plus a slack $\epsilon^{\prime}$.
-The overall FairPO objective for generation would then employ the GRPO minimax formulation to balance $\mathcal{L}_{\mathcal{P}}$ (Eq.
-19) and $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq.
-20), aiming for robust alignment across both privileged and non-privileged attributes identified in the preference data.
-Planned Task and Potential Datasets: Future work will focus on evaluating FairPO for generating text $y$ based on a prompt $x$, aiming for alignment with attributes $\mathcal{A}=\mathcal{P}\cup\bar{\mathcal{P}}$.
-This would require preference datasets $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}$ where the driving attribute $j_{i}$ for the preference is known.
-Potential datasets for such an investigation include:
-  • Publicly available RLHF datasets (e.g., subsets of Anthropic HH-RLHF [Bai et al., 2022], Summarization preferences [Stiennon et al., 2020]), where attributes like ‘Helpfulness’, ‘Harmlessness’, ‘Conciseness’, ‘Faithfulness’ would need to be partitioned into $\mathcal{P}$ and $\bar{\mathcal{P}}$.
-  • Synthetic or newly annotated preference data where attribute labels are explicitly assigned during collection, tailored for this specific fairness investigation.
-Prospective Model and Implementation Strategy: We plan to use pretrained generative language models (e.g., Gemma [Team et al., 2024], Llama [Touvron et al., 2023], Mistral [Jiang et al., 2023]) as the base policy $\pi_{\theta}$.
-These models would then be fine-tuned using the FairPO objective proposed for generation (Equations 19 and 20).
-The reference policy $\pi_{\text{ref}}$ would typically be the SFT version of the base model.
-Proposed Evaluation and Baselines: Evaluation in this future work would primarily rely on preference-based metrics. Comparisons would include:
-  • Win rates (human or model-based, e.g., using GPT-4) for preferences driven by privileged attributes $j\in\mathcal{P}$.
-  • Win rates for preferences driven by non-privileged attributes $k\in\bar{\mathcal{P}}$.
-  • The worst-group win rate across the two attribute groups.
-Baselines for comparison would likely include standard non-robust DPO fine-tuning and the SFT reference model.
-Future experiments within this generative setting would also investigate the impact of hyperparameters like $\beta$ and $\epsilon^{\prime}$, and the potential effectiveness of adapting the alternative SimPO and CPO loss formulations to the FairPO framework for generation.
+This penalizes the model only if its preference modeling for non-privileged attributes degrades substantially beyond baseline performance (represented by $\log 2$ for random preference) plus a slack $\epsilon^{\prime}$.
+The overall FairPO objective would then use GRPO to balance these two losses.
+Planned Experiments: The planned task is generating text from a prompt, requiring preference datasets where the driving attribute $j_{i}$ is known.
+This could involve partitioning attributes like ‘Helpfulness’ or ‘Harmlessness’ from existing RLHF datasets (e.g., Anthropic HH-RLHF [Bai and Others, 2022], Summarization preferences [Stiennon et al., 2022]) or creating new, explicitly annotated data.
+We plan to fine-tune pretrained language models (e.g., Gemma [Team, 2024], Llama [Touvron and Others, 2023], Mistral [Jiang and Others, 2023]) with this objective.
+Evaluation would rely on preference-based metrics, such as win rates for privileged and non-privileged attribute groups, compared against baselines like standard DPO and SFT.
+Future experiments would also explore hyperparameter effects and adapting the SimPO/CPO loss formulations for generation.
 Appendix B FairPO Algorithm
-The FairPO framework is trained iteratively to solve the minimax objective presented in Eq. 18. The detailed procedure, inspired by the DPO-based variant of FairPO, is provided in Algorithm 2.
-The core training loop proceeds as follows:
-  1. 1.
-    Initialization: Model parameters $\{\mathbf{w}_{t}|\exists t\in\mathcal{T}\}$ are initialized (e.g., from a reference model $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$), and the adaptive group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are typically set to uniform values (e.g., 0.5 each).
-  2. 2.
-    Iterative Updates: For each training step:
-    • An instance $(x_{i},y_{i})$ is sampled from the dataset $\mathcal{D}$. A label $r\in\mathcal{T}$ is then randomly selected for processing from this instance.
-    • Privileged Label Processing (if $r\in\mathcal{P}$): The algorithm first identifies if any confusing examples $S_{il}$ exist for the chosen privileged label $l=r$ with respect to other labels (Lines 11-17 in Alg.
-2).
-If $S_{il}\neq\emptyset$, a DPO-inspired preference loss (Eq.
-8 or 9) is computed between label $l$ and a sampled confusing example $k\in S_{il}$ (Lines 19-27).
-This loss encourages the model to correctly rank $l$ against $k$.
-If $S_{il}=\emptyset$ (no confusing examples found), a standard base classification loss (e.g., BCE, Eq.
-10) is used for label $l$ to ensure continued learning (Lines 29-31).
-The resulting loss contributes to $\mathcal{L}_{\mathcal{P}}^{(s)}$.
-    • Non-Privileged Label Processing (if $r\in\bar{\mathcal{P}}$): For a chosen non-privileged label $j=r$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed as per Eq.
-17, penalizing deviations from the reference model’s performance beyond a slack $\epsilon$ (Lines 34-37).
-    • Adaptive Weight Update (GRPO): The group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated using a mirror ascent step (exponential weighting based on the current step’s group losses $\mathcal{L}_{\mathcal{P}}^{(s)}$ and $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ should be scaled for stable training as explained in Section 4) and then normalized to sum to one (Lines 39-41).
-This dynamically adjusts the focus on the group currently performing worse.
-    • Model Parameter Update: The model parameters $\mathbf{w}_{t}$ for all labels $t\in\mathcal{T}$ are updated via a mirror descent step (gradient descent), using a combined gradient weighted by the updated $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ (Line 42).
-  3. 3.
-    The process repeats for a predefined number of iterations or until convergence.
-This iterative procedure allows FairPO to dynamically balance the specialized objectives for privileged labels against the performance constraints for non-privileged labels, guided by the GRPO mechanism to achieve robust fairness.
-For variants like FairPO-SimPO or FairPO-CPO, the DPO-inspired preference loss component (Lines 21-26) for privileged labels is replaced with their respective preference formulations (e.g., Eq.
-12 or 13, possibly combined with an NLL-like term as in Eq.
-14), while the overall GRPO structure and non-privileged handling remain consistent.
+The FairPO framework is trained iteratively to solve the minimax objective presented in Eq.
+18.
+The detailed procedure, which is inspired by the DPO-based variant of FairPO, is provided in Algorithm  2.
+Initialization: The training process begins by initializing the model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$, for instance by copying them from a pre-trained reference model $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$.
+The adaptive group weights, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$, are typically set to uniform values such as 0.5 each.
+Iterative Training Loop: The core of the framework is an iterative training loop.
+In each step, an instance $(x_{i},y_{i})$ is sampled from the dataset $\mathcal{D}$, and a single label $r\in\mathcal{T}$ is randomly selected from that instance for processing.
+The subsequent steps depend on whether this sampled label belongs to the privileged or non-privileged set.
+If the sampled label $r$ is in the privileged set $\mathcal{P}$, the algorithm first identifies if a confusing set $S_{il}$ exists for that label (where $l=r$), as detailed in Algorithm 2.
+The loss computation is then conditional on this set:
+  • If confusing examples exist ($S_{il}\neq\emptyset$), a DPO-inspired preference loss is computed between label $l$ and a randomly sampled confusing example $k\in S_{il}$.
+This preference loss directly encourages the model to improve its ranking of $l$ relative to its specific confounder $k$.
+  • If no confusing examples are found ($S_{il}=\emptyset$), the algorithm reverts to a standard base classification loss (e.g., BCE, Eq.
+10) for label $l$.
+This fallback is crucial as it ensures the model continues to receive a learning signal on easier instances, promoting stable training.
+The loss calculated from either of these cases contributes to the current step’s privileged group loss, $\mathcal{L}_{\mathcal{P}}^{(s)}$.
+Conversely, if the sampled label $r$ belongs to the non-privileged set $\bar{\mathcal{P}}$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed according to Eq.
+17.
+This loss penalizes the model only if its performance on the label $j=r$ deviates from the reference model’s performance by more than a predefined slack margin $\epsilon$.
+After the appropriate group loss is computed, the GRPO mechanism performs two key updates.
+First, the Adaptive Weight Update adjusts the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ using a mirror ascent step.
+This step uses an exponential weighting based on the current (and scaled) group losses, dynamically increasing the focus on the group that is currently performing worse (Lines 39-41).
+Second, the Model Parameter Update updates all model parameters $\mathbf{w}_{t}$ via a mirror descent step, using a combined gradient that is weighted by the newly updated adaptive weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$.
+This entire process repeats for a predefined number of iterations or until convergence, allowing FairPO to dynamically balance its objectives to achieve robust fairness.
+For variants like FairPO-SimPO or FairPO-CPO, the core logic remains identical; only the DPO-inspired preference loss component is replaced with their respective preference formulations (e.g., Eq.
+12 or 13).
+The overall GRPO structure and non-privileged handling are consistent across all variants.
 Algorithm 2 FairPO Algorithm for Multi-Label Classification (DPO-inspired)
-1:Initialize: model parameters $\{\mathbf{w}_{t}^{(0)}\in\mathbb{R}^{d}|\forall t\in\mathcal{T}\}$ (e.g., copy $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$), group weights $\alpha_{\mathcal{P}}^{(0)}\leftarrow 0.5,\alpha_{\bar{\mathcal{P}}}^{(0)}\leftarrow 0.5$.
-2:Choose: learning rates $\eta_{\mathbf{w}},\eta_{\alpha}$, DPO hyperparameter $\beta$, reference parameters $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$, constraint slack $\epsilon$.
+1:Initialize: $\{\mathbf{w}_{t}^{(0)}\in\mathbb{R}^{d}|\forall t\in\mathcal{T}\}$ (e.g., copy $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$), $\alpha_{\mathcal{P}}^{(0)}\leftarrow 0.5,\alpha_{\bar{\mathcal{P}}}^{(0)}\leftarrow 0.5$.
+2:Choose: $\eta_{\mathbf{w}},\eta_{\alpha}$, $\beta$, $\{\hat{\mathbf{w}}_{t}|\forall t\in\mathcal{T}\}$, $\epsilon$.
 3:for $s=0$ to $S$ (MaxIterations) do
@@
 9:    if $r\in\mathcal{P}$ then $\triangleright$ Handle privileged label
-10:          $l\leftarrow r$
-11:          $S_{il}^{\text{neg}}\leftarrow\emptyset$, $S_{il}^{\text{pos}}\leftarrow\emptyset$
-12:          if $y_{il}=+1$ then $\triangleright$ True Positive case for privileged label $l$
-13:              $S_{il}^{\text{neg}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\geq m(x_{i};\mathbf{w}_{l}^{(s)})\}$
-14:              $S_{il}\leftarrow S_{il}^{\text{neg}}$
-15:          else if $y_{il}=0$ then $\triangleright$ True Negative case for privileged label $l$
-16:              $S_{il}^{\text{pos}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\leq m(x_{i};\mathbf{w}_{l}^{(s)})\}$
-17:              $S_{il}\leftarrow S_{il}^{\text{pos}}$
-18:          end if
-19:          if $S_{il}\neq\emptyset$ then $\triangleright$ Confusing examples exist, use DPO-inspired loss
-20:              Sample $k\in S_{il}\sim Uniform(\frac{1}{|S_{il}|})$
-21:              if $y_{il}=+1$ then $\triangleright$ DPO for True Positive $l$ vs Confusing Negative $k$
-22:                  $h_{\mathbf{w}^{(s)}}(x_{i},l,k)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)$.
-23:                  $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},l,k)\right)$
-24:              else if $y_{il}=0$ then $\triangleright$ DPO for True Negative $l$ vs Confusing Positive $k$
-25:                  $h_{\mathbf{w}^{(s)}}(x_{i},k,l)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)$.
-26:                  $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},k,l)\right)$
-27:              end if
-28:              $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{pref}}$
-29:              $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{pref}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
-30:          else $\triangleright$ No confusing examples, use BCE loss for privileged label $l$
-31:              $\mathcal{L}_{\text{BCE}}\leftarrow-[y_{il}\log m(x_{i};\mathbf{w}_{l}^{(s)})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}^{(s)}))]$
-32:              $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{BCE}}$
-33:              $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{BCE}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
-34:          end if
-35:      else if $r\in\bar{\mathcal{P}}$ then $\triangleright$ Handle non-privileged label
-36:          $j\leftarrow r$
-37:          $\ell(\mathbf{w}_{j}^{(s)})\leftarrow-[y_{ij}\log(m(x_{i};\mathbf{w}_{j}^{(s)}))+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}^{(s)}))]$
-38:          $\ell(\hat{\mathbf{w}}_{j})\leftarrow-[y_{ij}\log(m(x_{i};\hat{\mathbf{w}}_{j}))+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))]$
-39:          $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow\max\left(0,\,\ell(\mathbf{w}_{j}^{(s)})-\ell(\hat{\mathbf{w}}_{j})-\epsilon\right)$
-40:          $g_{\bar{\mathcal{P}}}^{t}\leftarrow g_{\bar{\mathcal{P}}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
-41:      end if
-42:      $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\alpha_{\mathcal{P}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\mathcal{P},scaled}^{(s)})$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\alpha_{\bar{\mathcal{P}}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\bar{\mathcal{P}},scaled}^{(s)})$triangleright Mirror ascent
-43:      $Z\leftarrow\alpha_{\mathcal{P}}^{(s+1)}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}$
-44:      $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\frac{\alpha_{\mathcal{P}}^{(s+1)}}{Z}$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\frac{\alpha_{\bar{\mathcal{P}}}^{(s+1)}}{Z}$triangleright Weight normalization
-45:      $\mathbf{w}_{t}^{(s+1)}\leftarrow\mathbf{w}_{t}^{(s)}-\eta_{\mathbf{w}}(\alpha_{\mathcal{P}}^{(s+1)}g_{\mathcal{P}}^{t}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}g_{\bar{\mathcal{P}}}^{t})$$\forall t\in\mathcal{T}$$\triangleright$ Mirror descent
-46:end for
-47:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$
+10:     $l\leftarrow r$, $S_{il}^{\text{neg}}\leftarrow\emptyset$, $S_{il}^{\text{pos}}\leftarrow\emptyset$
+11:     if $y_{il}=+1$ then $\triangleright$ True Positive case for privileged label $l$
+12:      $S_{il}^{\text{neg}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\geq m(x_{i};\mathbf{w}_{l}^{(s)})\}$, $S_{il}\leftarrow S_{il}^{\text{neg}}$
+13:     else if $y_{il}=0$ then $\triangleright$ True Negative case for privileged label $l$
+14:      $S_{il}^{\text{pos}}\leftarrow\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k}^{(s)})\leq m(x_{i};\mathbf{w}_{l}^{(s)})\}$, $S_{il}\leftarrow S_{il}^{\text{pos}}$
+15:     end if
+16:     if $S_{il}\neq\emptyset$ then $\triangleright$ Confusing examples exist, use DPO-inspired loss
+17:      Sample $k\in S_{il}\sim Uniform(\frac{1}{|S_{il}|})$
+18:      if $y_{il}=+1$ then $\triangleright$ DPO for True Positive $l$ vs Confusing Negative $k$
+19:        $h_{\mathbf{w}^{(s)}}(x_{i},l,k)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)$.
+20:        $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},l,k)\right)$
+21:      else if $y_{il}=0$ then $\triangleright$ DPO for True Negative $l$ vs Confusing Positive $k$
+22:        $h_{\mathbf{w}^{(s)}}(x_{i},k,l)\leftarrow\left(\log\frac{m(x_{i};\mathbf{w}_{k}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l}^{(s)})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)$.
+23:        $\mathcal{L}_{\text{pref}}\leftarrow-\log\sigma\left(\beta\cdot h_{\mathbf{w}^{(s)}}(x_{i},k,l)\right)$
+24:      end if
+25:      $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{pref}}$, $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{pref}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
+26:     else $\triangleright$ No confusing examples, use BCE loss for privileged label $l$
+27:      $\mathcal{L}_{\text{BCE}}\leftarrow-[y_{il}\log m(x_{i};\mathbf{w}_{l}^{(s)})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}^{(s)}))]$
+28:      $\mathcal{L}_{\mathcal{P}}^{(s)}\leftarrow\mathcal{L}_{\text{BCE}}$, $g_{\mathcal{P}}^{t}\leftarrow g_{\mathcal{P}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\text{BCE}}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
+29:     end if
+30:    else if $r\in\bar{\mathcal{P}}$ then $\triangleright$ Handle non-privileged label
+31:     $j\leftarrow r$
+32:     $\ell(\mathbf{w}_{j}^{(s)})\leftarrow-[y_{ij}\log(m(x_{i};\mathbf{w}_{j}^{(s)}))+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}^{(s)}))]$
+33:     $\ell(\hat{\mathbf{w}}_{j})\leftarrow-[y_{ij}\log(m(x_{i};\hat{\mathbf{w}}_{j}))+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))]$
+34:     $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}\leftarrow\max\left(0,\,\ell(\mathbf{w}_{j}^{(s)})-\ell(\hat{\mathbf{w}}_{j})-\epsilon\right)$, $g_{\bar{\mathcal{P}}}^{t}\leftarrow g_{\bar{\mathcal{P}}}^{t}+\nabla_{\mathbf{w}_{t}}\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}|_{\mathbf{w}_{t}^{(s)}}\quad\forall t\in\mathcal{T}$.
+35:    end if
+36:    $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\alpha_{\mathcal{P}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\mathcal{P},scaled}^{(s)})$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\alpha_{\bar{\mathcal{P}}}^{(s)}\exp(\eta_{\alpha}\mathcal{L}_{\bar{\mathcal{P}},scaled}^{(s)})$ $\triangleright$ Mirror ascent
+37:    $Z\leftarrow\alpha_{\mathcal{P}}^{(s+1)}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}$, $\alpha_{\mathcal{P}}^{(s+1)}\leftarrow\frac{\alpha_{\mathcal{P}}^{(s+1)}}{Z}$ and $\alpha_{\bar{\mathcal{P}}}^{(s+1)}\leftarrow\frac{\alpha_{\bar{\mathcal{P}}}^{(s+1)}}{Z}$ $\triangleright$ Weight normalization
+38:    $\mathbf{w}_{t}^{(s+1)}\leftarrow\mathbf{w}_{t}^{(s)}-\eta_{\mathbf{w}}(\alpha_{\mathcal{P}}^{(s+1)}g_{\mathcal{P}}^{t}+\alpha_{\bar{\mathcal{P}}}^{(s+1)}g_{\bar{\mathcal{P}}}^{t})$$\forall t\in\mathcal{T}\triangleright$ Mirror descent
+39:end for
+40:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$
 Appendix C Dataset and Preprocessing Details
@@
 The BCE-SFT baseline represents standard supervised fine-tuning.
-  • Objective: The model is trained by minimizing the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels.
-For an instance $(x_{i},y_{i})$, the loss is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$, where $m(x_{i};\mathbf{w}_{t})$ is the model’s predicted probability for label $t$.
-  • Implementation: A separate non-linear classifier head (same MLP as used in FairPO’s main experiments) is trained on top of the frozen ViT features for each label $t$, with parameters $\mathbf{w}_{t}$.
-  • Hyperparameters: The primary hyperparameter tuned was the learning rate for the classifier heads.
-This was selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ based on the best overall mean Average Precision (mAP) on the validation set.
-Weight decay was set consistent with FairPO experiments (e.g., 0.01).
-  • Role: This model serves two purposes:
-    1. 1.
-      It acts as a direct performance baseline.
-    2. 2.
-      The final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used by the FairPO framework (for DPO-based losses and the non-privileged constraint).
-  • Training: Trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
+Its objective is to minimize the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels, where the loss for an instance $(x_{i},y_{i})$ is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$.
+For implementation, a separate non-linear MLP classifier head is trained for each label $t$ on top of frozen ViT features, consistent with the main FairPO architecture.
+The primary hyperparameter tuned was the learning rate for these classifier heads, selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ to achieve the best overall mean Average Precision (mAP) on the validation set.
+Weight decay was set to 0.01, matching the FairPO experiments.
+This model serves two critical roles in our study.
+First, it acts as a direct performance baseline against which we measure fairness improvements.
+Second, the final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used throughout the FairPO framework, both for the DPO-based losses and for the non-privileged loss constraint.
+The model was trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
 D.2 BCE-SFT + Privileged Re-Weighting (RW)
-This baseline modifies the standard BCE-SFT by applying static loss re-weighting.
-  • Objective: The goal is to improve performance on the privileged labels $\mathcal{P}$ by assigning a higher weight to their contribution in the total loss. The loss function is:
+This baseline modifies the standard BCE-SFT by applying static loss re-weighting to improve performance on the privileged labels $\mathcal{P}$.
+The objective is to assign a higher weight to their contribution in the total loss function:
 $\mathcal{L}_{\text{RW}}=\lambda_{\mathcal{P}}\sum_{l\in\mathcal{P}}\text{BCE}(m(x_{i};\mathbf{w}_{l}),y_{il})+\lambda_{\bar{\mathcal{P}}}\sum_{j\in\bar{\mathcal{P}}}\text{BCE}(m(x_{i};\mathbf{w}_{j}),y_{ij})$
-    Typically, $\lambda_{\bar{\mathcal{P}}}$ is set to $1$, and $\lambda_{\mathcal{P}}>1$.
-  • Implementation: Same architecture and per-label classifiers as BCE-SFT. The loss computation is modified to incorporate the weights.
-  • Hyperparameters:
-    • Learning rate: Tuned as for BCE-SFT.
-    • Privileged group weight $\lambda_{\mathcal{P}}$: This is the key hyperparameter.
-It was tuned via grid search on the validation set, selecting from values such as $\{2,3,5,8,10\}$.
-The value was chosen to maximize mAP on the privileged group $\mathcal{P}$ without causing an excessive drop in mAP on the non-privileged group $\bar{\mathcal{P}}$, or based on the best overall mAP.
-For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective.
-$\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$.
-  • Training: Trained similarly to BCE-SFT.
+For this setup, we use the same per-label classifier architecture as BCE-SFT, but the loss computation is modified to incorporate the weights.
+The key hyperparameter is the privileged group weight $\lambda_{\mathcal{P}}$, which was tuned via grid search from values $\{2,3,5,8,10\}$.
+The value was chosen to maximize mAP on the privileged group without an excessive drop in performance on the non-privileged group.
+For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective, while $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$.
+Training was conducted similarly to BCE-SFT.
 D.3 Group DRO + BCE
 This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss.
-  • Objective: To minimize the worst-case BCE loss across the predefined groups (privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$). The objective is:
+The objective is to minimize the worst-case BCE loss across the predefined privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) groups:
 $\min_{\{\mathbf{w}_{t}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\geq 0,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\text{BCE}}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})\right]$
-    where $\mathcal{L}_{\text{BCE}}(\mathcal{G})=\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}\text{BCE}(m(x_{i};\mathbf{w}_{g}),y_{ig})$ is the average BCE loss for labels within group $\mathcal{G}$.
-  • Implementation:
-    • Architecture: Same as BCE-SFT.
-    • Loss Calculation: Per-label BCE losses are computed.
-These are then aggregated (averaged) separately for all labels in $\mathcal{P}$ to get $\mathcal{L}_{\text{BCE}}(\mathcal{P})$ for the current batch/sample, and similarly for $\bar{\mathcal{P}}$ to get $\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})$.
-    • Optimization: Model parameters $\{\mathbf{w}_{t}\}$ are updated using AdamW.
-The group weights $\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step (exponential weighting based on the current group losses, followed by normalization):
-      $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$
-      The final loss for backpropagation through model parameters is $\alpha_{\mathcal{P}}^{(s+1)}\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}^{(s+1)}\mathcal{L}_{\text{BCE}}^{(s)}(\bar{\mathcal{P}})$.
-  • Hyperparameters:
-    • Learning rate for model parameters $\{\mathbf{w}_{t}\}$: Tuned as for BCE-SFT.
-    • Learning rate for group weights $\eta_{\alpha}$: Tuned from $\{0.01,0.05,0.1,0.2\}$ on the validation set, similar to FairPO’s $\eta_{\alpha}$.
-The selection aimed for stable convergence and good worst-group performance.
-For Group DRO with BCE, the raw group losses might be directly used in the exponential update, or a similar loss scaling/normalization approach as discussed for FairPO’s GRPO step could be employed for improved stability if raw loss magnitudes vary significantly between groups or over time.
-For the reported results, $\eta_{\alpha}$ was set to $0.05$ after tuning.
-  • Training: Trained similarly to BCE-SFT and FairPO regarding epochs and batching.
-This section provides further details on the experimental setup for the main FairPO variants (FairPO-DPO, FairPO-SimPO, and FairPO-CPO), complementing the information in Section 5 of the main paper.
+Here, $\mathcal{L}_{\text{BCE}}(\mathcal{G})$ is the average BCE loss for all labels within a group $\mathcal{G}$.
+The implementation uses the same BCE-SFT architecture.
+Per-label BCE losses are computed and then aggregated separately for the $\mathcal{P}$ and $\bar{\mathcal{P}}$ groups.
+Model parameters $\{\mathbf{w}_{t}\}$ are updated with AdamW, while the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step: $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$.
+The final loss for backpropagation is the adaptively weighted sum of the group losses.
+For hyperparameters, the model learning rate was tuned as for BCE-SFT.
+The group weight learning rate $\eta_{\alpha}$ was tuned from $\{0.01,0.05,0.1,0.2\}$, with $\eta_{\alpha}=0.05$ selected for its stable convergence and good worst-group performance.
+For the reported results, raw group losses were used in the exponential update, though we note that a loss scaling approach similar to that in FairPO could also be employed for enhanced stability.
+The training protocol regarding epochs and batching was identical to FairPO.
 Appendix E FairPO Experimental Details
 E.1 Common Setup for All FairPO Variants
-Unless specified otherwise, the following setup details apply to all FairPO variants:
-  • Base Model: A Vision Transformer (ViT) model, specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k, was used as the feature extractor.
-The ViT backbone (i.e., the ‘model.vit‘ component) was kept frozen during fine-tuning with FairPO, except for its final encoder layer, which was made trainable to allow for some adaptation of higher-level features.
-  • Datasets and Preprocessing: Experiments were conducted on MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009].
-Images were resized to $224\times 224$ pixels and normalized using ImageNet statistics.
-Standard data augmentations (random horizontal flips, random resized crops) were applied during training.
-Details on dataset splits and privileged/non-privileged partitioning (20% least frequent labels as privileged) are provided in Section 5 and Appendix C.
-  • Optimizer: The AdamW optimizer [Loshchilov and Hutter, 2019] was used for updating the trainable parameters (classifier heads and the last ViT encoder layer).
-  • Training Schedule: Models were trained for a maximum of 25 epochs with a batch size of 32.
-Early stopping was employed based on the overall mAP (average of mAP on privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets) on the validation set, with a patience of 5 epochs.
-  • Reference Model ($\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$): The reference parameters $\{\hat{\mathbf{w}}_{t}|\exists t\in\mathcal{T}\}$ required by FairPO-DPO and for the non-privileged loss constraint (Eq.
-17) in all variants were obtained from a BCE-SFT model (see Appendix D).
-This SFT model used the same ViT architecture and per-label MLP classifier heads (detailed below) and was trained with a standard BCE loss.
+Unless specified otherwise, a common setup was used for all FairPO variants to ensure fair comparison.
+The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k.
+During our fine-tuning, the ViT backbone was kept frozen, with the exception of its final encoder layer, which was made trainable to allow for adaptation of higher-level features.
+All experiments were conducted on the MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009] datasets.
+Images were resized to $224\times 224$ pixels, normalized using ImageNet statistics, and augmented with standard techniques like random horizontal flips and resized crops.
+The AdamW optimizer [Loshchilov and Hutter, 2019] was used to update all trainable parameters.
+Models were trained for a maximum of 25 epochs with a batch size of 32, and we employed an early stopping strategy with a patience of 5 epochs based on the overall mAP on the validation set.
+The reference model parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, required for FairPO-DPO and the non-privileged loss constraint in all variants, were obtained from a BCE-SFT model detailed in Appendix D.
 E.2 Per-Label Non-Linear MLP Classifier Head
-For each of the $T$ labels in the dataset, a dedicated, independent non-linear Multi-Layer Perceptron (MLP) head was used to predict the probability of that label being positive.
-This allows each label to have its own specialized classifier operating on the shared image features extracted by the ViT.
-  • Architecture: Each MLP head takes the $d$-dimensional output feature vector from the ViT’s [CLS] token (where $d$ is the hidden size of ViT, e.g., 768 for ViT-Base) as input and outputs a single logit for the corresponding label.
-The architecture of each MLP head is as follows:
+For each of the $T$ labels in a dataset, we employed a dedicated and independent non-linear Multi-Layer Perceptron (MLP) head to predict the probability of that label being positive.
+Using separate MLP heads allows for more complex, non-linear decision boundaries tailored to each label’s specific characteristics, which is particularly beneficial for labels with varying difficulty.
+Each MLP head takes the $d$-dimensional feature vector (where $d=768$ for ViT-Base) from the ViT’s [CLS] token as input and outputs a single logit.
+The final probability score $m(x_{i};\mathbf{w}_{t})$ is obtained by applying a sigmoid function to this logit.
+The specific architecture for each MLP head is as follows:
   1. 1.
-      Linear Layer: $d\rightarrow 256$ neurons
+    Linear Layer: $d\rightarrow 256$ neurons, followed by ReLU Activation
   2. 2.
-      ReLU Activation
+    Linear Layer: $256\rightarrow 64$ neurons, followed by ReLU Activation
   3. 3.
-      Linear Layer: $256\rightarrow 64$ neurons
+    Linear Layer: $64\rightarrow 16$ neurons, followed by ReLU Activation
   4. 4.
-      ReLU Activation
+    Linear Layer: $16\rightarrow 4$ neurons, followed by ReLU Activation
   5. 5.
-      Linear Layer: $64\rightarrow 16$ neurons
-    6. 6.
-      ReLU Activation
-    7. 7.
-      Linear Layer: $16\rightarrow 4$ neurons
-    8. 8.
-      ReLU Activation
-    9. 9.
     Linear Layer (Output): $4\rightarrow 1$ neuron (producing the logit)
-    The final probability score $m(x_{i};\mathbf{w}_{t})$ for label $t$ is obtained by applying a sigmoid function to this output logit.
-  • Parameters: The parameters $\mathbf{w}_{t}$ for each label $t$’s MLP head are specific to that label and are learned during the fine-tuning process.
-The set of all such parameters is denoted by $\{\mathbf{w}_{t}|\forall t\in\mathcal{T}\}$.
-  • Trainability: All parameters of these MLP heads were trainable during both SFT (for the reference model) and FairPO fine-tuning.
-  • Motivation: Using separate MLP heads rather than a single shared linear layer allows for more complex, non-linear decision boundaries tailored to each label’s specific characteristics and inter-dependencies with the input features.
-This can be particularly beneficial for labels with varying difficulty or those requiring nuanced feature interpretations.
+The parameters $\mathbf{w}_{t}$ for each label $t$’s MLP head are unique to that label.
+All parameters within these MLP heads were fully trainable during both the SFT pre-training (for the reference model) and the final FairPO fine-tuning.
 E.3 FairPO Variant-Specific Hyperparameters
-The core FairPO learning rates ($\eta_{\mathbf{w}}$ for model parameters, $\eta_{\alpha}$ for GRPO weights) were tuned via grid search.
-For $\eta_{\mathbf{w}}$ (AdamW LR for classifier heads and last ViT layer), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$.
-For $\eta_{\alpha}$, values from $\{0.01,0.05,0.1,0.2\}$ were explored.
-The final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$ after tuning.
-Specific hyperparameters for each FairPO variant’s privileged loss component were tuned as per Table 5 in Appendix F.
-All hyperparameter tuning was performed on the designated validation set for each dataset.
-The GRPO mechanism with scaled loss updates (as described in Section 4) was used for adaptively balancing $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$.
+The core FairPO learning rates, $\eta_{\mathbf{w}}$ for model parameters and $\eta_{\alpha}$ for GRPO weights, were tuned via grid search.
+For $\eta_{\mathbf{w}}$ (the AdamW LR), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$, and for $\eta_{\alpha}$, values were explored from $\{0.01,0.05,0.1,0.2\}$.
+After tuning, the final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$.
+The specific hyperparameters for each FairPO variant’s privileged loss component (e.g., $\beta$, $\gamma$, $\lambda_{\text{CPO}}$) were also tuned on the designated validation set for each dataset, with final values reported in Table 5 (Appendix F).
+Throughout all experiments, the GRPO mechanism with scaled loss updates, as described in Section 3, was used to adaptively balance the privileged and non-privileged loss terms.
 Appendix F Hyperparameter Details
 F.1 Hyperparameters Tuning
-For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019].
-The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$ and the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$.
-A batch size of 32 was used.
-The ViT model used was vit-base-patch16-224555https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k.
-All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP (average of $\mathcal{P}$ and $\bar{\mathcal{P}}$ mAP), with a patience of 5 epochs  [^6] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
-Specific hyperparameters for FairPO variants were tuned via grid search on the validation set:
-  • FairPO-DPO: $\beta$ (DPO strength parameter, see Eq. 8, 9) was chosen from $\{0.1,0.5,1.0\}$. The non-privileged slack $\epsilon$ (Eq. 17) was chosen from $\{0.01,0.05,0.1\}$.
-  • FairPO-SimPO: $\beta$ (SimPO preference term scaling, see Eq.
-12) was chosen from $\{0.1,0.5,1.0\}$.
-The margin $\gamma$ (SimPO, Eq.
-12) was chosen from $\{0.05,0.1,0.2\}$.
-The non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
-  • FairPO-CPO: $\beta$ (CPO preference term scaling, see Eq.
-13) was chosen from $\{0.1,0.5,1.0\}$.
-The NLL regularizer weight $\lambda_{\text{CPO}}$ (Eq.
-14) was chosen from $\{0.1,0.5,1.0\}$.
-The non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
+For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019] with a batch size of 32.
+The base model was a vit-base-patch16-224222https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k.
+The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$, while the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$.
+All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP, using a patience of 5 epochs [^3] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
+Specific hyperparameters for each FairPO variant were tuned via grid search on the validation set.
+For FairPO-DPO, the strength parameter $\beta$ was chosen from $\{0.1,0.5,1.0\}$, and the non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
+For FairPO-SimPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the margin $\gamma$ from $\{0.05,0.1,0.2\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$.
+For FairPO-CPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the NLL regularizer weight $\lambda_{\text{CPO}}$ from $\{0.1,0.5,1.0\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$.
 The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
+For our baselines, the loss weight for privileged labels in BCE-SFT + Privileged Re-Weighting was set to 5 after tuning from $\{2,3,5,10\}$.
+The group weight learning rate $\eta_{\alpha}$ for GDRO + BCE was tuned similarly to FairPO.
+F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
+We investigated the sensitivity of FairPO-CPO’s performance on MS-COCO to variations in the preference strength parameter $\beta$ (from Eq. 13) and the non-privileged slack $\epsilon$ (from Eq. 17).
 Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE.
 Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$.
-Base learning rate $\eta_{\mathbf{w}}$ was $41e-$ 4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
+Base learning rate $\eta_{\mathbf{w}}$ was 1e41e-4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
 Dataset Method $\beta$ / $\lambda_{\text{CPO}}$ $\gamma$ (SimPO only) $\epsilon$
@@
 FairPO-CPO $\beta=0.1,\lambda_{\text{CPO}}=0.5$ N/A $0.05$
-For the baseline BCE-SFT + Privileged Re-Weighting, the loss weight for privileged labels was set to 5 after tuning from $\{2,3,5,10\}$.
-For GDRO + BCE, the $\eta_{\alpha}$ was tuned similarly to FairPO.
-F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
-We investigated the sensitivity of FairPO-CPO’s performance on MS-COCO to variations in the preference strength parameter $\beta$ (from Eq. 13) and the non-privileged slack $\epsilon$ (from Eq. 17).
 Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values.
 Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results.
-Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed.
-A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.
 Table 6: Sensitivity of FairPO-CPO mAP on MS-COCO to preference strength $\beta$ (with $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$).
@@
 2.0 88.67 90.11
+Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed.
+A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.
 Table 7: Sensitivity of FairPO-CPO mAP on MS-COCO to non-privileged slack $\epsilon$ (with $\beta=0.5,\lambda_{\text{CPO}}=0.5$).
@@
 0.2 89.55 89.87
-Appendix G Computational Resources and Training Time
-All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM).
-  • Training Time:
-    • BCE-SFT (Reference Model Training): Approximately 12-13 hours for MS-COCO and 14-15 hours for NUS-WIDE on a single GPU.
-    • FairPO Variants (DPO, SimPO, CPO): Fine-tuning with FairPO typically took slightly longer per epoch than BCE-SFT due to the additional computations for identifying confusing sets and the GRPO updates.
-FairPO-DPO requires an additional forward pass for the reference model if reference logits are not pre-computed and stored.
-For FairPO-SimPO and FairPO-CPO, this is not needed.
-Total training time for FairPO variants was roughly 13-14 hours for MS-COCO and 15-17 hours for NUS-WIDE on a single GPU, including early stopping.
-  • Total Compute: The entire research project, including hyperparameter tuning for all baselines and FairPO variants across both datasets, involved multiple runs.
-A rough estimate for the total GPU-hours would be in the range of 400-500 GPU-hours.
-The memory footprint for training was manageable within a single 80GB A100 GPU.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: The abstract and introduction clearly state the FairPO framework, its methodology for addressing fairness in MLC, and the key findings regarding improved privileged label performance while maintaining non-privileged performance, which are supported by the experimental results in Section 6.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: The Discussion section (Section 8) explicitly addresses technical limitations of the FairPO framework, including aspects like the dynamic confusing set, reference model dependency, and GRPO dynamics.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This paper primarily presents an empirical framework and experimental results; it does not introduce new theoretical results requiring formal proofs.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: Section 5 details the datasets, model architecture, baselines, and evaluation metrics.
-Appendix F and C provide further specifics on hyperparameters and data processing.
-Algorithm 1 and 2 describe the method.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: The paper uses publicly available datasets (MS-COCO, NUS-WIDE) as detailed in Section 5 and Appendix C.
-An anonymized code repository link is provided in the abstract footnote, with plans for public release upon acceptance.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: Experimental settings are described in Section 5.
-Further details on hyperparameters, their selection via validation, optimizer choice, and data splits are provided in Appendix F and C.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34.
-Justification: The current version of the paper reports mean performance metrics from single runs after hyperparameter tuning.
-Error bars or statistical significance tests across multiple runs were not included due to computational resource constraints for comprehensive multi-run evaluations of all variants and baselines.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39.
-Justification: Appendix G provides details on the GPU type, software versions, estimated training times per dataset for baselines and FairPO variants, and an overall estimate of total compute for the project.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research aims to improve fairness in machine learning models, uses public datasets, and does not involve human subjects in a way that would contravene the Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: The core motivation of FairPO is to address fairness disparities in MLC, which is a positive societal impact.
-While not explicitly detailing negative impacts, the work contributes to responsible AI by design.
-The limitations section (Section 8) also implicitly touches on areas where the method might not generalize perfectly, which could have downstream effects if misapplied.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54.
-Justification: The paper focuses on a training framework for multi-label classification models using standard public image datasets and a pretrained ViT.
-It does not introduce new large-scale models or datasets deemed high-risk for misuse requiring special safeguards beyond standard responsible research practices.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59.
-Justification: The datasets MS-COCO (Section 5, [Lin et al., 2015]) and NUS-WIDE (Section 5, [Chua et al., 2009]) are standard public benchmarks and are cited appropriately.
-The ViT model [Dosovitskiy et al., 2021] is also cited.
-These assets are used in accordance with their publicly available terms.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64.
-Justification: The paper introduces a new methodological framework (FairPO) and provides its algorithm.
-It does not introduce new datasets or standalone models as primary assets for release.
-The code for the framework will be released.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: This research does not involve crowdsourcing or direct experiments with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: This research does not involve direct experiments with human subjects requiring IRB approval.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79.
-Justification: LLMs were not used as an important, original, or non-standard component of the core methodology presented in this research.
-The base model is a Vision Transformer (ViT).
-However, as mentioned in Appendix A, the future work will be tested on the LLMs.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v4

@@
-FairPO: Robust Preference Optimization for Fair Multi-Label Learning
+
 Abstract
-Multi-label classification (MLC) often exhibits performance disparities, especially for infrequent or sensitive label categories.
-We propose FairPO, a novel framework that integrates preference-based loss formulations with group-robust optimization to improve fairness in multi-label classification (MLC), particularly targeting underperforming and sensitive label groups.
-FairPO partitions labels into a privileged set, targeted for enhanced performance, and a non-privileged set, where baseline performance is maintained.
-For privileged labels, a preference-based loss, inspired by Direct Preference Optimization (DPO), encourages model scores for true positives to be significantly higher than for their confusing negative counterparts, and scores for true negatives to be significantly lower than for their confusing positive counterparts—addressing hard examples that challenge standard classifiers.
-For non-privileged labels, a constrained objective ensures performance does not degrade substantially below a reference model.
-These group-specific objectives are balanced using a Group Robust Preference Optimization (GRPO) formulation, adaptively mitigating bias.
-Our experiments also include FairPO variants built on recent reference-free preference optimization techniques, namely Contrastive Preference Optimization (CPO) and Simple Preference Optimization (SimPO), which further highlight FairPO’s versatility [^1] Our code is available at GitHub: https://anonymous.4open.science/r/FairPO.
+Multi-label classification (MLC) often suffers from performance disparities across labels.
+We propose FairPO, a framework combining preference-based loss and group-robust optimization to improve fairness by targeting underperforming labels.
+FairPO partitions labels into a privileged set for targeted improvement and a non-privileged set to maintain baseline performance.
+For privileged labels, a DPO-inspired preference loss addresses hard examples by correcting ranking errors between true labels and their confusing counterparts.
+A constrained objective maintains performance for non-privileged labels, while a Group Robust Preference Optimization (GRPO) formulation adaptively balances both objectives to mitigate bias.
+We also demonstrate FairPO’s versatility with reference-free variants using Contrastive (CPO) and Simple (SimPO) Preference Optimization [^1] Our code is available at GitHub: https://github.com/soumenkm/FairPO.
+00footnotetext: Equal contribution
 1 Introduction
-Multi-label classification (MLC)—assigning a subset of labels from a universe $\mathcal{T}$ to an instance xx—is a pervasive task in fields like image annotation (Wang et al., 2016) and document categorization (Zangari et al., 2024; Schietgat et al., 2010).
-Typically, models are trained by minimizing per-label binary cross-entropies (BCE) over a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (Zhang and Zhou, 2014; Sorower, 2010; Tarekegn et al., 2024).
-However, this global optimization often creates fairness issues (Mehrabi et al., 2022a; Chouldechova, 2017), as imbalanced label frequencies and varying importance lead to disparate performance (Mehrabi et al., 2022b).
-For instance, a model might excel on common labels while consistently failing on rare but critical ones.
-This disparity is particularly problematic when certain labels correspond to protected attributes or groups, demanding equitable treatment.
-Figure 1: FairPO’s methodology
-A key challenge lies in the model’s ability to discriminate effectively, especially for hard examples.
-Standard losses like BCE (Ruby and Yendapalli, 2020) may provide insufficient signal for nuanced distinctions.
-For a given label $l$, if it is truly positive ($y_{il}=+1$), the model might still assign a low score $m(x_{i};\mathbf{w}_{l})$ or, more critically, assign a higher score to an incorrect label $k$ such that $y_{ik}=0$, i.e., a confusing negative that the model mistakenly ranks above the true label.
-Conversely, if label $l$ is truly negative ($y_{il}=0$), the model might erroneously assign it a high score $m(x_{i};\mathbf{w}_{l})$ (a confusing positive).
-Existing fairness interventions often target single-label settings (Zafar et al., 2017; Hardt et al., 2016; Dwork et al., 2011) and are not readily adapted to the multi-label context and these symmetric discriminative challenges.
-To address this, we introduce FairPO (Fair Preference Optimization), a novel framework that integrates preference-based learning with group robust optimization.
-Inspired by the success of Direct Preference Optimization (DPO) in aligning models with human preferences (Rafailov et al., 2024), FairPO recasts parts of the MLC task as learning explicit preferences.
-We partition the label set $\mathcal{T}$ into a privileged set $\mathcal{P}$, where enhanced, fair performance is paramount, and a non-privileged set $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, where the goal is to maintain robust baseline performance.
-For all the privileged labels $l\in\mathcal{P}$, FairPO employs a conditional objective.
-If a confusing set exists for $(x_{i},l)$—meaning there are confusing negatives $k\in S_{il}^{\text{neg}}$ (where $y_{ik}=0,m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=+1$) or confusing positives $k^{\prime}\in S_{il}^{\text{pos}}$ (where $y_{ik^{\prime}}=+1,m(x_{i};\mathbf{w}_{k^{\prime}})\leq m(x_{i};\mathbf{w}_{l})$ when $y_{il}=0$)—a preference loss (inspired by DPO (Rafailov et al., 2024), SimPO (Meng et al., 2024), or CPO (Xu et al., 2024)) is applied.
-This loss encourages $m(x_{i};\mathbf{w}_{l})\gg m(x_{i};\mathbf{w}_{k})$ for the $(l,k)$ pair if $y_{il}=+1$, or $m(x_{i};\mathbf{w}_{l})\ll m(x_{i};\mathbf{w}_{k^{\prime}})$ for the $(l,k^{\prime})$ pair if $y_{il}=0$.
-If no such confusing examples are found for $(x_{i},l)$, a standard BCE loss for label $l$ is applied to ensure consistent learning.
-For non-privileged labels $j\in\bar{\mathcal{P}}$, FairPO enforces a constraint that their classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ does not substantially exceed that of a reference model $\hat{\mathbf{w}}_{j}$, i.e., $\ell(\mathbf{w}_{j})\leq\ell(\hat{\mathbf{w}}_{j})+\epsilon$.
-To manage the trade off between these two distinct objectives, FairPO leverages a group specific robust optimization technique—Group Robust Preference Optimization (GRPO) (Ramesh et al., 2024).
-The overall learning problem is formulated as a standard minimax optimization: $\min_{\{\mathbf{w}_{t}\}_{t\in\mathcal{T}}}\max_{\alpha\in\Delta_{1}}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right]$, where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ are the aggregate losses for the privileged and non-privileged groups respectively, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ are adaptive weights in the simplex $\Delta_{1}$.
-This dynamically balances training, preventing performance degradation in one group for gains in another, thereby promoting fairness across the labels (see Figure 1).
-Our primary contributions are a novel framework for fair MLC using preference signals, a conditional objective targeting hard examples, and a robust optimization strategy to manage fairness trade-offs.
-Our comprehensive experiments show that FairPO, particularly the CPO-variant, achieves significant gains—for instance, up to 3.44% mAP on the least frequent labels of COCO (Lin et al., 2015)—while robustly maintaining performance on non-privileged labels.
-We also outline the framework’s applicability to attribute generation in Appendix A.
-2 Preliminaries: Preference Optimization Methods
-Our framework builds upon recent preference optimization techniques. We briefly review the key formulations.
-Direct Preference Optimization (DPO): DPO (Rafailov et al., 2024) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$.
-Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
-$h_{\pi_{\theta}}(x,y_{w},y_{l})$ $=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)},$ (1)
-$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(h_{\pi_{\theta}}(x,y_{w},y_{l})\right)\right].$ (2)
-where $\sigma$ is the sigmoid function and $\beta$ controls the deviation from $\pi_{\text{ref}}$.
-Group Robust Preference Optimization (GRPO): GRPO (Ramesh et al., 2024) extends preference optimization to handle diverse preference groups $\{D_{g}\}_{g=1}^{K}$.
-Instead of minimizing the average loss, GRPO minimizes the worst-case loss across groups using a robust objective:
-$\min_{\pi_{\theta}}\max_{\alpha\in\Delta_{K-1}}\sum_{g=1}^{K}\alpha_{g}L_{\text{Pref}}(\pi_{\theta};\pi_{\text{ref}},D_{g}),$ (3)
-where $L_{\text{Pref}}$ is a base preference loss (like $L_{\text{DPO}}$), and $\alpha=(\alpha_{1},...,\alpha_{K})$ are adaptive weights in the probability simplex $\Delta_{K-1}$.
-The optimization dynamically increases weights $\alpha_{g}$ for groups with higher current loss, focusing learning on the worst-performing groups.
-Simple Preference Optimization (SimPO): SimPO (Meng et al., 2024) aims to align the implicit reward with generation metrics and eliminates the need for $\pi_{\text{ref}}$.
-It uses the length-normalized average log-likelihood as the reward: $r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y|x)$.
-It also introduces a target margin $\gamma>0$ into the preference model.
-The resulting SimPO loss is:
-$L_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$ (4)
-Contrastive Preference Optimization (CPO): CPO (Xu et al., 2024) also removes the dependency on $\pi_{\text{ref}}$ for efficiency, approximating the DPO objective.
-It combines a reference-free preference loss with a negative log-likelihood (NLL) regularizer on preferred responses $y_{w}$ to maintain generation quality:
-$L_{\text{prefer}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)\right]$ (5)
-$L_{\text{NLL}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w})\sim D}[\log\pi_{\theta}(y_{w}|x)]$ (6)
-$L_{\text{CPO}}(\pi_{\theta})$ $=L_{\text{prefer}}+L_{\text{NLL}}.$ (7)
-This formulation avoids the computational cost of the reference model pass.
-3 Methodology: Fair Preference Optimization (FairPO)
+The Challenge of Performance Disparity in MLC.
+Standard multi-label classification (MLC) models assume all labels are equal, minimizing an aggregate loss over the entire label set (kiritchenko2006hierarchical; DBLP:journals/corr/WeiXLZ0G16; schietgat2010predicting).
+This assumption fails in real-world scenarios where data asymmetries cause significant performance disparities.
+These disparities stem from several factors, including class imbalance, where models favor frequent labels over rare ones (Charte2015AddressingII; 6471714), and varying real-world importance, such as distinguishing a critical disease from a benign one (rajpurkar2017chexnetradiologistlevelpneumoniadetection; 10368872).
+Disparities also arise from semantic complexity, where models learn simple concepts but fail on nuanced ones.
+Consequently, training becomes dominated by the easy majority, yielding systems that are unreliable for the very cases that matter most (swayamdipta2020datasetcartographymappingdiagnosing; pleiss2020identifyingmislabeleddatausing).
+Limitations of Conventional Methods.
+Conventional methods are limited by the aggregate nature of the Binary Cross-Entropy (BCE) loss, which allows the learning signal for underrepresented labels to be drowned out by easy ones (bce_loss; Charte2015AddressingII).
+While solutions like Focal Loss help, they are fundamentally restricted because they treat each label’s prediction in isolation (lin2018focallossdenseobject).
+They do not teach the model to perform discrimination by resolving confusion between a true label and a specific negative alternative.
+This reveals a gap for a framework that can learn relative preferences, improving performance on difficult labels while maintaining it for others.
+A preference-based objective forces the model to learn that a true label’s score is explicitly preferred over that of a confusing competitor.
+Our Approach: The FairPO Framework.
+We introduce FairPO (Fair Preference Optimization), a framework to manage performance disparities by partitioning labels into a privileged group for targeted improvement and a non-privileged group for performance maintenance.
+We prioritize learning on the privileged group using a preference-based objective, a novel adaptation of techniques from generative model alignment for a discriminative task.
+For each privileged true label, FairPO dynamically identifies its confusing counterparts, incorrect labels with misleadingly high scores and trains the model to prefer the true label, creating a large discriminative margin.
+Concurrently, a constrained objective safeguards the non-privileged group against performance degradation relative to a reference model.
+These competing objectives are dynamically balanced using a Group Robust Preference Optimization (GRPO) formulation (ramesh2024group).
+Figure 1: FairPO Framework
+Application and Contributions.
+Our general FairPO framework can address any MLC task requiring label prioritization due to factors like class imbalance or real-world importance.
+To demonstrate its effectiveness, we apply it to class imbalance, defining infrequent tail labels as the privileged group and frequent head labels as the non-privileged one.
+Our main contributions are:
+  1. 1.
+    We propose FairPO, a novel framework for targeted performance improvement in MLC applicable to various sources of disparity.
+  2. 2.
+    We are the first to adapt preference optimization techniques (meng2024simposimplepreferenceoptimization; xu2024contrastivepreferenceoptimizationpushing) from generative modeling to create discriminative objectives for MLC.
+  3. 3.
+    We successfully apply GRPO to solve the trade-off between objectives for the privileged and non-privileged groups, a novel application in this context.
+Experiments show FairPO significantly boosts performance on infrequent labels by up to 3.59% mAP while robustly maintaining performance on frequent ones, outperforming several baselines.
+2 Methodology: The FairPO Framework
+Our methodology builds upon a foundation of preference optimization techniques.
+To provide the necessary background on these methods, namely DPO, CPO, SimPO, and GRPO, we have moved the detailed review of their core formulations to Appendix A.
 2.1 Problem Setup and Fairness Goals
-Given a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, we fine-tune parameters $\mathbf{w}_{t}$ for a per-label classifier that assigns a score $m(x_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$.
-Our framework leverages a reference model with parameters $\hat{\mathbf{w}}_{t}$, typically obtained from standard supervised fine-tuning (SFT).
-Our framework addresses the problem of group fairness, where the goal is to ensure equitable performance across different predefined groups.
-In our MLC context, we define these groups not by instance attributes (e.g., demographics), but by partitioning the labels themselves into a privileged set $\mathcal{P}$ (e.g., rare, critical, or historically underperforming labels) and a non-privileged set $\bar{\mathcal{P}}$.
-Our fairness objective is a nuanced form of equitable treatment: we aim to significantly improve performance for the underperforming privileged group while ensuring that performance for the non-privileged group is not substantially harmed.
-This goal of targeted improvement without undue harm to other groups is enforced by our robust optimization framework.
-For any privileged label $l\in\mathcal{P}$, the model must accurately discriminate its true state from confusing alternatives.
-For any non-privileged label $j\in\bar{\mathcal{P}}$, performance should not degrade significantly from the reference model.
-To formalize this, we define two types of confusing sets for a privileged label $l\in\mathcal{P}$: the set of confusing negatives, $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(x_{i};\mathbf{w}_{k})\geq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=+1$, and the set of confusing positives, $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=+1\text{ and }m(x_{i};\mathbf{w}_{k})\leq m(x_{i};\mathbf{w}_{l})\}$ when $y_{il}=0$.
-The overall confusing set is $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
-3.2 Objective for Privileged Labels ($l\in\mathcal{P}$)
-Confusing Examples Exist ($S_{il}\neq\emptyset$): We employ a DPO-inspired (Rafailov et al., 2024) preference loss.
-If $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$ is sampled, we prefer $l$ over $k$ ($l\succ k$).
-The DPO term is:
-$h_{\mathbf{w}}(x_{i},l,k)=\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right).$ (8)
-The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},l,k))$.
-If $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$ is sampled, we prefer $l$ over $k$ (i.e., the true negative $l$ is preferred over the confusing positive $k$, meaning $m(x_{i};\mathbf{w}_{l})$ should be lower than $m(x_{i};\mathbf{w}_{k})$ relative to the reference).
-The DPO term is (note the swapped order for $k$ and $l$ to reflect $k\succ l$ effectively aiming for $m(x_{i};\mathbf{w}_{l})<m(x_{i};\mathbf{w}_{k})$):
-$h_{\mathbf{w}}(x_{i},k,l)=\left(\log\frac{m(x_{i};\mathbf{w}_{k})}{m(x_{i};\hat{\mathbf{w}}_{k})}\right)-\left(\log\frac{m(x_{i};\mathbf{w}_{l})}{m(x_{i};\hat{\mathbf{w}}_{l})}\right).$ (9)
-The loss contribution is $\log\sigma(\beta\cdot h_{\mathbf{w}}(x_{i},k,l))$.
-No Confusing Examples ($S_{il}=\emptyset$): We revert to a standard base classification loss (e.g., Binary Cross-Entropy, BCE) for label $l$.
-This BCE fallback ensures that the model continues to receive learning signals even for labels without any identified confusing counterparts, thereby avoiding training stagnation.
-$\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})=-[y_{il}\log m(x_{i};\mathbf{w}_{l})+(1-y_{il})\log(1-m(x_{i};\mathbf{w}_{l}))].$ (10)
-The overall Privileged Loss $\mathcal{L}_{\mathcal{P}}$ is the expectation over these conditional losses:
-$\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t}|t\in\mathcal{T}\},\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\})$ $=\mathbb{E}_{(x_{i},l)\text{ s.t.
-}l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot(-\log\sigma(\beta\cdot h_{\text{DPO}}))\right$.
-$\quad\left.+\mathbf{1}_{S_{il}=\emptyset}\cdot\ell_{\text{BCE}}(x_{i},y_{il};\mathbf{w}_{l})\right]$ (11)
-where $h_{\text{DPO}}$ refers to the appropriate term from Eq. 8 or Eq. 9. The hyperparameter $\beta$ controls the preference strength.
-SimPO-Inspired Loss: We also adapt insights from Simple Preference Optimization (SimPO) (Meng et al., 2024).
-Original SimPO for generative models uses length-normalized log-likelihoods and a target margin $\gamma$ in its preference scoring.
-Since our multi-label classification (MLC) setup deals with individual label scores $m(x_{i};\mathbf{w}_{t})$ where sequence length is inapplicable, we adapt SimPO’s core concept of a target margin $\gamma$ to our context.
-This results in a reference-free preference component that aims to ensure a minimum separation for preferred label scores, omitting length normalization but retaining the margin:
-$L_{\text{pref}}^{\text{SimPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)-\gamma\right),$ (12)
-where $(m(x_{i};\text{preferred}),m(x_{i};\text{dispreferred}))$ are $(m(x_{i};\mathbf{w}_{l}),m(x_{i};\mathbf{w}_{k}))$ if $y_{il}=+1$ and $k\in S_{il}^{\text{neg}}$, or $(m(x_{i};\mathbf{w}_{k}),m(x_{i};\mathbf{w}_{l}))$ if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$.
-The term $\beta\log(\cdot)$ captures the relative preference, while $\gamma$ enforces a desired separation margin.
-This $L_{\text{pref}}^{\text{SimPO}}$ replaces the DPO term in Eq.
-11 when $S_{il}\neq\emptyset$, with the BCE fallback (Eq.
-10) for $S_{il}=\emptyset$ remaining.
-More broadly, this adaptation leverages SimPO’s margin mechanism for more distinct score separation in challenging MLC cases.
-CPO-Inspired Loss: Adapting Contrastive Preference Optimization (Xu et al., 2024). If $S_{il}\neq\emptyset$, the preference component is:
-$L_{\text{pref}}^{\text{CPO}}(x_{i},l,k)=-\log\sigma\left(\beta\left(\log\frac{m(x_{i};\text{preferred})}{m(x_{i};\text{dispreferred})}\right)\right).$ (13)
-(using the same preferred/dispreferred logic as SimPO).
-CPO also includes an NLL regularizer.
-Thus, the overall CPO-inspired privileged loss, $\mathcal{L}_{\mathcal{P}}^{\text{CPO}}$, could be formulated by combining this preference term (when $S_{il}\neq\emptyset$) with a consistent NLL/BCE component for all true ground truth labels within $\mathcal{P}$:
-[…]
-where $k_{s}$ is a confusing example sampled from $S_{il}$, and $\lambda_{\text{CPO}}$ is a weighting hyperparameter.
-Note that the $\ell_{\text{BCE}}$ term is always applied, serving as both the NLL regularizer and the fallback for $S_{il}=\emptyset$ if the first term becomes zero.
-These alternatives offer different ways to model preferences.
-The non-privileged loss $\mathcal{L}_{\bar{\mathcal{P}}}$ (Eq.
-17) remains unchanged regardless of the privileged loss choice.
-3.3 Constrained Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$)
-For non-privileged labels $j\in\bar{\mathcal{P}}$, our goal is to maintain performance relative to the reference model $\hat{\mathbf{w}}_{j}$, preventing degradation due to the focus on privileged labels.
-We use a standard base classification loss $\ell(\mathbf{w}_{j};x_{i},y_{ij})$ (e.g., Binary Cross-Entropy) defined as:
-$\ell(\mathbf{w}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\mathbf{w}_{j})+(1-y_{ij})\log(1-m(x_{i};\mathbf{w}_{j}))],$ (15)
-$\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})$ $=-[y_{ij}\log m(x_{i};\hat{\mathbf{w}}_{j})+(1-y_{ij})\log(1-m(x_{i};\hat{\mathbf{w}}_{j}))].$ (16)
-The Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$ employs a hinge mechanism, penalizing the model only if its loss $\ell(\mathbf{w}_{j})$ exceeds the reference model’s loss $\ell(\hat{\mathbf{w}}_{j})$ by more than a predefined slack margin $\epsilon\geq 0$:
-$\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})=\mathbb{E}_{(x_{i},j)\text{ s.t.
-}j\in\bar{\mathcal{P}}}\left[\max(0,\ell(\mathbf{w}_{j};x_{i},y_{ij})-\ell(\hat{\mathbf{w}}_{j};x_{i},y_{ij})-\epsilon)\right].$ (17)
-This ensures $\mathbf{w}_{j}$ is primarily updated only if performance on label $j$ drops significantly below the reference baseline plus $\epsilon$.
-3.4 Group Robust Optimization Formulation
-To effectively balance the objective for privileged labels ($\mathcal{L}_{\mathcal{P}}$) and the constraint for non-privileged labels ($\mathcal{L}_{\bar{\mathcal{P}}}$), we employ the Group Robust Preference Optimization (GRPO) framework (Ramesh et al., 2024).
-We define two distinct groups corresponding to our label partitions:
-  • $G_{\mathcal{P}}$: Associated with the Privileged Loss $\mathcal{L}_{\mathcal{P}}$, focusing on triplets $(x_{i},l,k)$.
-  • $G_{\bar{\mathcal{P}}}$: Associated with the Non-Privileged Loss $\mathcal{L}_{\bar{\mathcal{P}}}$, focusing on pairs $(x_{i},j)$.
-The FairPO objective becomes the GRPO minimax problem:
-$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\,\geq\,0,\,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\,=\,1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\{\mathbf{w}_{t},\hat{\mathbf{w}_{t}}|t\in\mathcal{T}\})\right].$ (18)
-Here, $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are adaptive weights representing the importance assigned to each group’s loss.
-The inner maximization finds the worst-case distribution over group losses (by increasing the weight $\alpha$ for the group with higher current loss), while the outer minimization seeks model parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}$ that perform well even under this worst-case weighting.
-This formulation inherently promotes robustness and fairness by preventing the optimization from disproportionately favouring one group at the expense of the other.
+Let us consider a standard multi-label classification setting with a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}$ is an input instance and $y_{i}\in\{0,1\}^{|\mathcal{T}|}$ is the corresponding multi-label vector over a universe of $|\mathcal{T}|$ labels.
+The goal is to train a model, parameterized by weights $\mathbf{w}$, that produces a set of per-label scores $m(\mathbf{x}_{i};\mathbf{w}_{t})$ for each label $t\in\mathcal{T}$.
+Our framework also utilizes a pre-trained reference model with parameters $\hat{\mathbf{w}}_{t}$.
+Conventional methods often train such models by minimizing a single, aggregate loss where each label contributes equally.
+However, as motivated in our introduction, real-world applications frequently demand a more nuanced approach where certain labels are prioritized.
+Our framework, FairPO, is explicitly designed for these scenarios.
+The core idea is to partition the total label set $\mathcal{T}$ into two disjoint subsets:
+  • A privileged group $\mathcal{P}\subset\mathcal{T}$, which contains labels for which we seek to significantly enhance the model’s performance.
+  • A non-privileged group $\bar{\mathcal{P}}=\mathcal{T}\setminus\mathcal{P}$, which contains the remaining labels, for which our objective is to robustly maintain at least a baseline level of performance.
+This partitioning is a key strength of our framework’s general design.
+The criteria for assigning labels to the privileged group are flexible and can be adapted to the specific problem at hand, such as using label frequency for class imbalance, domain knowledge for real-world importance, or annotation consistency for data quality issues.
+In this work, to demonstrate FairPO’s efficacy, we focus on the class imbalance problem, where the privileged group consists of the least frequent labels.
+We defer the precise details of this setup to Section 3.
+2.2 FairPO Objectives
+FairPO’s methodology is built on three core components: a conditional objective for privileged labels, a constrained objective for non-privileged labels, and a robust optimization framework to balance them.
+The detailed mathematical formulations for each component are provided in Appendix B.
+Objective for Privileged Labels ($l\in\mathcal{P}$): To improve performance on critical labels, we employ a conditional objective designed to target hard-to-discriminate cases.
+When a confusing set of labels exists for an instance (i.e., high-scoring negatives or low-scoring positives), a preference loss inspired by DPO (rafailov2023direct) is applied to enforce correct relative rankings.
+In the absence of such confusing examples, the objective reverts to a standard BCE loss, which acts as a crucial fallback for stable training.
+Our framework is versatile and also supports reference-free preference loss variants inspired by CPO (xu2024contrastivepreferenceoptimizationpushing) and SimPO (meng2024simposimplepreferenceoptimization).
+Objective for Non-Privileged Labels ($j\in\bar{\mathcal{P}}$): To maintain baseline performance on the remaining labels, we use a constrained objective.
+This objective employs a hinge mechanism that only incurs a penalty if the model’s performance on a label drops significantly below that of a reference model by a predefined slack margin $\epsilon$.
+This acts as a protective measure, preventing substantial performance degradation for the non-privileged group.
+Group Robust Optimization: These two distinct objectives for the privileged and non-privileged groups are adaptively balanced using the Group Robust Preference Optimization (GRPO) framework (ramesh2024group).
+The overall learning problem is formulated as the following minimax objective:
+$\min_{\{\mathbf{w}_{t}|t\in\mathcal{T}\}}\max_{\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}(\cdot)+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}(\cdot)\right].$ (1)
+This formulation dynamically adjusts the training focus between the privileged and non-privileged groups, seeking a solution that is robust to the worst-case group loss and thereby managing the fairness-performance trade-off.
 2.3 Optimization Algorithm
-The FairPO framework solves the minimax objective (Eq.
-18) iteratively.
-A high-level overview is in Algorithm 1, with full details in Algorithm 2 (see Appendix B).
-In each training iteration, an instance and a label $r$ are sampled.
-The corresponding loss ($\mathcal{L}_{\mathcal{P}}^{(s)}$ or $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ via Eq.
-11 or Eq.
-17) and its gradient are computed based on whether $r$ is in the privileged set $\mathcal{P}$ or non-privileged set $\bar{\mathcal{P}}$.
-The group weights $\alpha^{(s+1)}$ are updated using a mirror ascent step.
-A crucial aspect for the stability and effectiveness of this step is loss scaling.
-Instead of directly using the raw current loss values $\mathcal{L}_{g}^{(s)}$ (for $g\in\{\mathcal{P},\bar{\mathcal{P}}\}$) in the exponent, we use a scaled version.
-Specifically, for each group $g$, we maintain a running average of its loss from previous steps, denoted $\bar{\mathcal{L}}_{g}^{\text{avg}}$.
-The term used to update the weights is then based on the relative change of the current loss from this average, for example, a scaled difference like $(\mathcal{L}_{g}^{(s)}-\bar{\mathcal{L}}_{g}^{\text{avg}})/(\bar{\mathcal{L}}_{g}^{\text{avg}}+\delta)$, where $\delta$ is a small constant for numerical stability.
-This normalization prevents instability due to scale disparities in group loss values and makes the optimization more sensitive to relative improvements or regressions, rather than absolute magnitudes.
-This scaled loss, say $\Delta\tilde{\mathcal{L}}_{g}^{(s)}$, is then used in the exponential update: $\alpha_{g}^{\text{new}}\leftarrow\alpha_{g}^{(s)}\exp(\eta_{\alpha}\Delta\tilde{\mathcal{L}}_{g}^{(s)})$, followed by normalization so $\sum_{g}\alpha_{g}^{\text{new}}=1$.
-This scaling normalizes loss magnitudes across groups, makes weight updates sensitive to significant performance changes relative to recent history, and improves overall stability of the $\alpha$ dynamics.
-Subsequently, model parameters $\mathbf{w}^{(s+1)}$ are updated via mirror descent, using a gradient that combines group gradients weighted by the adaptively balanced $\alpha^{(s+1)}$.
+We solve the minimax objective in Eq.
+17 iteratively using alternating mirror descent (Algorithm 1).
+A key component for stability is loss scaling during the update of the group weights $\alpha$.
+Since the group losses, $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$, can have different scales and variances, using raw loss values for the mirror ascent step is unstable.
+To mitigate this, we update $\alpha$ based on the relative change of each group’s loss from its running average, $\bar{\mathcal{L}}_{g}^{\text{avg}}$, ensuring a more stable optimization.
+Further details are in Appendix E.
 Algorithm 1 FairPO Training Overview (DPO-inspired)
@@
 5:    If $r\in\mathcal{P}$:
-6:     Compute privileged loss $\mathcal{L}_{\mathcal{P}}^{(s)}$ for $(x_{i},r)$ (Eq. 11, conditionally using preference loss like Eq. 8/9 or BCE Eq. 10).
+6:     Compute privileged loss $\mathcal{L}_{\mathcal{P}}^{(s)}$ for $(x_{i},r)$ (DPO Eq. 13, or BCE Eq. 11).
 7:    Else if $r\in\bar{\mathcal{P}}$:
@@
 12:Return Optimized parameters $\{\mathbf{w}_{t}|t\in\mathcal{T}\}^{(S)}$.
-13:Full details are in Algorithm 2 (see Appendix B).
+13:Full details are in Algorithm 2 (see Appendix E).
 3 Experimental Setup
-We evaluate FairPO on two standard multi-label image classification benchmarks: MS-COCO 2014 (Lin et al., 2015) and NUS-WIDE (Chua et al., 2009).
-For our fairness setup, we define the privileged group ($\mathcal{P}$) as the 20% least frequent labels in each training set, with the remaining 80% forming the non-privileged group ($\bar{\mathcal{P}}$).
-Our base model is a Vision Transformer (ViT) (Dosovitskiy et al., 2021), where we fine-tune label-specific classifier heads on top of its features.
-The reference model parameters $\hat{\mathbf{w}}$ are obtained from a standard supervised fine-tuning (SFT) of this architecture with a Binary Cross-Entropy (BCE) loss.
-Performance is assessed separately for $\mathcal{P}$ and $\bar{\mathcal{P}}$ sets using standard metrics: Mean Average Precision (mAP), Sample F1, Exact Match Ratio (EMR), and Accuracy.
-We compare FairPO against three baselines: (1) BCE-SFT, which also serves as our reference model; (2) BCE-SFT + Re-Weighting (RW), a simple loss up-weighting for $\mathcal{P}$ labels; and (3) Group DRO + BCE (Sagawa et al., 2020), which minimizes the worst-group loss.
-We test three variants of our framework, differing in their preference loss for privileged labels: FairPO-DPO (Eq.
-11), FairPO-SimPO (Eq.
-12), and FairPO-CPO (Eq.
-14).
-Comprehensive details on datasets, preprocessing, model architecture, baseline implementations, and hyperparameter tuning are deferred to Appendices C, D, and E.
+We evaluate FairPO on two multi-label benchmarks: MS-COCO 2014 (lin2015microsoftcococommonobjects) and NUS-WIDE (chua2009nus).
+To address class imbalance, we partition labels by frequency: the 20) and the remaining 80).
+While this partitioning is effective, the FairPO framework is general and could use other criteria like domain importance.
+We assess performance using mAP, Sample F1, and Accuracy, focusing on the mAP gain on the privileged set, , relative to the BCE-SFT baseline.
+We compare against baselines including Group DRO + BCE (sagawa2019distributionally) and Focal Loss (lin2018focallossdenseobject).
+Our base model is a frozen Vision Transformer (ViT) (dosovitskiy2021imageworth16x16words) with separate non-linear MLP heads per label, trained with AdamW (loshchilov2019decoupledweightdecayregularization).
+The trained BCE-SFT baseline provides the reference model parameters.
+We test three framework variants: FairPO-DPO, FairPO-SimPO, and FairPO-CPO.
+Full details are in Appendix F.
 4 Results and Analysis
-Table 1: Performance comparison on MS-COCO.
-$\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%).
-Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
-Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 86.32 90.65 61.43 64.89 94.89 98.12 35.78 36.98 Ref BCE SFT + RW 87.25 89.85 62.68 64.11 95.95 97.93 47.43 33.81 +0.93 GDRO + BCE 87.92 90.41 62.31 64.75 95.72 98.05 46.12 35.16 +1.60 FairPO-DPO 88.02 89.97 63.45 63.65 97.89 98.04 62.19 35.12 +1.70 FairPO-SimPO 87.67 88.76 62.34 63.12 95.69 97.78 45.32 32.34 +1.35 FairPO-CPO 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44
-Table 2: Performance comparison on NUS-WIDE.
-$\mathcal{P}$ denotes the privileged label set (20% least frequent), and $\bar{\mathcal{P}}$ denotes the non-privileged set (remaining 80%).
-Best results for $\mathcal{P}$ metrics and $\Delta\text{mAP}(\mathcal{P})$ are in bold.
-Method mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ BCE SFT 63.53 70.24 48.12 55.83 91.51 95.22 19.32 11.56 Ref BCE SFT + RW 65.12 69.14 49.51 54.73 92.33 94.81 21.23 10.32 +1.59 GDRO + BCE 64.84 69.91 49.13 55.62 92.11 95.13 21.02 11.34 +1.31 FairPO-DPO 66.34 69.05 51.71 54.52 93.92 95.04 27.91 11.21 +2.81 FairPO-SimPO 64.11 68.03 48.82 53.81 91.94 94.52 20.18 10.19 +0.58 FairPO-CPO 67.12 69.83 52.21 55.24 94.31 95.12 31.87 11.25 +3.59
-Tables 1 and 2 detail our findings.
-The standard BCE-SFT baseline confirms the fairness problem central to our motivation: on MS-COCO, a significant performance gap exists between the privileged ($\mathcal{P}$) labels (86.32% mAP) and the non-privileged ($\bar{\mathcal{P}}$) labels (90.65% mAP).
-This disparity highlights how standard aggregate losses can fail to ensure equitable performance for rare or difficult categories.
-Our results demonstrate that FairPO variants, particularly FairPO-CPO, effectively address this challenge by navigating the fairness-performance trade-off in a controlled manner.
-On MS-COCO, FairPO-CPO achieves a substantial +3.44% mAP gain for $\mathcal{P}$ labels, while performance on $\bar{\mathcal{P}}$ labels dips by a negligible 0.31%.
-Similarly, on NUS-WIDE, it yields a +3.59% mAP gain for a minor 0.41% drop.
-This is not an uncontrolled side effect but a managed outcome, enforced by the GRPO mechanism and the constrained loss (Eq.
-17), which together prevent significant harm to the non-privileged group while targeting improvements for the privileged one.
-By directly optimizing the model to rank true labels higher than their specific, dynamically identified confusing counterparts, FairPO sharpens discriminative power for the most challenging cases.
-Among the variants, the reference-free FairPO-CPO proves most effective.
-This directly addresses a potential limitation of DPO-style methods, which can be sensitive to the quality of the reference model.
-FairPO-CPO’s strong performance, likely due to its dual objective of optimizing both relative preference and absolute correctness (via its NLL-like regularizer), demonstrates its robustness and versatility, making it a more practical choice than the reference-dependent FairPO-DPO or the fixed-margin FairPO-SimPO.
+As shown in Tables 1 and 2, FairPO-CPO consistently achieves the highest performance on the privileged group ($\mathcal{P}$), confirming its effectiveness.
+On MS-COCO, it delivers a statistically significant gain of +3.44% $\Delta\text{mAP}(\mathcal{P})$ over strong baselines like BCE-SFT.
+Crucially, this targeted improvement comes with only a minor and statistically insignificant performance drop on the non-privileged group ($\bar{\mathcal{P}}$), demonstrating FairPO’s ability to reallocate model capacity without causing measurable harm.
+The superiority of the CPO variant stems from its robust design: unlike DPO, it is reference-free, and its integrated BCE regularizer provides a crucial signal for absolute correctness that SimPO lacks.
+This dual objective of learning both relative preferences and absolute scores yields a more stable and effective training process, making FairPO-CPO the most practical choice.
+[…]
+Method mAP Sample F1 Accuracy $\Delta$ mAP ($\mathcal{P}$) $\Delta$ mAP ($\bar{\mathcal{P}}$) $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ (vs.
+Focal) (vs.
+BCE) BCE-SFT (bce_loss) 86.32±0.11 90.65±0.08 61.43±0.15 64.89±0.12 94.89±0.09 98.12±0.05 -2.03 Ref GDRO + BCE (sagawa2019distributionally) 87.92±0.12 90.41±0.09 62.31±0.20 64.75±0.13 95.72±0.10 98.05±0.05 -0.43 -0.24 Focal Loss (lin2018focallossdenseobject) 88.35±0.14 89.81±0.12 63.15±0.16 64.18±0.15 96.11±0.12 97.90±0.07 Ref -0.84 FairPO-DPO 88.02±0.15 89.97±0.11 63.45±0.17 63.65±0.16 97.89±0.13 98.04±0.06 -0.33 -0.68 FairPO-SimPO 87.67±0.18 88.76±0.21 62.34±0.22 63.12±0.19 95.69±0.15 97.78±0.09 -0.68 -1.89 FairPO-CPO 89.76±0.09 90.34±0.07 64.01±0.13 64.32±10 98.03±0.07 98.06±0.05 +1.41 -0.31
+Table 2: Performance comparison on NUS-WIDE. Notations are similar to Table 1.
+Method mAP Sample F1 Accuracy $\Delta$ mAP ($\mathcal{P}$) $\Delta$ mAP ($\bar{\mathcal{P}}$) $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ (vs.
+Focal) (vs.
+BCE) BCE SFT (bce_loss) 63.53±0.15 70.24±0.11 48.12±0.21 55.83±0.14 91.51±0.12 95.22±0.08 -2.28 Ref GDRO + BCE (sagawa2019distributionally) 64.84±0.16 69.91±0.12 49.13±0.22 55.62±0.15 92.11±0.13 95.13±0.09 -0.97 -0.33 Focal Loss (lin2018focallossdenseobject) 65.81±0.17 68.95±0.15 50.33±0.20 54.31±0.18 93.05±0.11 94.75±0.11 Ref -1.29 FairPO-DPO 66.34±0.20 69.05±0.14 51.71±0.25 54.52±0.19 93.92±0.16 95.04±0.09 +0.53 -1.19 FairPO-SimPO 64.11±0.22 68.03±0.24 48.82±0.28 53.81±0.22 91.94±0.18 94.52±0.13 -1.70 -2.21 FairPO-CPO 67.12±0.14 69.83±0.10 52.21±0.19 55.24±0.13 94.31±0.10 95.12±0.08 +1.31 -0.41
 5 Ablation Studies
+Ablation studies on MS-COCO with FairPO-CPO confirm that each component is critical (Tables 3 & 4).
+Removing the preference loss, the non-privileged constraint, or the adaptive GRPO balancing all lead to significant performance degradation.
+A non-targeted Global CPO variant, while better than standard BCE, is substantially less effective than the full FairPO model, highlighting the necessity of our targeted approach.
+The design of the preference objective itself is also vital: restricting the preference signal to only Confusing Negatives or removing the BCE Fallback for non-confusing cases both degrade performance and stability.
+These results demonstrate that while preference optimization is powerful, its effectiveness hinges on the complete, balanced FairPO framework.
 Table 3: Ablation on core components of FairPO-CPO (MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT. Parentheses show change vs Full FairPO-CPO.
-Method Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 48.51 35.30 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) (-16.92) (+0.07) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 63.12 31.95 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) (-2.31) (-3.28) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 56.70 34.15 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) (-8.73) (-1.08) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 55.20 37.15 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (-0.47) (-1.08) (+0.05) (-10.23) (+1.92)
-Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT (86.32). Parentheses show change vs Full FairPO-CPO.
-Preference Detail Variant mAP Sample F1 Accuracy EMR $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 65.43 35.23 +3.44 (Conf.
-Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 22.54 35.10 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) (-42.89) (-0.13) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 60.75 34.90 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07) (-4.68) (-0.33)
-We conduct ablation studies on MS-COCO using FairPO-CPO to assess component contributions (Tables 3 and 4).
-Our analysis of core components (Table 3) confirms the criticality of each part.
-Replacing the preference loss with BCE reduces privileged gains, removing the $\bar{\mathcal{P}}$ constraint harms non-privileged performance, and using fixed weights instead of GRPO degrades both groups.
-To further isolate the value of our fairness framework, we tested a Global CPO variant that applies the preference objective to all labels without the privileged/non-privileged split or GRPO.
-While this proves the preference objective is a strong general-purpose loss (achieving a +2.23% mAP gain over BCE), the full FairPO framework is decisively superior for the targeted fairness task.
-By applying CPO globally, the performance gains are unfocused, resulting in a 1.21 point lower mAP on the privileged set.
-This demonstrates that the full FairPO architecture is crucial for focusing the model’s capacity to resolve the hardest discrimination challenges within the privileged group.
-Furthermore, the design of the preference objective itself is crucial (Table 4).
-Restricting it to only Confusing Negatives (for $y_{il}=+1$) causes a profound performance drop, as this neglects the far more common scenario of ranking true negatives below confusing positives for rare labels, making the learning signal exceptionally sparse.
-Similarly, removing the BCE Fallback for non-confusing instances degrades performance, proving that a standard classification signal on easier cases is vital for model stability.
-Both aspects of our full formulation are thus essential, demonstrating that while targeted preference optimization is powerful, it must be complemented by standard losses and robust balancing for stable, effective training.
+Method Variant mAP Sample F1 Accuracy $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 +3.44 w/o Preference Loss 88.12 90.45 62.45 64.80 95.80 98.09 +1.80 ($\mathcal{L}_{\mathcal{P}}$ is BCE) (-1.64) (+0.11) (-1.56) (+0.48) (-2.23) (+0.03) w/o $\bar{\mathcal{P}}$ Constraint 89.55 88.98 63.70 62.95 97.90 97.55 +3.23 ($\mathcal{L}_{\bar{\mathcal{P}}}$ is BCE) (-0.21) (-1.36) (-0.31) (-1.37) (-0.13) (-0.51) w/o GRPO 88.48 89.75 62.88 63.50 96.53 97.88 +2.16 (Fixed 0.5/0.5 weights) (-1.28) (-0.59) (-1.13) (-0.82) (-1.50) (-0.18) Global CPO (on all labels) 88.55 90.68 62.75 64.85 96.95 98.11 +2.23 (No $\mathcal{P}$$\bar{\mathcal{P}}$ split or GRPO) (-1.21) (+0.34) (-1.26) (+0.47) (-1.08) (+0.05)
+Table 4: Ablation on preference formulation (FairPO-CPO, MS-COCO). $\Delta\text{mAP}(\mathcal{P})$ vs BCE SFT. Parentheses show change vs Full FairPO-CPO.
+Preference Detail Variant mAP Sample F1 Accuracy $\Delta\text{mAP}(\mathcal{P})$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ $\mathcal{P}$ $\bar{\mathcal{P}}$ FairPO-CPO (Full) 89.76 90.34 64.01 64.32 98.03 98.06 +3.44 (Conf.
+Neg & Pos, BCE Fallback) Only Confusing Negatives 73.15 90.25 47.88 64.20 94.67 98.01 -13.17 (-16.61) (-0.09) (-16.13) (-0.12) (-3.36) (-0.05) w/o BCE Fallback 89.05 90.21 63.20 64.10 97.55 97.99 +2.73 (No loss if $S_{il}=\emptyset$) (-0.71) (-0.13) (-0.81) (-0.22) (-0.48) (-0.07)
 6 Related Work
-Recent efforts in fair MLC address complex challenges like label imbalance impacting tail labels (Guo and Wang, 2021), subjective fairness (Liu et al., 2023), and class-incremental learning (Dong et al., 2025).
-FairPO contributes a novel approach by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets.
-It applies distinct, fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and manages the trade-off with a robustness framework, differentiating our targeted approach from prior work.
-We adapt recent advances in preference optimization, originally developed for aligning LLMs (Ouyang and Others, 2022; Christiano et al., 2023).
-Techniques like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and its reference-free variants CPO (Xu et al., 2024) and SimPO (Meng et al., 2024) optimize policies from preference pairs.
-Rather than ranking entire outputs, we repurpose these methods to specifically differentiate true label scores from their dynamically identified confusing counterparts within the privileged set, thereby sharpening critical decision boundaries.
-To balance our objectives, we employ a Group Robust Optimization strategy inspired by Group DRO (Sagawa et al., 2020; Rice et al., 2021) and GRPO (Ramesh et al., 2024).
-While these methods typically balance performance across data or preference groups, FairPO uniquely defines its groups by our label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$).
-It then uses GRPO’s adaptive weighting to balance their distinct, custom-formulated loss objectives, providing a principled mechanism for managing the specific fairness-performance trade-offs in our MLC context.
+Our work is situated at the intersection of three research areas, which we detail further in Appendix C.
+First, while recent efforts in fair MLC have addressed challenges such as tail labels (wu2023longtailed) and subjective fairness (gao2023learning), FairPO contributes a novel approach by partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets and applying targeted, distinct objectives to each.
+Second, we adapt modern preference optimization techniques, originally developed for aligning LLMs like DPO (rafailov2023direct), CPO (xu2024contrastivepreferenceoptimizationpushing), and SimPO (meng2024simposimplepreferenceoptimization).
+Instead of ranking entire outputs, we repurpose these methods to resolve ambiguities between true labels and their dynamically identified confusing counterparts.
+Finally, to manage the trade-off between our objectives, we employ Group Robust Optimization.
+Inspired by Group DRO (sagawa2020distributionallyrobustneuralnetworks) and GRPO (ramesh2024group), we uniquely define the groups by our label partitions ($\mathcal{P}$ and $\bar{\mathcal{P}}$) and use GRPO’s adaptive weighting to balance their custom-formulated losses.
 7 Discussion
-In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in multi-label classification.
-Our experiments, particularly with the FairPO-CPO variant, highlight the value of nuanced preference signals for navigating complex fairness-performance trade-offs in challenging discrimination tasks.
-While promising, FairPO has technical limitations: its dynamic confusing set can lead to instability or sparse signals for rare labels; the DPO-based variant and non-privileged constraint rely on a well-calibrated reference model; and GRPO’s balancing of heterogeneous losses requires careful tuning and lacks theoretical convergence guarantees.
-These challenges also define our future work, which includes extending FairPO’s principles to multi-label attribute generation (Appendix A), conducting comprehensive empirical validation on more datasets and modalities, analyzing alternative label partitioning strategies, and pursuing theoretical insights into the framework’s convergence properties.
-Appendix A Adapting FairPO for Multi-Attribute Generation
+In conclusion, we introduced FairPO, a novel framework that effectively integrates preference optimization with group robustness to enhance fairness in MLC, with our CPO variant proving particularly effective.
+While promising, FairPO has limitations, including the potential instability of its dynamic ‘confusing set‘, the DPO variant’s reliance on a reference model, and the need for careful tuning of GRPO’s balancing act.
+These challenges directly motivate our future work, which will focus on extending FairPO to attribute generation (Appendix D), performing broader empirical validation, exploring alternative label partitioning strategies, and developing theoretical insights into the framework’s convergence.
+Appendix A Preliminaries: Preference Optimization Methods
+Our framework builds upon recent preference optimization techniques. We briefly review the key formulations.
+Direct Preference Optimization (DPO): DPO (rafailov2023direct) directly optimizes a policy $\pi_{\theta}$ using preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$ for prompt $x$.
+Assuming a Bradley-Terry preference model tied to an implicit reward function related to $\pi_{\theta}$ and a reference policy $\pi_{\text{ref}}$, DPO maximizes the likelihood of observed preferences, resulting in the loss:
+$h_{\pi_{\theta}}(x,y_{w},y_{l})$ $=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)},$ (2)
+$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(h_{\pi_{\theta}}(x,y_{w},y_{l})\right)\right].$ (3)
+where $\sigma$ is the sigmoid function and $\beta$ controls the deviation from $\pi_{\text{ref}}$.
+Group Robust Preference Optimization (GRPO): GRPO (ramesh2024group) extends preference optimization to handle diverse preference groups $\{D_{g}\}_{g=1}^{K}$.
+Instead of minimizing the average loss, GRPO minimizes the worst-case loss across groups using a robust objective:
+$\min_{\pi_{\theta}}\max_{\alpha\in\Delta_{K-1}}\sum_{g=1}^{K}\alpha_{g}L_{\text{Pref}}(\pi_{\theta};\pi_{\text{ref}},D_{g}),$ (4)
+where $L_{\text{Pref}}$ is a base preference loss (like $L_{\text{DPO}}$), and $\alpha=(\alpha_{1},...,\alpha_{K})$ are adaptive weights in the probability simplex $\Delta_{K-1}$.
+The optimization dynamically increases weights $\alpha_{g}$ for groups with higher current loss, focusing learning on the worst-performing groups.
+Simple Preference Optimization (SimPO): SimPO (meng2024simposimplepreferenceoptimization) aims to align the implicit reward with generation metrics and eliminates the need for $\pi_{\text{ref}}$.
+It uses the length-normalized average log-likelihood as the reward: $r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y|x)$.
+It also introduces a target margin $\gamma>0$ into the preference model.
+The resulting SimPO loss is:
+$L_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)\right].$ (5)
+Contrastive Preference Optimization (CPO): CPO (xu2024contrastivepreferenceoptimizationpushing) also removes the dependency on $\pi_{\text{ref}}$ for efficiency, approximating the DPO objective.
+It combines a reference-free preference loss with a negative log-likelihood (NLL) regularizer on preferred responses $y_{w}$ to maintain generation quality:
+$L_{\text{prefer}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)\right]$ (6)
+$L_{\text{NLL}}(\pi_{\theta})$ $=-\mathbb{E}_{(x,y_{w})\sim D}[\log\pi_{\theta}(y_{w}|x)]$ (7)
+$L_{\text{CPO}}(\pi_{\theta})$ $=L_{\text{prefer}}+L_{\text{NLL}}.$ (8)
+This formulation avoids the computational cost of the reference model pass.
+Appendix B FairPO: Detailed Methodology
+B.1 Objective for the Privileged Group: Preference-Based Discrimination
+For the privileged group $\mathcal{P}$, simply adjusting the weight of a standard BCE loss is insufficient.
+The primary challenge for these labels is often not basic classification, but fine-grained discrimination against closely related, incorrect alternatives.
+To address this directly, we reformulate the learning objective for this group as a preference learning task as motivated in our introduction.
+Defining Confusing Counterparts.
+For a given instance $\mathbf{x}_{i}$ and a privileged label $l\in\mathcal{P}$, we dynamically identify a set of confusing counterparts based on the model’s current predictions.
+We define two such sets:
+  • When the true label is positive ($y_{il}=1$), the set of confusing negatives is composed of incorrect labels that the model scores higher than or equal to the true label:
+    $S_{il}^{\text{neg}}=\{k\in\mathcal{T}\mid y_{ik}=0\text{ and }m(\mathbf{x}_{i};\mathbf{w}_{k})\geq m(\mathbf{x}_{i};\mathbf{w}_{l})\}.$ (9)
+  • When the true label is negative ($y_{il}=0$), the set of confusing positives is composed of correct labels that the model scores lower than or equal to the true negative:
+    $S_{il}^{\text{pos}}=\{k\in\mathcal{T}\mid y_{ik}=1\text{ and }m(\mathbf{x}_{i};\mathbf{w}_{k})\leq m(\mathbf{x}_{i};\mathbf{w}_{l})\}.$ (10)
+The overall confusing set for $(i,l)$ is denoted by $S_{il}=S_{il}^{\text{neg}}\cup S_{il}^{\text{pos}}$.
+Conditional Objective with a Stablility Fallback.
+Our objective for the privileged group is designed to be adaptive.
+When a confusing set $S_{il}$ is non-empty, we apply a preference loss $\mathcal{L}_{\text{pref}}$ to directly target these hard discriminative cases.
+However, as the model learns and becomes more accurate, the confusing set for many examples will naturally become sparse or empty.
+Relying solely on the preference loss in such a scenario would lead to an unstable, sparse gradient signal, effectively stalling the training process.
+To ensure continuous and stable learning, we incorporate a crucial stability fallback.
+If $S_{il}=\emptyset$ for a given instance, we revert to a standard Binary Cross-Entropy (BCE) loss:
+$\ell_{\text{BCE}}(\mathbf{x}_{i},y_{il};\mathbf{w}_{l})=-[y_{il}\log m(\mathbf{x}_{i};\mathbf{w}_{l})+(1-y_{il})\log(1-m(\mathbf{x}_{i};\mathbf{w}_{l}))].$ (11)
+This fallback mechanism is a critical design choice for robustness.
+Initially, the model makes many mistakes, leading to large confusing sets and frequent application of the preference loss.
+As training progresses, the preference loss successfully resolves these hard cases, and the BCE fallback takes over to fine-tune the decision boundaries on the now easier examples.
+Our analysis of the training dynamics confirms that a sufficient preference signal exists throughout training, with the fallback rate gradually increasing as the model improves.
+The total loss for the privileged group, $\mathcal{L}_{\mathcal{P}}$, is the expectation over this adaptive, conditional choice.
+In our framework, we explore several modern preference optimization techniques to instantiate $\mathcal{L}_{\text{pref}}$.
+FairPO-DPO Variant.
+To define our preference losses consistently, let us denote the score of the preferred label in a pair as $m(\mathbf{x}_{i};\mathbf{w}_{p})$ and the score of the dispreferred label as $m(\mathbf{x}_{i};\mathbf{w}_{d})$.
+The assignment of these roles depends on the ground truth.
+For example, if $y_{il}=1$ and $k\in S_{il}^{\text{neg}}$, the true label $l$ is preferred ($p=l$) and the confusing negative $k$ is dispreferred ($d=k$).
+Conversely, if $y_{il}=0$ and $k\in S_{il}^{\text{pos}}$, the true negative $l$ is preferred ($p=l$) and the confusing positive $k$ is dispreferred ($d=k$).
+Inspired by DPO (rafailov2023direct), this variant computes a preference loss relative to a reference model $\hat{\mathbf{w}}$.
+The preference loss is the negative log-likelihood of the preference, modeled by the difference between the log-probability ratios:
+$\mathcal{L}_{\text{pref}}^{\text{DPO}}(\mathbf{x}_{i},p,d)=-\log\sigma\left(\beta\left(\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{p})}{m(\mathbf{x}_{i};\hat{\mathbf{w}}_{p})}-\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{d})}{m(\mathbf{x}_{i};\hat{\mathbf{w}}_{d})}\right)\right).$ (12)
+The full DPO-based loss for the privileged group, $\mathcal{L}_{\mathcal{P}}^{\text{DPO}}$, combines this preference loss with the BCE fallback:
+$\mathcal{L}_{\mathcal{P}}^{\text{DPO}}=\mathbb{E}_{(\mathbf{x}_{i},l)\text{ s.t.
+}l\in\mathcal{P}}\left[\mathbf{1}_{S_{il}\neq\emptyset}\cdot\mathcal{L}_{\text{pref}}^{\text{DPO}}(\mathbf{x}_{i},p,d)+\mathbf{1}_{S_{il}=\emptyset}\cdot\ell_{\text{BCE}}(\mathbf{x}_{i},y_{il};\mathbf{w}_{l})\right],$ (13)
+where for each sampled privileged label $l$ with a non-empty confusing set, a confusing counterpart $k$ is sampled from $S_{il}$, and the pair $(p,d)$ is determined based on $(l,k)$ and the ground truth to compute the loss.
+FairPO-SimPO Variant (Reference-Free with Margin).
+This variant adapts SimPO (meng2024simposimplepreferenceoptimization) to create a reference-free preference loss that incorporates an explicit target margin $\gamma>0$.
+We adapt SimPO’s core concept, which was originally designed for sequences, to our per-label score setting.
+The resulting preference loss is:
+$\mathcal{L}_{\text{pref}}^{\text{SimPO}}(\mathbf{x}_{i},p,d)=-\log\sigma\left(\beta\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{p})}{m(\mathbf{x}_{i};\mathbf{w}_{d})}-\gamma\right).$ (14)
+The term $\beta\log(\cdot)$ captures the relative preference between the scores, while the margin term $\gamma$ enforces a stronger separation, pushing the model to not just rank the labels correctly but to do so by a significant amount.
+This preference loss is used in place of the DPO term within the overall conditional objective for the privileged group (analogous to Eq.
+13).
+FairPO-CPO Variant (Reference-Free with BCE Regularization).
+Our final and most effective variant adapts CPO (xu2024contrastivepreferenceoptimizationpushing).
+This approach is unique in that it integrates a BCE regularizer directly into the preference objective to ensure model stability and a sense of absolute correctness.
+When a confusing set is found, the preference loss is defined as:
+$\mathcal{L}_{\text{pref}}^{\text{CPO}}(\mathbf{x}_{i},p,d,l)=-\log\sigma\left(\beta\log\frac{m(\mathbf{x}_{i};\mathbf{w}_{p})}{m(\mathbf{x}_{i};\mathbf{w}_{d})}\right)+\lambda_{\text{CPO}}\cdot\ell_{\text{BCE}}(\mathbf{x}_{i},y_{il};\mathbf{w}_{l}).$ (15)
+For each sampled label $l$, the pair $(p,d)$ is determined from a confusing counterpart $k\in S_{il}$ if the set is non-empty.
+Here, the first term is a margin-free preference objective, while the second term, weighted by a hyperparameter $\lambda_{\text{CPO}}$, is the standard BCE loss for the privileged label $l$.
+This BCE component is crucial, acting as a regularizer that grounds the model in learning absolute scores.
+The full privileged loss objective is then analogous to the DPO formulation (Eq.
+13): the comprehensive $\mathcal{L}_{\text{pref}}^{\text{CPO}}$ is used when a confusing set is found, and only the standard BCE term (Eq.
+11), is used as a fallback otherwise.
+Advantages of the Preference-Based Objective.
+Our choice to adapt modern preference optimization techniques is motivated by the fundamental limitations of standard point-wise losses like BCE and Focal Loss for fine-grained discrimination.
+These losses evaluate each label’s score independently, answering the question, “Is the score for label $l$ correct in isolation?".
+While effective for general classification, this approach provides an indirect and often insufficient signal for resolving confusion between two closely-scored labels (lin2018focallossdenseobject).
+For example, a model might correctly learn to assign a high score (e.g., 0.8) to a true positive label and a slightly lower score (e.g., 0.7) to a confusing negative.
+A point-wise loss would provide only a small error signal for this negative case.
+The preference-based objective in FairPO is fundamentally different.
+It operates on pairs of labels and directly answers a more relational question: “Is the score for the correct label $l$ decisively higher than the score for its specific confusing competitor kk?".
+By optimizing the log-ratio of these scores, our framework generates a strong, targeted gradient to explicitly drive their values apart.
+This provides a far more direct and effective mechanism for resolving the most challenging discriminative cases within the privileged group, a capability that point-wise losses lack by design.
+B.2 Objective for the Non-Privileged Group: Constrained Performance
+For the non-privileged group $\bar{\mathcal{P}}$, the primary objective is different.
+While the model dedicates its capacity to improving the challenging privileged labels, we must ensure that this focus does not come at an unacceptable cost to the performance on the remaining labels.
+Therefore, our goal for this group is not to aggressively maximize performance, but to preserve it by preventing significant degradation relative to a reliable baseline.
+To achieve this, we introduce a constrained objective that leverages a pre-trained reference model with parameters $\hat{\mathbf{w}}$.
+We use the standard BCE loss (Eq.
+11), as our objective for a given label $j\in\bar{\mathcal{P}}$.
+The model is penalized only if the BCE loss of the current model, $\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\mathbf{w}_{j})$, exceeds the loss of the reference model, $\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\hat{\mathbf{w}}_{j})$, by more than a small, predefined slack margin $\epsilon\geq 0$.
+This is implemented using a hinge loss mechanism:
+$\mathcal{L}_{\bar{\mathcal{P}}}=\mathbb{E}_{(\mathbf{x}_{i},j)\text{ s.t.
+}j\in\bar{\mathcal{P}}}\left[\max(0,\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\mathbf{w}_{j})-\ell_{\text{BCE}}(\mathbf{x}_{i},y_{ij};\hat{\mathbf{w}}_{j})-\epsilon)\right].$ (16)
+The gradient for any non-privileged label is zero as long as its performance is good enough (i.e., close to or better than the reference model).
+A learning signal is generated only when performance on a label $j$ drops below this safety threshold.
+B.3 Group Robust Optimization Formulation
+Having defined two distinct objectives, a preference-based loss $\mathcal{L}_{\mathcal{P}}$ for the privileged group and a constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}$ for the non-privileged group, the final challenge is to balance them.
+As these objectives are often in competition, a simple weighted sum is insufficient.
+Instead, we formulate the overall training objective as a minimax game, adapting the principles of Group Robust Preference Optimization (GRPO) (ramesh2024group).
+We treat the privileged and non-privileged label sets as two distinct groups.
+The goal is to train a model, parameterized by $\mathbf{w}$, that is robust to the worst-case distribution of losses across these groups.
+This is expressed as the following minimax problem:
+$\min_{\{\mathbf{w}_{t}\}}\max_{\alpha\in\Delta}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\bar{\mathcal{P}}}\right],$ (17)
+where $\mathcal{L}_{\mathcal{P}}$ and $\mathcal{L}_{\bar{\mathcal{P}}}$ represent the expected losses for their respective groups, and $\alpha=(\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}})$ is a vector of adaptive weights in the probability simplex $\Delta$.
+This formulation provides a principled way to manage the performance trade-off between the two groups.
+Appendix C Extended Related Works
+Fairness in Multi-Label Classification: Ensuring fairness in MLC (zhang2021survey; tantithamthavorn2018impact) is complex due to multi-faceted label relationships.
+Recent efforts address MLC-specific fairness challenges, such as tackling label imbalance impacting tail labels (wu2023longtailed), learning instance and class-level subjective fairness (gao2023learning), or incorporating fairness in dynamic learning settings like class-incremental MLC (masud2023fairnessaware).
+FairPO contributes by explicitly partitioning labels into privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets, applying distinct fairness-motivated objectives to each—notably using preference signals for $\mathcal{P}$—and managing them via a robustness framework.
+This targeted approach to enhancing performance for pre-defined critical labels, while safeguarding others, differentiates our work.
+Preference Optimization: Preference optimization, especially for aligning LLMs (ouyang2022training; christiano2017deep), has rapidly advanced.
+Direct Preference Optimization (DPO) (rafailov2023direct) and its reference-free variants like CPO (xu2024contrastivepreferenceoptimizationpushing) and SimPO (meng2024simposimplepreferenceoptimization) optimize policies directly from preference pairs.
+The field continues to evolve with methods like Identity Preference Optimization (IPO) for stability (azhar2023ipo), Kahneman-Tversky-based optimization (KTO) (ethayarajh2024kto), simple yet effective Rejection Sampling Fine-Tuning (RFT) (dong2023rejection), and ongoing theoretical analyses of these methods (zhao2024whyisdpo).
+We use preferences not to rank entire outputs, but to specifically differentiate true label scores from those of their confusing positive/negative counterparts within the privileged label set, thereby sharpening the model’s decision boundaries for critical distinctions.
+Group Robust Optimization: Distributionally Robust Optimization (DRO) (ben2009robust), particularly Group DRO (sagawa2019distributionally), aims to improve worst-case performance across predefined data groups, enhancing fairness and robustness.
+This concept sees continued development, for instance, in improving its practicality (zhang2022fromaverage).
+Group Robust Preference Optimization (GRPO) (ramesh2024group) extended this to LLM preference learning, balancing performance across preference groups.
+FairPO directly employs GRPO’s adaptive optimization strategy.
+However, our groups are defined by the label partition ($\mathcal{P}$ and $\bar{\mathcal{P}}$), and GRPO balances their distinct, custom-formulated loss objectives.
+This provides a principled mechanism for robustly managing the specific fairness-performance trade-offs in our MLC context.
+Appendix D Adapting FairPO for Multi-Attribute Generation
 This section outlines our planned extension of FairPO to multi-attribute generation, a conceptual direction for future work.
@@
 The core idea involves partitioning $\mathcal{A}$ into privileged $\mathcal{P}$ and non-privileged $\bar{\mathcal{P}}$ sets and retaining the GRPO minimax structure (Eq.
-18).
+17).
 The group losses would be defined over a preference dataset $\mathcal{D}_{pref}=\{(x_{i},y_{wi},y_{li},j_{i})\}_{i=1}^{M}$, with preference losses like DPO applied to the log-probabilities of entire generated sequences rather than individual label scores.
@@
 The overall FairPO objective would then use GRPO to balance these two losses.
-Planned Experiments: The planned task is generating text from a prompt, requiring preference datasets where the driving attribute $j_{i}$ is known.
-This could involve partitioning attributes like ‘Helpfulness’ or ‘Harmlessness’ from existing RLHF datasets (e.g., Anthropic HH-RLHF [Bai and Others, 2022], Summarization preferences [Stiennon et al., 2022]) or creating new, explicitly annotated data.
-We plan to fine-tune pretrained language models (e.g., Gemma [Team, 2024], Llama [Touvron and Others, 2023], Mistral [Jiang and Others, 2023]) with this objective.
-Evaluation would rely on preference-based metrics, such as win rates for privileged and non-privileged attribute groups, compared against baselines like standard DPO and SFT.
-Future experiments would also explore hyperparameter effects and adapting the SimPO/CPO loss formulations for generation.
-Appendix B FairPO Algorithm
+Appendix E FairPO Algorithm
 The FairPO framework is trained iteratively to solve the minimax objective presented in Eq.
-18.
+17.
 The detailed procedure, which is inspired by the DPO-based variant of FairPO, is provided in Algorithm  2.
@@
   • If no confusing examples are found ($S_{il}=\emptyset$), the algorithm reverts to a standard base classification loss (e.g., BCE, Eq.
-10) for label $l$.
+11) for label $l$.
 This fallback is crucial as it ensures the model continues to receive a learning signal on easier instances, promoting stable training.
@@
 Conversely, if the sampled label $r$ belongs to the non-privileged set $\bar{\mathcal{P}}$, the constrained loss $\mathcal{L}_{\bar{\mathcal{P}}}^{(s)}$ is computed according to Eq.
-17.
+16.
 This loss penalizes the model only if its performance on the label $j=r$ deviates from the reference model’s performance by more than a predefined slack margin $\epsilon$.
@@
 For variants like FairPO-SimPO or FairPO-CPO, the core logic remains identical; only the DPO-inspired preference loss component is replaced with their respective preference formulations (e.g., Eq.
-12 or 13).
+14 or 15).
 The overall GRPO structure and non-privileged handling are consistent across all variants.
@@
 40:return $\{\mathbf{w}_{t}^{(S)}|\forall t\in\mathcal{T}\}$
-Appendix C Dataset and Preprocessing Details
-MS-COCO 2014 [Lin et al., 2015]: We used the official 2014 train/val splits.
+Appendix F Dataset and Preprocessing Details
+MS-COCO 2014 (lin2015microsoftcococommonobjects): We used the official 2014 train/val splits.
 The training set contains 82,783 images and the validation set (used as our test set) contains 40,504 images.
@@
 The remaining 64 labels formed $\bar{\mathcal{P}}$.
-NUS-WIDE [Chua et al., 2009]: This dataset contains 269,648 images with 81 concept labels.
+NUS-WIDE (chua2009nus): This dataset contains 269,648 images with 81 concept labels.
 We used the common split of 161,789 images for training and 107,859 for testing.
@@
 Standard data augmentations like random horizontal flips and random resized crops were applied during training.
-Appendix D Baseline Experimental Details
-This section provides further details on the implementation and hyperparameter tuning for the baseline methods used in our experiments.
-Unless otherwise specified, all baselines were trained using the same base Vision Transformer (ViT) architecture, optimizer (AdamW), number of epochs, and batch size as the main FairPO experiments for fair comparison.
-Hyperparameters specific to each baseline were tuned on the validation set of MS-COCO and NUS-WIDE.
-The privileged group $\mathcal{P}$ comprised the 20% least frequent labels, and the non-privileged group $\bar{\mathcal{P}}$ comprised the remaining 80%.
-D.1 BCE-SFT (Reference Model)
-The BCE-SFT baseline represents standard supervised fine-tuning.
-Its objective is to minimize the sum of independent Binary Cross-Entropy (BCE) losses for each of the $T$ labels, where the loss for an instance $(x_{i},y_{i})$ is $\mathcal{L}_{\text{BCE-SFT}}=\sum_{t=1}^{T}\text{BCE}(m(x_{i};\mathbf{w}_{t}),y_{it})$.
-For implementation, a separate non-linear MLP classifier head is trained for each label $t$ on top of frozen ViT features, consistent with the main FairPO architecture.
-The primary hyperparameter tuned was the learning rate for these classifier heads, selected from $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$ to achieve the best overall mean Average Precision (mAP) on the validation set.
-Weight decay was set to 0.01, matching the FairPO experiments.
-This model serves two critical roles in our study.
-First, it acts as a direct performance baseline against which we measure fairness improvements.
-Second, the final trained parameters $\{\mathbf{w}_{t}\}$ from this SFT process become the reference parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}=\{\hat{\mathbf{w}}_{t}\}$ used throughout the FairPO framework, both for the DPO-based losses and for the non-privileged loss constraint.
-The model was trained for the same number of epochs as FairPO, with early stopping based on overall validation mAP.
-D.2 BCE-SFT + Privileged Re-Weighting (RW)
-This baseline modifies the standard BCE-SFT by applying static loss re-weighting to improve performance on the privileged labels $\mathcal{P}$.
-The objective is to assign a higher weight to their contribution in the total loss function:
-$\mathcal{L}_{\text{RW}}=\lambda_{\mathcal{P}}\sum_{l\in\mathcal{P}}\text{BCE}(m(x_{i};\mathbf{w}_{l}),y_{il})+\lambda_{\bar{\mathcal{P}}}\sum_{j\in\bar{\mathcal{P}}}\text{BCE}(m(x_{i};\mathbf{w}_{j}),y_{ij})$
-For this setup, we use the same per-label classifier architecture as BCE-SFT, but the loss computation is modified to incorporate the weights.
-The key hyperparameter is the privileged group weight $\lambda_{\mathcal{P}}$, which was tuned via grid search from values $\{2,3,5,8,10\}$.
-The value was chosen to maximize mAP on the privileged group without an excessive drop in performance on the non-privileged group.
-For our reported results, a weight of $\lambda_{\mathcal{P}}=5$ was found to be effective, while $\lambda_{\bar{\mathcal{P}}}$ was fixed at $1$.
-Training was conducted similarly to BCE-SFT.
-D.3 Group DRO + BCE
-This baseline applies the Group Distributionally Robust Optimization technique to the standard BCE loss.
-The objective is to minimize the worst-case BCE loss across the predefined privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) groups:
-$\min_{\{\mathbf{w}_{t}\}}\max_{\alpha_{\mathcal{P}},\alpha_{\bar{\mathcal{P}}}\geq 0,\alpha_{\mathcal{P}}+\alpha_{\bar{\mathcal{P}}}=1}\left[\alpha_{\mathcal{P}}\mathcal{L}_{\text{BCE}}(\mathcal{P})+\alpha_{\bar{\mathcal{P}}}\mathcal{L}_{\text{BCE}}(\bar{\mathcal{P}})\right]$
-Here, $\mathcal{L}_{\text{BCE}}(\mathcal{G})$ is the average BCE loss for all labels within a group $\mathcal{G}$.
-The implementation uses the same BCE-SFT architecture.
-Per-label BCE losses are computed and then aggregated separately for the $\mathcal{P}$ and $\bar{\mathcal{P}}$ groups.
-Model parameters $\{\mathbf{w}_{t}\}$ are updated with AdamW, while the group weights $\alpha_{\mathcal{P}}$ and $\alpha_{\bar{\mathcal{P}}}$ are updated iteratively using a mirror ascent step: $\alpha_{\mathcal{G}}^{(s+1)}\propto\alpha_{\mathcal{G}}^{(s)}\exp(\eta_{\alpha}\cdot\mathcal{L}_{\text{BCE}}^{(s)}(\mathcal{G}))$.
-The final loss for backpropagation is the adaptively weighted sum of the group losses.
-For hyperparameters, the model learning rate was tuned as for BCE-SFT.
-The group weight learning rate $\eta_{\alpha}$ was tuned from $\{0.01,0.05,0.1,0.2\}$, with $\eta_{\alpha}=0.05$ selected for its stable convergence and good worst-group performance.
-For the reported results, raw group losses were used in the exponential update, though we note that a loss scaling approach similar to that in FairPO could also be employed for enhanced stability.
-The training protocol regarding epochs and batching was identical to FairPO.
-Appendix E FairPO Experimental Details
-E.1 Common Setup for All FairPO Variants
+Appendix G FairPO Experimental Details
+G.1 Common Setup for All FairPO Variants
 Unless specified otherwise, a common setup was used for all FairPO variants to ensure fair comparison.
-The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 [Dosovitskiy et al., 2021], which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k.
+The base model for feature extraction was a Vision Transformer (ViT), specifically vit-base-patch16-224 (dosovitskiy2021imageworth16x16words), which was pre-trained on ImageNet-21k and fine-tuned on ImageNet-1k.
 During our fine-tuning, the ViT backbone was kept frozen, with the exception of its final encoder layer, which was made trainable to allow for adaptation of higher-level features.
-All experiments were conducted on the MS-COCO 2014 [Lin et al., 2015] and NUS-WIDE [Chua et al., 2009] datasets.
+All experiments were conducted on the MS-COCO 2014 (lin2015microsoftcococommonobjects) and NUS-WIDE (chua2009nus) datasets.
 Images were resized to $224\times 224$ pixels, normalized using ImageNet statistics, and augmented with standard techniques like random horizontal flips and resized crops.
-The AdamW optimizer [Loshchilov and Hutter, 2019] was used to update all trainable parameters.
+The AdamW optimizer (loshchilov2019decoupled) was used to update all trainable parameters.
 Models were trained for a maximum of 25 epochs with a batch size of 32, and we employed an early stopping strategy with a patience of 5 epochs based on the overall mAP on the validation set.
-The reference model parameters $\{\hat{\mathbf{w}}_{t}|t\in\mathcal{T}\}$, required for FairPO-DPO and the non-privileged loss constraint in all variants, were obtained from a BCE-SFT model detailed in Appendix D.
-E.2 Per-Label Non-Linear MLP Classifier Head
+G.2 Per-Label Non-Linear MLP Classifier Head
 For each of the $T$ labels in a dataset, we employed a dedicated and independent non-linear Multi-Layer Perceptron (MLP) head to predict the probability of that label being positive.
@@
 All parameters within these MLP heads were fully trainable during both the SFT pre-training (for the reference model) and the final FairPO fine-tuning.
-E.3 FairPO Variant-Specific Hyperparameters
-The core FairPO learning rates, $\eta_{\mathbf{w}}$ for model parameters and $\eta_{\alpha}$ for GRPO weights, were tuned via grid search.
-For $\eta_{\mathbf{w}}$ (the AdamW LR), values were explored from $\{1e-5,5e-5,1e-4,5e-4\}$, and for $\eta_{\alpha}$, values were explored from $\{0.01,0.05,0.1,0.2\}$.
-After tuning, the final reported results used $\eta_{\mathbf{w}}=1e-4$ and $\eta_{\alpha}=0.05$.
-The specific hyperparameters for each FairPO variant’s privileged loss component (e.g., $\beta$, $\gamma$, $\lambda_{\text{CPO}}$) were also tuned on the designated validation set for each dataset, with final values reported in Table 5 (Appendix F).
-Throughout all experiments, the GRPO mechanism with scaled loss updates, as described in Section 3, was used to adaptively balance the privileged and non-privileged loss terms.
-Appendix F Hyperparameter Details
-F.1 Hyperparameters Tuning
-For all experiments, we used the AdamW optimizer [Loshchilov and Hutter, 2019] with a batch size of 32.
-The base model was a vit-base-patch16-224222https://huggingface.co/docs/transformers/en/model_doc/vit pretrained on ImageNet-21k and fine-tuned on ImageNet-1k.
-The initial learning rate for model parameters ($\eta_{\mathbf{w}}$) was selected from $\{1e-5,5e-5,1e-4,5e-4\}$, while the learning rate for GRPO’s alpha weights ($\eta_{\alpha}$) was selected from $\{0.01,0.05,0.1,0.2\}$.
-All models were trained for a maximum of 25 epochs with early stopping based on the validation set’s overall mAP, using a patience of 5 epochs [^3] All experiments were conducted on a machine equipped with NVIDIA A100 GPUs (80GB VRAM)..
-Specific hyperparameters for each FairPO variant were tuned via grid search on the validation set.
-For FairPO-DPO, the strength parameter $\beta$ was chosen from $\{0.1,0.5,1.0\}$, and the non-privileged slack $\epsilon$ was chosen from $\{0.01,0.05,0.1\}$.
-For FairPO-SimPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the margin $\gamma$ from $\{0.05,0.1,0.2\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$.
-For FairPO-CPO, the preference scaling $\beta$ was chosen from $\{0.1,0.5,1.0\}$, the NLL regularizer weight $\lambda_{\text{CPO}}$ from $\{0.1,0.5,1.0\}$, and the slack $\epsilon$ from $\{0.01,0.05,0.1\}$.
-The final selected hyperparameters for each dataset and FairPO variant are reported in Table 5.
-For our baselines, the loss weight for privileged labels in BCE-SFT + Privileged Re-Weighting was set to 5 after tuning from $\{2,3,5,10\}$.
-The group weight learning rate $\eta_{\alpha}$ for GDRO + BCE was tuned similarly to FairPO.
-F.2 Sensitivity to Key Hyperparameters $\beta$ and $\epsilon$
-We investigated the sensitivity of FairPO-CPO’s performance on MS-COCO to variations in the preference strength parameter $\beta$ (from Eq. 13) and the non-privileged slack $\epsilon$ (from Eq. 17).
-Table 5: Final selected hyperparameters for FairPO variants on MS-COCO and NUS-WIDE.
-Note: For FairPO-CPO, the first value under $\beta/\lambda_{\text{CPO}}$ is $\beta$ and the second is $\lambda_{\text{CPO}}$.
-Base learning rate $\eta_{\mathbf{w}}$ was 1e41e-4 and $\eta_{\alpha}$ was $0.05$ for all reported results after tuning.
-Dataset Method $\beta$ / $\lambda_{\text{CPO}}$ $\gamma$ (SimPO only) $\epsilon$
-MS-COCO FairPO-DPO $\beta=0.5$ N/A $0.05$
-FairPO-SimPO $\beta=0.5$ $0.1$ $0.05$
-FairPO-CPO $\beta=0.5,\lambda_{\text{CPO}}=0.5$ N/A $0.01$
-NUS-WIDE FairPO-DPO $\beta=0.1$ N/A $0.05$
-FairPO-SimPO $\beta=0.1$ $0.05$ $0.1$
-FairPO-CPO $\beta=0.1,\lambda_{\text{CPO}}=0.5$ N/A $0.05$
-Table 6 shows the mAP for privileged ($\mathcal{P}$) and non-privileged ($\bar{\mathcal{P}}$) sets as $\beta$ varies, keeping other hyperparameters (including $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$) fixed to their optimal values.
-Performance is relatively stable across a range of $\beta$, though very small or very large values can lead to suboptimal results.
-Table 6: Sensitivity of FairPO-CPO mAP on MS-COCO to preference strength $\beta$ (with $\epsilon=0.01,\lambda_{\text{CPO}}=0.5$).
-$\beta$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
-0.05 88.95 90.28
-0.1 89.32 90.31
-0.5 (Optimal) 89.76 90.34
-1.0 89.51 90.25
-2.0 88.67 90.11
-Table 7 shows the mAP as $\epsilon$ varies, keeping other hyperparameters (including $\beta=0.5,\lambda_{\text{CPO}}=0.5$) fixed.
-A moderate $\epsilon$ helps balance the objectives; too small an $\epsilon$ can be overly restrictive on $\bar{\mathcal{P}}$ performance, while too large an $\epsilon$ might allow too much degradation.
-Table 7: Sensitivity of FairPO-CPO mAP on MS-COCO to non-privileged slack $\epsilon$ (with $\beta=0.5,\lambda_{\text{CPO}}=0.5$).
-$\epsilon$ mAP($\mathcal{P}$) mAP($\bar{\mathcal{P}}$)
-0.001 89.65 90.40
-0.01 (Optimal) 89.76 90.34
-0.05 89.81 90.22
-0.1 89.70 90.05
-0.2 89.55 89.87